### v1

# A Closer Look at TabPFN v2: Strength, Limitation, and Extension

Abstract
Tabular datasets are inherently heterogeneous, posing significant challenges for developing pre-trained foundation models. The recently introduced transformer-based Tabular Prior-data Fitted Network v2 (TabPFN v2) achieves unprecedented in-context learning accuracy across multiple tabular datasets, marking a pivotal advancement in tabular foundation models. In this paper, we comprehensively evaluate TabPFN v2 on over 300 datasets, confirming its exceptional generalization capabilities on small- to medium-scale tasks. Our analysis identifies randomized feature tokens as a key factor behind TabPFN v2’s success, as they unify heterogeneous datasets into a fixed-dimensional representation, enabling more effective training and inference. To further understand TabPFN v2’s predictions, we propose a leave-one-fold-out approach, transforming TabPFN v2 into a feature extractor and revealing its capability to simplify data distributions and boost accuracy. Lastly, to address TabPFN v2’s limitations in high-dimensional, large-scale, and many-category tasks, we introduce a divide-and-conquer mechanism inspired by Chain-of-Thought prompting, enabling scalable inference. By uncovering the mechanisms behind TabPFN v2’s success and introducing strategies to expand its applicability, this study provides key insights into the future of tabular foundation models.
Machine Learning, ICML
1 Introduction
Tabular data is ubiquitous across diverse applications, including healthcare (Hyland et al., 2020), finance (Kovalerchuk & Vityaev, 2005), and scientific research (Ivanciuc et al., 2007; Hyland et al., 2020). In this format, each instance is represented as a vector of attributes, and the task of a machine learning model is to map these vector inputs to their corresponding labels (Borisov et al., 2024). Traditionally, tree-based models (Prokhorenkova et al., 2018; Chen & Guestrin, 2016) have dominated the tabular domain, but deep tabular models are increasingly showing potential to advance the field and close the performance gap (Gorishniy et al., 2021; Holzmüller et al., 2024; Ye et al., 2025).
However, unlike vision and language data, where pre-trained foundation models have driven significant progress (Kirillov et al., 2023; Zhou et al., 2024), tabular data is still desperately awaiting a similar breakthrough (Somepalli et al., 2022; Hollmann et al., 2023; Onishi et al., 2023; Zhu et al., 2023; Ye et al., 2023; van Breugel & van der Schaar, 2024). A primary challenge arises from the inherent heterogeneity of tabular datasets, which often vary in dimensionality and attribute meanings, making the development of effective and versatile foundation models difficult. Additionally, there is an urgent need for such models, as many tabular datasets are small-scale—such as medical data with limited patient numbers. Training individual models from scratch for these datasets is highly sensitive to hyperparameters and struggles with insufficient data for generalization (Feurer et al., 2015; Guyon et al., 2019; Han et al., 2024).
Recently, the Tabular Prior-Fitted Network v2 (TabPFN v2), proposed by Hollmann et al. (2025), has emerged as a significant step forward. Built on transformer architectures (Vaswani et al., 2017) and pre-trained on gigantic synthetic datasets (Hollmann et al., 2023, 2025), TabPFN v2 can be directly applied to downstream tasks without the need for additional tuning, effectively addressing the heterogeneity of tabular data. Specifically, TabPFN v2 takes both a labeled training set and an unlabeled test instance as input, predicting the test label in an “in-context learning” manner. When evaluated on multiple datasets covering both classification and regression tasks, TabPFN v2 achieved unprecedented accuracy compared to previous methods.
Motivated by the remarkable performance of TabPFN v2, this paper pursues four timely objectives to deepen our understanding and advance the field. First, we aim to expand the evaluation of TabPFN v2 across a diverse set of scenarios to investigate its generalizability. Second, we seek to understand the underlying mechanisms by which TabPFN v2 effectively handles data heterogeneity. Third, we strive to uncover how TabPFN v2 achieves its high prediction accuracy. Last but not least, we aim to extend TabPFN v2’s applicability beyond the comfort zone defined in Hollmann et al. (2025)—datasets with no more than 10,000 samples, 500 dimensions, and 10 classes.
TabPFN v2 consistently outperforms existing methods on small- to medium-scale datasets. We evaluate TabPFN v2 extensively on over 300 tabular datasets spanning various domains, scales, dimensionalities, and tasks (Grinsztajn et al., 2022; McElfresh et al., 2023; Ye et al., 2024b; Rubachev et al., 2025). Our results demonstrate TabPFN v2’s exceptional generalizability on small- to medium-scale datasets, outperforming both tree-based models and state-of-the-art deep tabular models. However, TabPFN v2’s performance diminishes on large-scale and high-dimensional datasets, indicating areas where further improvements are needed.
TabPFN v2 effectively handles data heterogeneity through randomized feature tokens. A key strength of TabPFN v2 is its ability to manage heterogeneous feature spaces, enabling applications to diverse downstream datasets without additional tuning. Given a $d$-dimensional input vector $\in\mathbb{R}^{d}$, TabPFN v2 transforms it into a $d\times k$ matrix and leverages a transformer architecture to handle variability in $d$, following (Song et al., 2019; Gorishniy et al., 2021; Yan et al., 2024). The critical innovation lies in the vector-to-matrix transformation: TabPFN v2 uses a “shared” $k$-dimensional vector to lift each attribute value into a $k$-dimensional space. To differentiate attributes, a random $k$-dimensional perturbation is added to the shared vector for each attribute, consistent within a dataset but distinct across datasets [^1] This detail was identified through the supplementary material and code of (Hollmann et al., 2025).. We view this transformation as a variant of token-based tabular methods, where TabPFN v2 removes the need for dataset- and attribute-specific feature token learning, a limitation that has hindered the transferability of previous methods to diverse downstream datasets. Our analysis confirms that these feature tokens require no semantic meaning as long as they are distinct across attributes, demonstrating their efficacy.
TabPFN v2 simplifies feature distributions among categories. To deepen our understanding of TabPFN v2, we investigate whether its superior in-context learning capability is accompanied by the creation of more separable feature representations. This analysis is challenging due to the distinct roles of labeled training data and unlabeled test data in TabPFN v2’s in-context learning process, resulting in non-comparable embeddings. To overcome this, we propose a leave-one-fold-out strategy to extract features from the training data that align more closely with the test data. Our findings reveal that TabPFN v2 effectively transforms tabular datasets into a nearly separable embedding space. Remarkably, training a linear model on TabPFN v2’s extracted features achieves accuracy comparable to its in-context learner, underscoring TabPFN v2’s potential as a feature encoder. This approach not only offers valuable insights into TabPFN v2 but also paves the way for broader applications and further analysis of its capabilities.
Divide-and-conquer extends TabPFN v2’s applicability. As noted in Hollmann et al. (2025) and confirmed by our study, TabPFN v2 encounters challenges with large-scale, high-dimensional, and many-category datasets. To address these limitations without extensive retraining, we introduce post-hoc divide-and-conquer strategies, reminiscent of Chain-of-Thought (CoT) prompting used in large-language models (Wei et al., 2022). For large-scale datasets, we split the training data into support and query sets, using the support set to extract TabPFN v2’s features for the query and test sets. The resulting features then form a new tabular dataset compatible with linear models, akin to representation learning strategies in the early days (Vincent et al., 2010). For high-dimensional datasets, we subsample attributes into subsets (Breiman, 2001), apply TabPFN v2 to each subset in parallel, and ensemble the predictions, similar to random forests. For a multi-class task with over 10 categories, we introduce decimal encoding inspired by Error Correcting Output Codes (ECOC) (Dietterich & Bakiri, 1995), decomposing it into multiple 10-class problems. Empirical results validate these strategies, demonstrating significant gains in TabPFN v2’s accuracy across these data regimes and highlighting the potential of advanced post-hoc methods to further expand the capabilities of tabular foundation models.
Contributions. This paper provides a timely and in-depth investigation into TabPFN v2’s strengths, limitations, and potential extensions, aiming to offer valuable insights to advance tabular foundation models. In particular:
  • We confirm TabPFN v2’s remarkable in-context learning capability across diverse small- to medium-scale tasks, positioning it as a new go-to method for these data regimes.
  • We highlight randomized feature tokens as the key driver behind TabPFN v2’s handling of heterogeneous datasets and show their applicability to other deep tabular models.
  • We introduce a simple leave-one-fold-out approach to turn TabPFN v2 into a feature encoder, enhancing its understanding and suggesting broader applications (e.g., for visualization and error diagnosis).
  • We demonstrate the effectiveness of post-hoc mechanisms to extend TabPFN v2 beyond its designated data regimes, similar to prompting strategies like CoT for LLMs.
2 Related Work
Foundation Tabular Models. Pre-trained models have revolutionized vision and language domains (Kirillov et al., 2023; Zhou et al., 2024), but their adoption in tabular data remains limited due to the inherent heterogeneity of tabular datasets. Differences in feature spaces, dimensionalities, and class distributions pose significant challenges for joint training and transferability. One solution is to utilize semantic meanings of features, as demonstrated by methods that transform tabular examples into textual representations for large language models (LLMs) (Hegselmann et al., 2023; Zhang et al., 2023a; Wang et al., 2023; Wen et al., 2024). Pre-computing feature tokens based on semantic embeddings has also been explored to improve transferability (Yan et al., 2024; Kim et al., 2024).
However, semantic meanings are often unavailable due to privacy concerns, annotation costs, or lack of describability. To address this, Ye et al. (2023) propose representing examples through similarity relationships with a fixed number of reference examples, mapping them into a latent space with a fixed dimension. The TabPFN family of models (Hollmann et al., 2023, 2025) leverages the in-context learning capabilities of transformers, directly predicting labels by positioning test instances within the context of training examples. While TabPFN v1 pads features to a fixed dimension, TabPFN v2 introduces a specialized feature tokenizer to better handle heterogeneity. Meta-learning has also been explored to obtain the weights of tabular models for downstream tasks with limited examples (Iwata & Kumagai, 2020; Bonet et al., 2024). Other pre-trained models require lightweight fine-tuning on downstream datasets to adapt to heterogeneous feature and class spaces (Liu et al., 2022; Zhang et al., 2023b; Shen et al., 2023; Zhu et al., 2023).
Variants of TabPFN. TabPFN’s success stems from its pre-training on gigantic synthetic datasets, enabling strong in-context learning performance on small-scale classification tasks (Hollmann et al., 2023). Inspired by its capabilities, various applications have been investigated, including tabular data generation (Ma et al., 2024a), anomaly detection (Ruiz-Villafranca et al., 2024), and time series forecasting (Hoo et al., 2025). The efficiency of TabPFN is highly sensitive to context size, prompting strategies to enhance scalability and performance (Feuer et al., 2023; Xu et al., 2024). Some approaches enhance TabPFN’s performance on downstream tasks by adapting the context (Thomas et al., 2024), fine-tuning specific parts of the model (Feuer et al., 2024; Liu & Ye, 2025), or even pre-training the architecture on real-world datasets (Ma et al., 2024b). Nagler (2023) analyzes TabPFN from a bias-variance perspective, shedding light on its generalization capabilities. The recently introduced TabPFN v2 extends support to regression tasks and handles larger contexts (Hollmann et al., 2025). In this paper, we conduct a comprehensive evaluation of TabPFN v2, analyze its strengths and limitations, and propose solutions to overcome scalability and applicability challenges.
3 Preliminary
We formalize the problem of learning with tabular datasets and briefly describe the core mechanism of TabPFN v2.
Figure 1: Illustration of the TabPFN v2 mechanism (Hollmann et al., 2025) for classification. $\{A_{1},\ldots,A_{d}\}$ denote $d$ attributes of the task. Training examples and test instances are combined into a tabular context set and transformed into a ${(N+1)\times d\times k}$ tensor using a combination of learnable and randomized tokens. Two types of self-attentions are applied alternately along rows (inter-sample) and columns (inter-feature). The output token corresponding to the (dummy) label of the test instance is processed through an MLP to generate a 10-class logit.
Learning with a Single Tabular Dataset. A tabular dataset ${\mathcal{D}}=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{N}$ contains $N$ training examples, corresponding to the rows in a table. Each instance ${\bm{x}}_{i}$ is represented by $d$ features or attributes (columns in the table); its label $y_{i}$ belongs to $[C]=\{1,\ldots,C\}$ for a classification task or is a numerical value for a regression task. We assume all attributes of an instance are numerical (continuous). If categorical (discrete) attributes are present, they are transformed with ordinal/one-hot encoding beforehand. The goal of tabular machine learning is to construct a mapping $f$ from an instance to its label. Specifically, given an unseen instance ${\bm{x}}^{*}\in{\mathbb{R}}^{d}$ sampled from the same distribution as ${\mathcal{D}}$, the learned mapping $f$ predicts its label as $\hat{y}^{*}=f({\bm{x}}^{*}\mid{\mathcal{D}})$. A lower prediction discrepancy between $\hat{y}^{*}$ and the true label $y^{*}$ indicates a stronger generalization ability of $f$.
TabPFN Variants. The original TabPFN implements $f$ for classification via a Transformer-like architecture (Hollmann et al., 2023). Both training and test instances are padded to a fixed dimension $k^{\prime}$ (e.g., 100) with zeros. Then, ${\bm{x}}_{i}$ and $y_{i}$ are further mapped to $\tilde{{\bm{x}}}_{i}\in{\mathbb{R}}^{k}$ and $\tilde{{\bm{y}}}_{i}\in{\mathbb{R}}^{k}$ with linear transformations. TabPFN takes both a labeled training set and an unlabeled test instance as input, predicting the test label in an “in-context learning” manner. The task context is formulated as:
${\mathcal{C}}=\{(\tilde{{\bm{x}}}_{1}+\tilde{{\bm{y}}}_{1}),\ldots,(\tilde{{\bm{x}}}_{N}+\tilde{{\bm{y}}}_{N}),(\tilde{{\bm{x}}}^{*})\}\in{\mathbb{R}}^{k\times(N+1)}\;,$
with $N+1$ tokens of $k$ dimensions each. These tokens are processed through multiple transformer layers, which support variable-length inputs. The output token corresponding to the test instance is further mapped to a 10-class logit using a multi-layer perceptron (MLP).
The recently introduced TabPFN v2 incorporates several modifications to the original TabPFN. For example, each attribute in ${\mathcal{D}}$ is mapped to a $k$-dimensional space through linear projection. Random perturbations are added to differentiate attributes in the $k$-dimensional space, similar to a kind of positional encoding. The core idea is illustrated in Figure 1. We view this strategy as a variant of the token-based method (Song et al., 2019; Gorishniy et al., 2021; Yan et al., 2024) and will discuss the construction of tokens and analyze how they handle the challenges of heterogeneous attribute spaces in section 5. Together with the mapped label embedding $\tilde{{\bm{y}}}_{i}$, an instance ${\bm{x}}_{i}$ is transformed into tokens with size $(d+1)\times k$. For a test instance ${\bm{x}}^{*}$, since its label is unknown, a dummy label (e.g., the average of training set labels) is used to generate the label embedding $\tilde{{\bm{y}}}^{*}$.
Subsequently, the set of training and test instances is transformed into a ${(N+1)\times(d+1)\times k}$ tensor. Two types of self-attentions are alternately applied for inter-sample and inter-feature in-context learning. Finally, the output token corresponding to the test instance’s dummy label $\tilde{{\bm{y}}}^{*}$ is extracted and mapped to either a 10-class logit for classification or a single-value logit for regression.
The weights in TabPFN v2 are pre-trained on synthetic datasets generated using structural causal models (SCMs); the checkpoint is selected based on real-world datasets. This pre-training enables TabPFN v2 to be directly applied to downstream tasks without additional tuning, effectively addressing the heterogeneity of tabular data. However, the complexity of transformers with respect to the input size limits TabPFN v2’s scalability to small and medium datasets (e.g., $N<10,000$) and its ability to handle classification with fewer than 10 classes. Please refer to Hollmann et al. (2025) for further details, including feature pre-processing, feature grouping, acceleration, and post-hoc ensembling.
4 Comprehensive Evaluation of TabPFN v2
Building upon the remarkable performance of TabPFN v2, we conduct an extensive evaluation across diverse scenarios to assess its generalizability. Specifically, we evaluate TabPFN v2 on over 300 tabular datasets spanning a wide range of domains, scales, dimensionalities, and tasks.
Figure 2: Wilcoxon-Holm test at a significance level of 0.05 over 261 small- to medium-scale datasets. We omit the 12 datasets with more than 10 classes and the 27 datasets used to select the TabPFN v2 checkpoint from the 300 datasets in (Ye et al., 2024b).
4.1 Setups
To ensure a comprehensive evaluation, we adopt the benchmark proposed in Ye et al. (2024b), which includes 300 datasets covering 120 binary classification tasks, 80 multi-class classification tasks, and 100 regression tasks across various domains, dataset sizes, and feature types. This benchmark addresses common issues, such as mislabeled datasets and dataset redundancies due to multi-version overlaps (Kohli et al., 2024), ensuring high-quality evaluations.
Out of the 300 datasets, 12 classification datasets contain more than 10 classes, and 27 datasets overlap with the validation set used for checkpoint selection during TabPFN v2’s pre-training (Hollmann et al., 2025). To avoid evaluation bias, we exclude these datasets and report comparison results on the remaining 261 datasets.
Following the protocol in Gorishniy et al. (2021, 2024), each dataset is randomly split into training, validation, and test partitions in proportions to 64%/16%/20%. TabPFN v2 predicts the test set labels directly in an in-context learning fashion without further parameter and hyper-parameter tuning. Compared (deep) tabular methods perform hyper-parameter tuning using Optuna (Akiba et al., 2019) with 100 trials over the training set, and are early stopped on the validation set. All methods are evaluated with 15 random seeds, and the average performance across seeds is reported.
For classification tasks, we use accuracy (higher is better) as the evaluation metric, while regression tasks are evaluated using Root Mean Square Error (RMSE, lower is better).
Figure 3: Probability of Achieving the Best Accuracy/RMSE (PAMA) across 261 small- to medium-scale datasets. The values inside the rectangles represent the percentage of datasets where each method achieves the best accuracy.
4.2 Results
We compare TabPFN v2 against 26 representative (deep) tabular methods on 261 datasets (detailed references are in the appendix). To assess statistical significance, we apply the Wilcoxon-Holm test at a 0.05 significance level (Demsar, 2006). The critical difference analysis results are presented in Figure 2.
The results demonstrate that TabPFN v2 consistently outperforms both tree-based methods, such as CatBoost (Prokhorenkova et al., 2018), and deep tabular models, including RealMLP (Holzmüller et al., 2024), ModernNCA (Ye et al., 2025), TabR (Gorishniy et al., 2024), and FT-Transformer (Gorishniy et al., 2021), with statistically significant differences.
To further analyze performance, we consider the Probability of Achieving the Best Accuracy/RMSE (PAMA) (Delgado et al., 2014), which measures the percentage of datasets where a method achieves the best performance. As shown in Figure 3, TabPFN v2 achieves the highest PAMA score, with 25.06% of datasets yielding the best results, outperforming other methods like ModernNCA (with a percentage of 13.73%) and CatBoost (9.48%). These results validate TabPFN v2’s exceptional generalizability on small- to medium-scale tabular datasets, further enhanced by its efficiency and lack of reliance on hyper-parameter tuning.
Table 1: Average rank (lower is better) of TabPFN v2 and representative comparison methods on 18 high-dimensional datasets and 18 large-scale datasets. Full results with our extensions are in Figure 6 and Figure 8.
$\downarrow$ TabPFN v2 CatBoost MNCA RealMLP LR
High-Dim 3.36 2.82 4.41 2.14 2.27
Large-Scale 3.97 1.89 2.27 1.94 4.47
4.3 Additional Evaluations
To examine TabPFN v2’s performance on larger and more complex datasets, we conduct additional evaluations on 18 large-scale datasets with $N\times d>1,000,000$ and 18 high-dimensional datasets with $d\geq 2,000$ (Jiang et al., 2024). For high-dimensional datasets, we follow the same evaluation protocol as before. For large-scale datasets, due to the prohibitive cost of hyper-parameter tuning, default hyper-parameters are used for all methods. The average ranks of some representative methods are summarized in Table 1. Full results are in section 7 with our extensions.
While TabPFN v2 performs superiorly on small- to medium-scale datasets, its performance degrades on large-scale and high-dimensional datasets. For instance, on large-scale datasets, TabPFN v2 ranks lower than both CatBoost and RealMLP, highlighting its scalability limitations. On high-dimensional datasets, TabPFN v2 fails to outperform even the simple Logistic Regression (LR) model, further emphasizing the need for enhancements to better handle these challenging scenarios.
These results highlight TabPFN v2’s limitations in handling high-dimensional and large-scale datasets, which may result from two key factors. First, TabPFN v2 is pre-trained exclusively on small- and medium-scale synthetic datasets, creating a mismatch when applied to large-scale datasets. Second, the complexity of transformers in TabPFN v2 scales with both $N$ (the number of instances) and $d$ (the number of features), meaning that increasing either value significantly can lead to a substantial rise in computational complexity, which in turn may reduce the model’s effectiveness.
In summary, while TabPFN v2 demonstrates exceptional performance on small- to medium-scale datasets, scaling it up to handle larger datasets, higher-dimensional feature spaces, and multi-class tasks exceeding 10 classes is a crucial and promising direction for future research.
5 Heterogeneity Mechanisms Analysis
As described in section 3, TabPFN v2 handles data heterogeneity by embedding a $d$-dimensional input vector into a $k$-dimensional space with size $d\times k$. This enables TabPFN v2 to adapt to diverse downstream datasets with varying feature dimensions seamlessly. In this section, we analyze this mechanism and revisit it as a variant of the token-based deep tabular method. Further analyses of the key components in TabPFN v2 are provided in Appendix D.
5.1 Analysis of the Heterogeneity Handling Strategy
Traditional deep tabular models often struggle with dataset heterogeneity due to their reliance on fixed feature semantics. For example, token-based approaches like AutoInt (Song et al., 2019) and FT-Transformer (FT-T) (Gorishniy et al., 2021) learn dataset- and attribute-specific feature tokens during training, which become ineffective when applied to new datasets with different feature meanings. These methods typically generate $d$ attribute-specific tokens $[{\bm{r}}_{1},\ldots,{\bm{r}}_{d}]\in{\mathbb{R}}^{k\times d}$, one for each dimension, and transform an instance ${\bm{x}}_{i}$ into $[{x_{i}^{1}}\cdot{\bm{r}}_{1},\ldots,{x_{i}^{d}}\cdot{\bm{r}}_{d}]\in{\mathbb{R}}^{k\times d}$, where $x_{i}^{j}$ is the $j$-th element of ${\bm{x}}_{i}$. However, when transitioning to a new dataset with $d^{\prime}$ features, these learned tokens cannot be reused, requiring feature tokens to be learned anew. Previous methods address this by relying on semantic descriptions of attributes (Yan et al., 2024) or assuming overlap between feature sets to construct a mapping (Onishi et al., 2023; Zhou et al., 2023).
In contrast, TabPFN v2 eliminates this dependency through randomized feature tokens. Specifically, TabPFN v2 applies a shared $k$-dimensional vector ${\bm{u}}\in\mathbb{R}^{k}$ to lift raw attribute values into a $k$-dimensional space. To differentiate attributes, TabPFN v2 adds a random $k$-dimensional perturbation for each attribute, consistent within a dataset but distinct across datasets. Formally, the randomized tokens are denoted as:
${\bm{R}}={\bm{W}}{\bm{P}}\in{\mathbb{R}}^{k\times d},$ (1)
where ${\bm{P}}\in{\mathbb{R}}^{k^{\prime}\times d}$ is randomly generated, and ${\bm{W}}\in{\mathbb{R}}^{k\times k^{\prime}}$ is a learnable pre-condition projection matrix that maps the random values to the required dimensionality ($k^{\prime}<k$).The $j$-th column of ${\bm{R}}$, ${\bm{r}}_{j}\in{\mathbb{R}}^{k}$, serves as the specific token for the $j$-th attribute and is combined with the shared vector ${\bm{u}}$. Then, an instance ${\bm{x}}_{i}$ is transformed through
$[{x_{i}^{1}}\cdot({\bm{u}}+{\bm{r}}_{1}),\ldots,{x_{i}^{d}}\cdot({\bm{u}}+{\bm{r}}_{d}),\tilde{{\bm{y}}}_{i}]\in{\mathbb{R}}^{k\times(d+1)}\;.$ (2)
This randomized tokenization ensures that distinct attributes—irrespective of their semantic meanings—are mapped to unique, nearly orthogonal directions in the high-dimensional space. This is akin to one-hot encoding but operates in a continuous over-parameterized regime, where random vectors naturally approximate orthogonality.
Through its transformer layers, TabPFN v2 dynamically combines these basis vectors during inference, weighting their contributions based on the input context. This process is analogous to inverse PCA: while PCA decomposes data into orthogonal components, TabPFN v2 reconstructs heterogeneous features by mixing randomized bases scaled by their attribute values. This approach avoids the need for pre-defined feature semantics, while the shared vector ${\bm{u}}$ introduces a stable inductive bias that supports cross-dataset generalization. Pre-training on diverse and abundant synthetic datasets further equips TabPFN v2 with the capability to adapt these bases across diverse tasks, unifying disparate tabular datasets into a coherent latent framework.
This randomized token strategy effectively removes the need for dataset- and attribute-specific feature token learning, addressing a significant limitation of prior methods and improving transferability to diverse downstream datasets.
Figure 4: The change of ranks between FT-T, MNCA, and the variants with randomized tokens over 45 tabular datasets. (The rank is calculated together with the 26 compared baselines.) Values in the parentheses denote the average rank.
5.2 Effectiveness of Randomized Feature Tokens
To further evaluate the effectiveness of randomized tokens, we integrate this strategy with two representative methods: FT-Transformer (FT-T) by Gorishniy et al. (2021) and ModernNCA (MNCA) by Ye et al. (2025).
FT-T with Randomized Tokens (FT-T). FT-T is a token-based method that learns specific tokens $\{{\bm{r}}_{1},\ldots,{\bm{r}}_{d}\}\in{\mathbb{R}}^{d\times k}$ for each attribute. After transforming an instance into a set of tokens, a transformer processes these tokens together with a learned class token. The output token corresponding to the class token is then fed into a linear classifier.
We replace FT-T’s learned tokens with the randomized tokens used in TabPFN v2, retrain FT-T (denoted as FT-T), and evaluate the resulting performance.
(a) Raw Feature
(b) Vanilla Extraction
(c) Layer 3
(d) Layer 6
(e) Layer 9
(f) Layer 12
Figure 5: Visualization of extracted embeddings for four datasets: churn (first row, two classes), bank (second row, two classes), KDD (third row, two classes), and website_phishing (fourth row, three classes). We use crosses to denote training examples and circles to denote test examples. (a) shows the raw features, while (b) presents the embeddings extracted using the vanilla strategy. (c)-(f) depict the embeddings obtained using our proposed methods across different layers. The accuracy value is calculated by training a linear model (logistic regression) over the extracted embeddings on the training set and predicting on the test set.
MNCA with Randomized Tokens (MNCA). ModernNCA is a neighborhood-based method that predicts an instance based on its similarity to training instances in a latent space. While ModernNCA typically operates on raw features, we adapt it into a token-based method by applying ModernNCA to the token output from FT-T before the final linear layer, incorporating the randomized tokens as above.
We evaluate these FT-T and MNCA variants over the tiny-bechmark2 from Ye et al. (2024b), which contains 45 representative tabular datasets. The results in Figure 4 demonstrate how the performance rank changes when replacing the specific tokens with the randomized tokens.
Remark. Our experiments confirm that feature tokens do not significantly rely on semantic meanings. Using randomized tokens only slightly increases the average ranks as long as they are distinct across attributes. This highlights the robustness and universality of the randomized token strategy, demonstrating its effectiveness in handling heterogeneous datasets and its potential to improve generalization across diverse tasks once pre-trained.
6 TabPFN v2 as a Feature Encoder
To better understand TabPFN v2, we examine whether its in-context learning capability also produces separable and meaningful feature representations.
6.1 Challenges in Embedding Extraction
As described in section 3, the output tokens from the multiple transformer layers in TabPFN v2 correspond one-to-one with the input tokens, forming a tensor of size $(N+1)\times(d+1)\times k$. The final token, corresponding to the (dummy) label embedding $\tilde{{\bm{y}}}^{*}$ of the test instance, is passed through an MLP block to produce the output. An intuitive idea is to treat the output tokens associated with the training label embeddings $\{{\bm{y}}_{i}\}_{i=1}^{N}$—prior to the MLP block—as the extracted embeddings for the training data.
However, this straightforward method has one fundamental limitation due to the distinct roles of labeled training and unlabeled test data in TabPFN v2’s in-context learning process. Specifically, the label embeddings for the training instances are derived from their true labels, while those for the test instances rely on dummy labels. This discrepancy results in output embeddings that are inherently non-comparable between the training and test instances.
6.2 Leave-one-fold-out Feature Extraction
To address this challenge, we propose a leave-one-fold-out strategy that enables the extraction of comparable embeddings for training and test data. Within the TabPFN v2 framework, we define the support set ${\mathcal{S}}$ as the examples with true labels and the query set ${\mathcal{Q}}$ as the examples with dummy labels. Under the standard configuration, ${\mathcal{S}}$ corresponds to the labeled training set, while ${\mathcal{Q}}$ corresponds to the unlabeled test instances.
A key trade-off arises: ${\mathcal{S}}$ needs to include as much labeled data as possible to propagate knowledge from ${\mathcal{S}}$ to ${\mathcal{Q}}$ effectively in the in-context learning process. However, to ensure that embeddings for training and test instances are comparable, training instances must also be included in ${\mathcal{Q}}$ and paired with dummy label embeddings.
To reconcile these requirements, we split the training set into multiple folds (e.g., 10 folds). One fold is used as ${\mathcal{Q}}$ for embedding extraction, while the remaining folds constitute ${\mathcal{S}}$ with true training labels. This ensures that ${\mathcal{S}}$ retains sufficient label information while allowing embeddings to be extracted for the training data in ${\mathcal{Q}}$. An extreme version of this strategy operates in a leave-one-out fashion, with only one training instance placed in ${\mathcal{Q}}$ at a time.
Results in Figure 5 (c)-(f) demonstrate that embeddings extracted using this strategy (with 10 folds) more effectively reveal dataset properties. We observe that TabPFN v2 simplifies tabular data distributions and transforms datasets into nearly separable embedding spaces, particularly in the embeddings after intermediate transformer layers.
Table 2: Average rank (lower is better) of TabPFN v2 and a linear classifier trained on the extracted embeddings across 29 classification datasets. “Combined” refers to an approach where embeddings from up to three layers (from the 12 available layers) are selected based on the validation set performance and concatenated.
$\downarrow$ TabPFN v2 Vanilla Layer 6 Layer 9 Layer 12 Combined
Rank 3.12 3.43 5.03 4.72 2.45 2.24
6.3 Validation of Embedding Quality
To validate the quality of the extracted embeddings, we train a logistic regression on top of the extracted embeddings. Specifically, the classifier is trained on embeddings derived from the training set and evaluated on test set embeddings. The average rank across 29 classification datasets from the tiny benchmark2 in Ye et al. (2024b) is reported in Table 2.
Remarkably, training a linear classifier on these extracted embeddings yields performance comparable to TabPFN v2’s in-context learner. Embeddings from intermediate layers (or their selected concatenations) sometimes achieve even better results. These findings highlight TabPFN v2’s potential as a robust feature encoder, offering valuable insights into its architecture and paving the way for broader applications in tabular data analysis and representation learning.
7 Extensions of TabPFN v2
We aim to expand TabPFN v2’s applicability beyond the boundaries outlined in Hollmann et al. (2025), which restricts its use to datasets with no more than 10,000 samples, 500 dimensions, and 10 classes. To address these limitations without requiring extensive retraining, we propose post-hoc divide-and-conquer strategies inspired by Chain-of-Thought (CoT) prompting in large-language models (Wei et al., 2022). We reformulate challenging tasks into multiple simpler subtasks that TabPFN v2 can effectively handle.
Figure 6: Mean accuracy on 18 high-dimensional datasets. “*” denote our extension of TabPFN v2, and “-PCA” is another variant using PCA to reduce dimensions.
7.1 High Dimension Datasets
High-dimensional datasets (Jiang et al., 2024) present a unique challenge due to the quadratic complexity of TabPFN v2 with respect to the number of dimensions. To mitigate this, we propose subsampling the feature space into smaller subsets, processing each subset independently, and combining the predictions in an ensemble (bagging) fashion, similar to random forests (Breiman, 2001).
In detail, we iteratively sample $m$ subsets (we set $m=4$ in experiments), each containing $d^{\prime}<d$ randomly selected attributes. For each subset, we leverage the TabPFN v2’s ability to handle lower-dimensional data to obtain predictions. The final result aggregates these outputs using averaging (for regression) or majority voting (for classification).
Figure 6 summarizes the results 18 high-dimensional classification datasets. TabPFN v2 with the proposed divide-and-conquer and then ensemble strategy significantly increases the mean accuracy (the second-highest) against the vanilla TabPFN v2, effectively extending TabPFN v2’s scalability to datasets with $d\geq 2000$. Another variant that utilizes PCA to reduce the dimensionality together with bagging also resolves the dimensionality issue to some extent.
Figure 7: Mean accuracy on 12 datasets with more than 10 classes. “PFNv2” denotes our extension, while “PFNv2-ECOC” denotes the multi-class ECOC strategy implemented by Hollmann et al. (2025) in their February 2025 code update.
7.2 Multi-Class Problems with More Than 10 Classes
To extend TabPFN v2 to tasks with more than 10 categories, we propose a decimal encoding approach that decomposes multi-class problems into multiple 10-class subproblems, ensuring compatibility with TabPFN v2’s constraints.
For a task with $C>10$ classes, we encode each label $y\in[C]$ as a $t$-digit decimal representation, where $t=\lceil\log_{10}C\rceil$. For each digit position $j\in\{1,\ldots,t\}$, we train a separate TabPFN v2 model $f_{j}$ to predict the $j$-th digit. For example, in a 15-class problem, we decompose it into two subtasks: the first predicts the tens digit (with classes ${0,1}$), while the second predicts the ones digit (with classes ${0,\dots,9}$). During inference, the predicted digits are reconstructed to obtain the final class label. To improve robustness, we permute the class order $\sqrt{C}$ times, leveraging the efficiency of TabPFN v2 to predict across all permutations. The final prediction is obtained by ensembling these results. As shown in Figure 7, this approach achieves the second-best mean accuracy on 12 datasets with more than 10 classes while preserving computational efficiency.
Figure 8: Average rank on 18 large-scale datasets. “*” indicates our extension. “-B” refers to the variant that randomly subsamples 10,000 training examples four times and aggregates their predictions. “-K” denotes the variant that selects a representative subset of 10,000 training examples based on proximity to prototypes obtained via KMeans. All variants improve TabPFN v2.
7.3 Large-Scale Datasets
For large-scale datasets, we randomly sample 10,000 training examples from the full training set as the support set and treat the remaining training examples and test instances as the query set. We extract their embeddings to form a new tabular dataset, on which a logistic regression classifier is trained to make predictions on the test set embeddings. This process is repeated four times, and the final predictions are aggregated, denoted as TabPFN v2-SQ.
We also investigate integrating TabPFN v2 with decision trees to handle large-scale tasks. Specifically, we first train a shallow decision tree, setting its maximum depth to 5 layers for datasets with fewer than 100,000 samples, and to 10 layers for larger datasets. The decision tree partitions the training set into smaller, more manageable subsets. Then, we apply TabPFN v2 to each leaf node, treating the examples in each node as a separate training set for TabPFN v2. During inference, a test instance is first passed through the shallow decision tree to a leaf node and then predicted by the corresponding TabPFN v2 model. We denote this extension as TabPFN v2-DT. We note that a similar strategy was mentioned in Hollmann et al. (2025) but for a drastically difference purpose. Specifically, Hollmann et al. (2025) use decision trees to handle within-dataset heterogeneous for some highly challenging tabular datasets, while we use decision trees for scalability.
Figure 8 shows the average rank results, including TabPFN v2-SQ and TabPFN v2-DT alongside variants using bagging and KMeans-based sampling. We observe that all variants improve upon the vanilla TabPFN v2 on large-scale datasets, with TabPFN v2-DT and TabPFN v2-SQ achieving the most significant improvement.
8 Conclusion
Our analysis highlights TabPFN v2’s potential as a pre-trained foundation model for tabular data. By leveraging randomized feature tokens, it effectively handles dataset heterogeneity and excels in in-context learning, outperforming other models on small- to medium-scale datasets. Additionally, its ability to create nearly separable embeddings suggests its potential role as a feature encoder. To address TabPFN v2’s limitations on high-dimensional, large-scale, and many-category datasets, we propose post-hoc divide-and-conquer strategies that improve scalability, expanding applicability while preserving its core strengths.
Appendix A Additional Related Work
Learning with Tabular Data. Tabular data is prevalent across diverse fields, including healthcare, finance, and education (Kovalerchuk & Vityaev, 2005; Hyland et al., 2020; Romero & Ventura, 2010; Amatriain et al., 2010). Tree-based models, such as XGBoost (Chen & Guestrin, 2016), LightGBM (Ke et al., 2017), and CatBoost (Prokhorenkova et al., 2018), have long dominated this domain. However, recent advances in deep neural networks (DNNs) have demonstrated strong potential for tabular data (Borisov et al., 2024). Popular architectures like multi-layer perceptrons (Gorishniy et al., 2021; Kadra et al., 2021) and Transformers (Huang et al., 2020) have been adapted to tabular tasks, alongside custom architectures designed specifically for tabular data (Klambauer et al., 2017; Wang et al., 2021).
Deep tabular methods can be broadly categorized into two types. The first type directly processes raw features (Holzmüller et al., 2024; Gorishniy et al., 2024; Ye et al., 2025), sometimes incorporating feature-specific encoding strategies (Gorishniy et al., 2022). The second type tokenizes features, transforming an example into a set of tokens (Song et al., 2019; Huang et al., 2020; Rubachev et al., 2022). Comprehensive benchmarks have been developed to evaluate these methods across diverse datasets (Grinsztajn et al., 2022; McElfresh et al., 2023; Ye et al., 2024b; Rubachev et al., 2025), highlighting the strengths and weaknesses of deep tabular models in various scenarios.
Appendix B Evaluation Details
Please refer Ye et al. (2024b) for details of the 300 small to medium datasets. For high-dimensional datasets, we selected 18 datasets with more than 2000 features from the scikit-feature repository. Detailed statistics of high-dimensional datasets and large-scale datasets are reported in Table 3 and Table 4.
We follow Ye et al. (2024b) and use different colors to represent various categories of methods in the result figures, ensuring clarity and easy comparison. In Figure 2 and Figure 3, we compare the following methods:
  • Classical Methods ( ): The classical methods include Dummy, Logistic Regression (LR), K-Nearest Neighbors (KNN), Support Vector Machines (SVM), Naive Bayes, Linear Regression, and DNNR (Nader et al., 2022), which serve as basic baselines for classification and regression tasks.
  • Tree-based Methods ( ): Tree-based methods such as Random Forest (Breiman, 2001), XGBoost (Chen & Guestrin, 2016), LightGBM (Ke et al., 2017), and CatBoost (Prokhorenkova et al., 2018) are known for their high performance on tabular data.
  • MLP variants ( ): MLP variants, including vanilla MLP, MLP-PLR, Self-Normalizing Neural Networks (Klambauer et al., 2017), and Residual Network (Gorishniy et al., 2021), enhance the flexibility and generalization of traditional MLP architectures through advanced regularization and residual connections.
  • Special Architectures ( ): Methods with specially designed architectures, such as DCNv2 (Wang et al., 2021), DANets (Chen et al., 2022), and TabCaps (Chen et al., 2023), focus on improving feature interaction and abstraction to capture complex relationships in tabular data.
  • Token-based Methods ( ): Token-based methods like AutoInt (Song et al., 2019), TabTransformer (Huang et al., 2020), FT-Transformer (Gorishniy et al., 2021), and ExcelFormer (Chen et al., 2024) represent features as tokens, enabling models to capture higher-order interactions through attention mechanisms.
  • Regularization-based Methods ( ): Regularization-based methods, including TANGOS (Jeffares et al., 2023), SwitchTab (Wu et al., 2024), and PTaRL (Ye et al., 2024a), aim to improve model generalization by incorporating regularization techniques during training to enhance the robustness of predictions.
  • Tree-mimic Methods ( ): Tree-mimic methods, such as NODE (Popov et al., 2020), GrowNet (Badirli et al., 2020), and TabNet (Arik & Pfister, 2021), combine the interpretability of decision trees with the power of deep learning, employing attention mechanisms to select important features.
  • Context-based Methods ( ): Context-based methods like TabR (Gorishniy et al., 2024) and ModernNCA (Ye et al., 2025) leverage contextual information from the training data to improve predictions by utilizing neighborhood-based and in-context learning strategies.
In addition to the aforementioned methods, for other experimental results, we will demonstrate the performance of TabPFN v2 and its variants, which are represented by emerald teal ( ), ensuring that their experimental effects are clearly distinguished from the other methods.
Table 3: Dataset Information for High-Dimensional Data Experiments: A collection of 18 datasets with varying numbers of instances, features, and classes used in our high-dimensional experiments.
Dataset #Instances #Features #Classes Dataset #Instances #Features #Classes
BASEHOCK 1993 4862 2 lung 203 3312 5
PCMAC 1943 3289 2 warpPIE10P 210 2420 10
RELATHE 1427 4322 2 orlraws10P 100 10304 10
ALLAML 72 7129 2 Prostate_GE 102 5966 2
CLL_SUB_111 111 11340 3 SMK_CAN_187 187 19993 2
colon 62 2000 2 warpAR10P 130 2400 10
GLI_85 85 22283 2 arcene 200 10000 2
GLIOMA 50 4434 4 gisette 7000 5000 2
leukemia 72 7070 2 TOX_171 171 5748 4
Table 4: Dataset Information for Large-scale Data Experiments.
Dataset #Instances #Features #Classes Dataset #Instances #Features #Classes
BNG(credit-a) 1,000,000 15 2 CDC_Indicators 253,680 21 2
Higgs 1,000,000 28 2 Smoking_signal 991,346 23 2
nomao 34,465 118 2 sf-police-incidents 2,215,023 8 2
Data_Crowdfunding 671,025 11 4 Fashion-MNIST 70,000 784 10
covertype 581,012 54 7 jannis 83,733 54 4
poker-hand 1,025,009 10 10 volkert 58,310 180 10
Airlines_DepDelay 10,000,000 9 - Wave_Energy_Farm 36,043 99 -
UJIndoorLoc 21,048 520 - blogfeedback 60,021 276 -
microsoft 1,200,192 136 - yahoo 709,877 699 -
Appendix C Detailed Results
We list the detailed results of TabPFN v2 and our extensions on various benchmarks.
  • We present the main results of TabPFN v2 on 300 datasets in LABEL:tab:main_results. The table includes accuracy for classification tasks and RMSE (Root Mean Squared Error) for regression tasks, along with the corresponding mean and standard deviation for each dataset. Notably, we excluded 27 datasets from these results in LABEL:tab:main_results, as they were used by TabPFN v2 to select the best checkpoint. These excluded datasets, which are not shown in Figure 2 and Figure 3, include:
    • (1)
      ada_prior, allbp, baseball, delta_ailerons, eye_movements, eye_movements_bin, GAMETES_Epistasis_2-Way_20atts_0.1H_EDM-1_1, hill-valley, JapaneseVowels, jungle_chess_2pcs_raw_endgame_complete, led24, longitudinal-survey, page-blocks, ringnorm, rl, thyroid-ann, waveform-5000,
    • (2)
      debutanizer, delta_elevators, mauna-loa-atmospheric, puma32H, stock_fardamento02, treasury, weather_izmir, wind.
  • In Table 6, we showcase the performance of various models on 18 high-dimensional datasets. The results display the mean accuracy of different models, including ModernNCA (MNCA), MLP, KNN, RealMLP, XGBoost (XGB), Random Forest (RForest), Logistic Regression (LogReg), and TabPFN v2 (PFN-v2), along with variants like TabPFN v2-pca and TabPFN v2*. This highlights the ability of these models to handle high-dimensional data with many features.
  • We demonstrate the performance of various models on 12 multi-class classification tasks with more than 10 classes in Table 7. The table provides the mean accuracy of models like KNN, TabPFN-v2*, XGBoost (XGB), CatBoost (CatB), Random Forest (RForest), ModernNCA (MNCA), MLP, Logistic Regression (LogReg), and RealMLP, showcasing how they perform on multi-class tasks with a larger number of classes. Additionally, we compare PFN-v2-ECOC, a multi-class classification solution provided by Hollmann et al. (2025). This method extends TabPFN-v2 by leveraging Error-Correcting Output Codes (ECOC) to enhance multi-class classification performance [^2] https://github.com/PriorLabs/tabpfn-community/blob/main/src/tabpfn_extensions/many_class/many_class_classifier.py.
  • In Table 8, we compare the performance of various models on 18 large-scale datasets. The results show the mean accuracy or RMSE for MLP, Logistic/Linear Regression (LR), KNN, XGBoost (XGB), Random Forest (RForest), CatBoost (CatB), ModernNCA (MNCA), RealMLP, and different versions of TabPFN v2 (PFNv2, PFNv2 with K-means, PFNv2 with Bagging, and PFNv2*). This illustrates the models’ performance on large-scale datasets.
  • We present a comparison of FT-T (Gorishniy et al., 2021) and MNCA (Ye et al., 2025) with and without randomized feature tokens on tiny benchmark datasets from Ye et al. (2024b) in Table 9. The first two columns represent FT-T and MNCA with randomized feature tokens, while the last two show the original FT-T and MNCA methods. The table demonstrates the effect of randomized feature tokens on model performance, with accuracy for classification tasks and RMSE for regression tasks. FT-T and MNCA implementations, including their hyperparameter search space, follow the approach described in Liu et al. (2024), and the results were obtained after 100 rounds of hyperparameter tuning.
  • In Table 10, we show the performance of TabPFN v2 and the extracted feature embeddings across 29 classification datasets. The table includes average classification accuracy for each dataset when using feature embeddings from different transformer layers (Layer 6, Layer 9, Layer 12), as well as a combined approach where embeddings from multiple layers are concatenated. The “selected layers” column indicates the layers chosen based on validation set performance, offering insights into how different layers contribute to overall model performance. In addition to evaluating the performance of TabPFN v2 and the extracted feature embeddings, we also compared the results with embeddings obtained using the vanilla strategy (Vanilla).
Table 5: Main results of TabPFN v2 on 300 datasets, including accuracy (for classification tasks) and RMSE (for regression tasks), along with the corresponding mean and standard deviation for each dataset. Among the 300 datasets, 200 are classification datasets, and 100 are regression datasets. The results demonstrate the effectiveness of TabPFN v2 across both classification and regression tasks.
Dataset Mean + Std Dataset Mean + Std
ASP-POTASSCO-classification 43.50 $\pm$ 1.27 Amazon_employee_access 94.22 $\pm$ 0.04
BLE_RSSI_localization 73.37 $\pm$ 0.15 BNG(breast-w) 98.56 $\pm$ 0.07
BNG(cmc) 57.69 $\pm$ 0.17 BNG(tic-tac-toe) 79.42 $\pm$ 0.26
Bank_Customer_Churn_Dataset 87.53 $\pm$ 0.12 Basketball_c 70.65 $\pm$ 0.47
California-Housing-Classification 91.47 $\pm$ 0.17 Cardiovascular-Disease-dataset 72.92 $\pm$ 0.13
Click_prediction_small 83.29 $\pm$ 0.03 Contaminant-10.0GHz 94.42 $\pm$ 0.36
Contaminant-10.5GHz 95.17 $\pm$ 0.32 Contaminant-11.0GHz 93.93 $\pm$ 0.50
Contaminant-9.0GHz 93.01 $\pm$ 0.47 Contaminant-9.5GHz 93.21 $\pm$ 0.50
Credit_c 69.98 $\pm$ 0.15 Customer_Personality_Analysis 90.03 $\pm$ 0.21
Diabetic_Retinopathy_Debrecen 72.81 $\pm$ 1.07 E-CommereShippingData 67.54 $\pm$ 0.21
Employee 84.80 $\pm$ 0.30 FICO-HELOC-cleaned 75.35 $\pm$ 0.21
FOREX_audcad-day-High 74.51 $\pm$ 0.51 FOREX_audcad-hour-High 71.01 $\pm$ 0.20
FOREX_audchf-day-High 76.66 $\pm$ 0.45 FOREX_audjpy-day-High 78.00 $\pm$ 0.28
FOREX_audjpy-hour-High 71.41 $\pm$ 0.32 FOREX_audsgd-hour-High 69.81 $\pm$ 0.39
FOREX_audusd-hour-High 69.57 $\pm$ 0.48 FOREX_cadjpy-day-High 71.68 $\pm$ 0.53
FOREX_cadjpy-hour-High 70.55 $\pm$ 0.40 Firm-Teacher-Direction 84.42 $\pm$ 0.47
Fitness_Club_c 79.67 $\pm$ 0.24 GAMETES_Epistasis 68.75 $\pm$ 0.82
GAMETES_Heterogeneity 65.90 $\pm$ 1.84 Gender_Gap_in_Spanish_WP 60.58 $\pm$ 0.24
GesturePhaseSegmentationProcessed 71.36 $\pm$ 1.15 HR_Analytics 80.02 $\pm$ 0.13
Heart-Disease-Dataset 91.23 $\pm$ 0.54 INNHotelsGroup 87.98 $\pm$ 0.23
Indian_pines 96.41 $\pm$ 0.23 Insurance 75.75 $\pm$ 0.00
Intersectional-Bias-Assessment 94.73 $\pm$ 0.13 Is-this-a-good-customer 88.41 $\pm$ 0.00
JapaneseVowels 99.68 $\pm$ 0.08 KDD 80.14 $\pm$ 0.46
KDDCup09_upselling 81.06 $\pm$ 0.26 Long 99.88 $\pm$ 0.00
MIC 90.20 $\pm$ 0.56 MagicTelescope 88.13 $\pm$ 0.21
Marketing_Campaign 88.11 $\pm$ 0.41 Mobile_Price_Classification 97.10 $\pm$ 0.29
Nutrition_Health_Survey 83.45 $\pm$ 0.22 Performance-Prediction 73.23 $\pm$ 0.61
PhishingWebsites 96.74 $\pm$ 0.13 PieChart3 87.31 $\pm$ 0.28
Pima_Indians_Diabetes_Database 75.93 $\pm$ 0.66 PizzaCutter3 88.20 $\pm$ 0.45
Pumpkin_Seeds 87.93 $\pm$ 0.21 QSAR_biodegradation 88.50 $\pm$ 0.50
Rain_in_Australia 83.88 $\pm$ 0.11 SDSS17 97.33 $\pm$ 0.06
Shipping 68.73 $\pm$ 0.40 Telecom_Churn_Dataset 95.18 $\pm$ 0.50
UJI_Pen_Characters 45.71 $\pm$ 2.16 VulNoneVul 98.95 $\pm$ 0.00
Water_Quality_and_Potability 65.49 $\pm$ 0.50 Waterstress 71.37 $\pm$ 0.96
Wilt 99.28 $\pm$ 0.06 abalone 63.58 $\pm$ 0.38
accelerometer 73.96 $\pm$ 1.32 ada 85.40 $\pm$ 0.25
ada_agnostic 83.99 $\pm$ 0.34 ada_prior 85.32 $\pm$ 0.19
adult 85.93 $\pm$ 0.12 airlines_2000 62.28 $\pm$ 0.48
allbp 97.85 $\pm$ 0.19 allrep 98.65 $\pm$ 0.12
analcatdata_authorship 99.72 $\pm$ 0.29 artificial-characters 73.90 $\pm$ 0.99
autoUniv-au4-2500 69.81 $\pm$ 1.00 autoUniv-au7-1100 41.18 $\pm$ 1.58
bank 90.86 $\pm$ 0.19 banknote_authentication 55.64 $\pm$ 0.18
baseball 93.81 $\pm$ 0.40 car-evaluation 98.29 $\pm$ 0.22
churn 96.33 $\pm$ 0.28 cmc 59.59 $\pm$ 0.49
company_bankruptcy_prediction 97.33 $\pm$ 0.07 compass 71.05 $\pm$ 0.29
connect-4 76.78 $\pm$ 0.35 contraceptive_method_choice 62.10 $\pm$ 0.37
credit 78.10 $\pm$ 0.11 credit-g 79.50 $\pm$ 0.81
customer_satisfaction_in_airline 94.79 $\pm$ 0.11 dabetes_130-us_hospitals 63.08 $\pm$ 0.07
default_of_credit_card_clients 82.63 $\pm$ 0.08 delta_ailerons 95.47 $\pm$ 0.09
dis 99.07 $\pm$ 0.14 dna 97.25 $\pm$ 0.20
drug_consumption 40.32 $\pm$ 0.00 dry_bean_dataset 92.76 $\pm$ 0.10
eeg-eye-state 98.34 $\pm$ 0.12 electricity 86.57 $\pm$ 0.45
estimation_of_obesity_levels 98.66 $\pm$ 0.24 eucalyptus 72.88 $\pm$ 1.17
eye_movements 77.03 $\pm$ 1.68 eye_movements_bin 67.28 $\pm$ 2.60
first-order-theorem-proving 61.12 $\pm$ 0.70 gas-drift 99.47 $\pm$ 0.04
golf_play_dataset_extended 92.60 $\pm$ 0.44 helena 33.32 $\pm$ 0.21
heloc 72.75 $\pm$ 0.20 hill-valley 98.33 $\pm$ 0.52
house_16H 88.55 $\pm$ 0.18 htru 97.95 $\pm$ 0.06
ibm-employee-performance 100.0 $\pm$ 0.00 in_vehicle_coupon 73.20 $\pm$ 0.35
internet_firewall 92.85 $\pm$ 0.30 internet_usage 54.34 $\pm$ 2.64
jasmine 81.34 $\pm$ 0.42 jm1 81.32 $\pm$ 0.10
jungle_chess_2pcs_raw_endgame 85.97 $\pm$ 1.82 kc1 86.65 $\pm$ 0.34
kdd_ipums_la_97-small 88.50 $\pm$ 0.12 kr-vs-k 78.46 $\pm$ 1.01
kr-vs-kp 99.64 $\pm$ 0.15 kropt 77.96 $\pm$ 0.63
law-school-admission-bianry 100.0 $\pm$ 0.00 led24 73.29 $\pm$ 0.62
led7 73.99 $\pm$ 0.31 letter 97.57 $\pm$ 0.10
madeline 90.72 $\pm$ 0.48 mammography 98.71 $\pm$ 0.05
maternal_health_risk 83.28 $\pm$ 0.64 mfeat-factors 96.98 $\pm$ 0.28
mfeat-fourier 89.85 $\pm$ 0.86 mfeat-karhunen 96.42 $\pm$ 0.24
mfeat-morphological 76.63 $\pm$ 0.50 mfeat-pixel 96.10 $\pm$ 0.32
mfeat-zernike 84.10 $\pm$ 0.87 mice_protein_expression 100.0 $\pm$ 0.00
microaggregation2 62.80 $\pm$ 0.14 mobile_c36_oversampling 98.11 $\pm$ 0.08
mozilla4 93.58 $\pm$ 0.16 naticusdroid+android+permissions 96.41 $\pm$ 0.10
national-longitudinal-survey-binary 100.0 $\pm$ 0.00 okcupid_stem 74.47 $\pm$ 0.12
one-hundred-plants-margin 88.56 $\pm$ 0.74 one-hundred-plants-shape 79.52 $\pm$ 0.72
one-hundred-plants-texture 90.94 $\pm$ 0.75 online_shoppers 90.65 $\pm$ 0.10
optdigits 98.59 $\pm$ 0.12 ozone-level-8hr 94.92 $\pm$ 0.25
ozone_level 97.86 $\pm$ 0.11 page-blocks 97.67 $\pm$ 0.10
pc1 93.51 $\pm$ 0.46 pc3 88.78 $\pm$ 0.27
pc4 90.87 $\pm$ 0.36 pendigits 99.56 $\pm$ 0.06
philippine 84.20 $\pm$ 1.24 phoneme 88.47 $\pm$ 0.35
pol 98.80 $\pm$ 0.09 predict_students_dropout 78.11 $\pm$ 0.38
rice_cammeo_and_osmancik 92.74 $\pm$ 0.23 ringnorm 98.00 $\pm$ 0.13
rl 86.04 $\pm$ 0.44 satimage 92.30 $\pm$ 0.29
segment 93.91 $\pm$ 0.19 seismic+bumps 93.40 $\pm$ 0.08
semeion 92.41 $\pm$ 0.94 shill-bidding 90.31 $\pm$ 0.18
shrutime 86.97 $\pm$ 0.11 shuttle 99.86 $\pm$ 0.04
spambase 94.85 $\pm$ 0.19 splice 96.61 $\pm$ 0.22
sports_articles 84.93 $\pm$ 0.40 statlog 72.13 $\pm$ 0.97
steel_plates_faults 84.68 $\pm$ 0.55 svmguide3 85.54 $\pm$ 0.54
sylvine 97.30 $\pm$ 0.27 taiwanese_bankruptcy_prediction 97.20 $\pm$ 0.07
telco-customer-churn 80.29 $\pm$ 0.28 texture 100.0 $\pm$ 0.00
W thyroid 99.48 $\pm$ 0.06 thyroid-ann 99.34 $\pm$ 0.08
thyroid-dis 68.75 $\pm$ 0.34 turiye_student_evaluation 51.74 $\pm$ 0.18
twonorm 97.94 $\pm$ 0.08 vehicle 84.31 $\pm$ 1.29
walking-activity 61.22 $\pm$ 0.22 wall-robot-navigation 99.44 $\pm$ 0.10
water_quality 90.12 $\pm$ 0.12 waveform-5000 86.29 $\pm$ 0.26
waveform_v1 86.59 $\pm$ 0.25 website_phishing 90.48 $\pm$ 0.48
wine 75.12 $\pm$ 0.72 wine-quality-red 58.35 $\pm$ 0.76
wine-quality-white 64.15 $\pm$ 0.69 yeast 60.18 $\pm$ 0.65
1000-Cameras-Dataset 607.71 $\pm$ 6.61 2dplanes 1.01 $\pm$ 0.00
RSSI_Estimation 0.00068 $\pm$ 0.00 RSSI_Estimation1 0.00092 $\pm$ 0.00
Abalone_reg 2.08 $\pm$ 0.00 Ailerons 0.00015 $\pm$ 0.00
Fiat 716.20 $\pm$ 4.05 BNG(echoMonths) 11.41 $\pm$ 0.03
BNG(lowbwt) 455.27 $\pm$ 0.78 BNG(mv) 4.63 $\pm$ 0.01
BNG(stock) 2.95 $\pm$ 0.02 Bias_correction_r 0.60 $\pm$ 0.01
Bias_correction_r_2 0.52 $\pm$ 0.01 Brazilian_houses_reproduced 0.01 $\pm$ 0.00
CPMP-2015-regression 478.02 $\pm$ 5.40 CPS1988 364.02 $\pm$ 0.24
CookbookReviews 1.52 $\pm$ 0.02 Data_Science_Salaries 60237.28 $\pm$ 102.97
Diamonds 533.30 $\pm$ 6.30 Facebook_Comment_Volume 23.16 $\pm$ 0.20
Food_Delivery_Time 7.55 $\pm$ 0.03 Goodreads-Computer-Books 0.43 $\pm$ 0.00
IEEE80211aa-GATS 0.02 $\pm$ 0.00 Job_Profitability 13.14 $\pm$ 0.02
bike_sharing_demand 68.41 $\pm$ 0.60 Laptop_Prices_Dataset 439.87 $\pm$ 3.10
Wave_Energy_Perth_100 15507.90 $\pm$ 104.31 Wave_Energy_Sydney_100 14737.67 $\pm$ 150.43
Wave_Energy_Sydney_49 4567.97 $\pm$ 64.02 MIP-2016-regression 20966.10 $\pm$ 454.90
MiamiHousing2016 83101.09 $\pm$ 507.30 Mobile_Phone_Market 714.87 $\pm$ 11.15
Moneyball 19.42 $\pm$ 0.08 NASA_PHM2008 40.24 $\pm$ 0.06
NHANES_age_prediction 15.47 $\pm$ 0.04 OnlineNewsPopularity 8606.54 $\pm$ 7.04
Parkinson_Sound_Record 14.58 $\pm$ 0.09 Parkinsons_Telemonitoring 0.60 $\pm$ 0.04
Physicochemical_r 3.45 $\pm$ 0.04 SAT11-HAND-runtime 1232.03 $\pm$ 58.01
Shop_Customer_Data 28.56 $\pm$ 0.01 Superconductivty 10.17 $\pm$ 0.07
Wine_Quality_red 0.65 $\pm$ 0.00 Wine_Quality_white 0.68 $\pm$ 0.00
airfoil_self_noise 1.16 $\pm$ 0.02 analcatdata_supreme 0.09 $\pm$ 0.00
archive2 342.64 $\pm$ 3.20 archive_r56_Portuguese 2.86 $\pm$ 0.02
auction_verification 1145.54 $\pm$ 146.94 avocado_sales 0.09 $\pm$ 0.00
bank32nh 0.08 $\pm$ 0.00 bank8FM 0.03 $\pm$ 0.00
boston 4.25 $\pm$ 0.19 chscase_foot 0.95 $\pm$ 0.00
colleges 0.14 $\pm$ 0.00 combined_cycle_power_plant 3.22 $\pm$ 0.05
communities_and_crime 0.13 $\pm$ 0.00 concrete_compressive_strength 4.63 $\pm$ 0.07
cpu_act 2.65 $\pm$ 0.03 cpu_small 3.06 $\pm$ 0.02
dataset_sales 4.04 $\pm$ 0.02 debutanizer 0.04 $\pm$ 0.00
delta_elevators 0.0014 $\pm$ 0.00 elevators 0.0019 $\pm$ 0.00
fifa 0.78 $\pm$ 0.00 fried 1.01 $\pm$ 0.00
garments_worker_productivity 0.13 $\pm$ 0.00 gas_turbine_emission 0.44 $\pm$ 0.00
healthcare_insurance_expenses 4716.87 $\pm$ 36.52 house_16H_reg 29631.75 $\pm$ 251.56
house_8L 28617.41 $\pm$ 202.41 house_prices_nominal 30676.02 $\pm$ 2455.48
house_sales_reduced 132655.03 $\pm$ 1847.33 houses 42559.98 $\pm$ 928.78
housing_price_prediction 1009361.62 $\pm$ 8758.05 kin8nm 0.08 $\pm$ 0.00
mauna-loa-atmospheric-co2 0.39 $\pm$ 0.01 mv 0.02 $\pm$ 0.00
pol_reg 3.84 $\pm$ 0.10 pole 3.21 $\pm$ 0.14
puma32H 0.01 $\pm$ 0.00 puma8NH 3.24 $\pm$ 0.00
qsar_aquatic_toxicity 1.05 $\pm$ 0.01 qsar_fish_toxicity 0.86 $\pm$ 0.01
satellite_image 0.65 $\pm$ 0.00 sensory 0.77 $\pm$ 0.01
socmob 19.53 $\pm$ 0.64 space_ga 0.09 $\pm$ 0.00
steel_industry_energy 0.37 $\pm$ 0.03 stock 0.65 $\pm$ 0.01
stock_fardamento02 17.57 $\pm$ 0.08 sulfur 0.03 $\pm$ 0.00
topo_2_1 0.03 $\pm$ 0.00 treasury 0.23 $\pm$ 0.00
us_crime 0.14 $\pm$ 0.00 volume 52.09 $\pm$ 0.34
weather_izmir 1.09 $\pm$ 0.01 wind 2.83 $\pm$ 0.00
wine+quality 0.72 $\pm$ 0.00 yprop_4_1 0.03 $\pm$ 0.00
Table 6: Performance of various models on 18 high-dimensional datasets. The results show the mean accuracy of different models, including ModernNCA (MNCA), MLP, KNN, RealMLP, XGBoost (XGB), Random Forest (RForest), Logistic Regression (LogReg), TabPFN v2 (PFN-v2), TabPFN v2 with PCA (v2-pca), TabPFN v2 with subsampling (v2*), ProtoGate (ProtoG), and CatBoost (CatB). The performance is evaluated on high-dimensional datasets, with the values representing mean accuracy for each model.
Dataset MNCA MLP KNN RealMLP XGB RForest LogReg PFN-v2 v2-pca v2* ProtoG CatB
CLL_SUB_111 62.90 72.46 57.39 70.43 73.04 70.14 73.91 70.14 57.68 73.17 65.51 71.59
BASEHOCK 96.31 97.01 71.88 97.46 95.29 96.73 96.99 69.09 97.41 97.47 96.32 95.87
Prostate_GE 81.27 86.98 80.00 87.94 89.52 87.94 91.43 95.24 88.57 94.60 84.13 94.92
PCMAC 88.21 88.53 66.48 90.15 91.64 92.20 87.15 92.70 90.76 86.86 88.21 92.01
GLI_85 81.57 85.49 76.47 89.80 82.35 83.92 90.59 80.39 86.27 92.55 81.96 80.78
RELATHE 88.18 90.54 75.03 90.23 87.11 87.30 90.49 86.36 87.65 86.81 89.92 90.35
SMK_CAN_187 63.51 66.84 69.47 69.82 66.49 70.70 72.11 71.05 71.75 73.28 70.71 71.40
warpPIE10P 98.41 99.05 92.38 100.0 94.92 98.57 100.0 100.0 100.0 99.84 97.79 98.89
leukemia 90.22 95.11 86.67 94.67 97.78 92.00 96.00 92.44 93.33 95.33 94.00 94.22
orlraws10P 97.67 98.33 92.00 99.00 84.33 99.00 99.00 92.00 99.33 98.33 92.67 99.00
GLIOMA 58.00 60.67 68.00 67.33 66.67 64.00 64.00 62.67 69.33 68.00 69.91 66.67
warpAR10P 83.08 85.64 53.08 97.44 81.28 87.18 97.69 90.77 95.38 92.05 90.04 87.44
TOX_171 76.00 88.19 70.86 90.48 78.10 78.67 90.29 80.95 82.48 86.86 85.52 83.05
lung 91.54 95.45 93.66 95.28 93.66 92.68 95.12 95.28 93.50 95.45 95.43 93.01
ALLAML 87.56 95.56 81.33 96.89 96.00 96.44 92.00 92.89 93.78 92.44 91.14 94.67
colon 78.46 78.97 76.92 83.08 74.87 82.56 86.15 81.54 78.46 80.00 78.46 77.95
gisette 97.21 97.57 95.04 97.86 97.55 96.82 97.51 97.35 97.26 97.37 97.18 97.78
arcene 81.67 85.50 84.50 81.00 75.00 86.83 88.00 83.67 88.33 90.67 85.33 85.00
Mean 82.86 87.11 77.29 88.83 84.76 86.87 89.36 85.25 87.29 88.95 86.37 87.48
Table 7: Performance of various models on 12 multi-class classification tasks with more than 10 classes. The results show the mean accuracy of different models, including KNN, PFN-v2*, PFN-v2-ECOC, XGB (XGBoost), CatBoost (CatB), Random Forest (RForest), ModernNCA (MNCA), Multi-layer Perceptron (MLP), Logistic Regression (LogReg), and RealMLP. The performance is evaluated on 12 multi-class datasets with more than 10 classes, with accuracy values presented for each model on the respective datasets.
Dataset KNN PFN-v2* PFN-v2-ECOC XGB CatB RForest MNCA MLP LogReg RealMLP
100-plants-texture 79.69 90.94 84.92 77.06 89.73 82.65 80.52 83.92 86.88 88.35
100-plants-margin 77.50 88.56 79.40 74.25 84.06 82.79 77.60 80.44 79.69 83.58
100-plants-shape 60.31 79.52 63.38 56.15 65.19 64.33 70.10 47.33 65.94 72.08
UJI_Pen_Characters 36.26 45.71 44.20 30.35 38.88 34.24 44.03 37.75 19.41 46.37
texture 98.45 100.0 100.0 98.55 99.13 96.76 99.68 99.40 99.64 99.95
letter 94.90 97.57 97.78 96.26 96.75 91.56 97.96 96.40 75.80 98.31
walking-activity 60.29 61.22 61.92 65.06 64.92 61.74 64.85 60.64 27.02 65.13
helena 28.94 33.31 19.20 32.42 37.90 33.91 36.58 37.91 33.40 38.55
internet_usage 30.17 54.34 50.86 51.08 37.90 33.91 52.09 43.00 37.73 52.23
kropt 71.22 77.96 77.11 86.95 79.26 71.77 78.27 64.45 28.08 92.03
kr-vs-k 70.78 78.46 76.29 87.26 74.81 71.60 76.83 65.03 28.03 91.85
ASP-POTASSCO 34.75 43.50 45.27 42.24 41.08 42.86 37.45 29.63 35.14 41.70
Mean 61.94 70.93 66.69 66.47 67.47 64.01 68.00 62.16 51.40 72.51
Table 8: Performance of various models on 18 large-scale datasets. The results show the mean accuracy/RMSE of different models, including MLP, Logistic Regression/Linear Regression (LR), KNN, XGBoost (XGB), Random Forest (RForest), CatBoost (CatB), ModernNCA (MNCA), RealMLP, and various versions of TabPFN v2: original TabPFN v2 (PFNv2), TabPFN v2 with K-means (PFNv2-K), TabPFN v2 with Bagging (PFNv2-B), and PFNv2* (TabPFNv2*).
Dataset MLP LR KNN XGB RForest CatB MNCA RealMLP PFNv2 PFNv2-K PFNv2-B PFNv2* PFNv2-DT
BNG(credit-a) 90.07 85.98 87.41 90.21 89.25 91.13 89.98 90.91 89.55 89.01 89.66 89.89 89.96
CDC_Indicators 86.79 86.55 86.39 86.76 86.60 86.78 86.76 86.76 86.65 86.68 86.69 86.74 86.22
Higgs 75.53 64.29 65.16 73.33 71.87 74.81 73.28 75.36 71.64 71.56 72.01 72.13 72.44
Smoking_signal 73.90 72.53 72.36 73.87 73.08 73.99 73.63 74.00 73.47 73.37 73.55 73.69 72.94
nomao 96.19 94.59 95.20 96.92 96.07 97.03 96.68 96.37 96.08 96.29 96.12 96.18 96.51
sf-police-incidents 87.84 87.84 85.87 87.68 87.84 87.87 - 87.84 87.84 87.84 87.84 87.84 87.84
Data_Crowdfunding 96.48 67.04 93.70 96.89 95.29 96.81 96.53 96.71 94.59 91.81 94.96 95.07 96.93
Fashion-MNIST 89.54 85.69 86.00 90.03 86.57 90.24 89.36 90.25 68.40 82.82 83.89 86.26 78.58
covertype 94.01 72.54 92.76 96.30 78.30 90.77 97.31 97.38 83.54 82.95 84.16 86.85 96.89
jannis 71.99 64.60 65.67 71.83 69.19 72.26 72.57 73.00 70.24 70.26 70.59 71.31 71.90
poker-hand 99.99 50.12 54.01 99.51 64.63 97.69 76.31 99.88 41.97 38.86 36.80 54.12 85.60
volkert 69.85 58.75 67.41 69.74 62.71 70.88 77.18 73.76 62.82 62.15 62.81 64.84 68.53
Airlines_DepDelay (×$\times 10^{1}$) 2.905 2.933 3.170 2.891 2.907 2.881 - 2.482 2.937 2.933 2.937 2.915 2.906
Wave_Energy_Farm (×$\times 10^{3}$) 8.199 13.19 32.29 6.917 7.294 7.173 6.148 59.05 7.214 8.375 7.063 10.506 6.489
UJIndoorLoc (×$\times 10^{0}$) 9.958 $\infty$ 9.004 10.47 23.19 9.139 5.990 65.34 66.49 7.825 7.435 9.538 7.024
blogfeedback (×$\times 10^{1}$ ) 2.387 $\infty$ 2.410 2.093 2.026 2.044 1.953 2.105 3.073 2.687 2.700 2.014 2.011
microsoft (×$\times 10^{-1}$ ) 7.577 7.782 8.284 7.514 7.566 7.453 7.573 5.077 7.735 7.981 7.720 7.612 7.952
yahoo (×$\times 10^{-1}$ ) 7.692 7.997 8.504 7.629 - 7.514 - 5.671 8.148 8.332 8.132 7.961 8.118
Table 9: Performance comparison of FT-T and MNCA with and without randomized feature tokens across tiny benchmark datasets of Ye et al. (2024b). The first two columns (FT-T and MNCA) show results using randomized tokens, while the last two columns (FT-T and MNCA) represent the original methods. All values are the mean of 15 runs with different random seeds. For classification tasks, the performance is measured by accuracy, while for regression tasks, the metric used is Root Mean Squared Error (RMSE). The table demonstrates the effect of randomized feature tokens on model performance in both types of tasks.
Dataset FT-T* MNCA* FT-T MNCA
BNG(breast-w) 98.62 98.66 98.60 98.59
BNG(cmc) 58.35 58.39 58.50 58.56
BNG(tic-tac-toe) 80.92 81.10 81.31 81.08
Cardiovascular-Disease-dataset 73.35 73.29 73.22 73.20
FOREX_audchf-day-High 52.90 72.93 51.79 75.44
FOREX_audsgd-hour-High 69.53 66.86 61.84 71.09
FOREX_cadjpy-hour-High 69.62 70.99 68.80 71.18
Gender_Gap_in_Spanish_WP 59.16 59.16 59.86 59.25
KDD 80.15 80.12 80.28 79.58
VulNoneVul 98.93 98.81 98.93 98.79
baseball 93.56 93.71 93.76 92.74
credit 75.83 75.39 75.88 72.98
dis 98.29 98.31 98.76 98.68
eye_movements_bin 59.01 61.31 59.45 90.88
jungle_chess_2pcs_raw_endgame_complete 97.11 99.67 97.61 99.50
law-school-admission-bianry 99.00 100.0 98.02 100.0
mfeat-fourier 86.28 85.78 85.23 86.48
online_shoppers 89.70 89.65 90.19 90.53
page-blocks 96.89 97.23 96.71 96.68
pc3 88.54 89.16 89.78 88.75
pendigits 99.20 99.39 99.35 99.40
rl 66.99 79.74 71.96 83.75
satimage 89.61 90.77 89.59 91.04
segment 91.49 91.40 91.37 92.77
sylvine 94.48 94.22 94.71 95.88
taiwanese_bankruptcy_prediction 96.90 96.68 96.74 96.74
waveform-5000 85.89 86.02 86.11 86.24
website_phishing 89.45 89.94 90.14 87.53
wine-quality-white 54.98 64.41 55.17 63.27
Ailerons 0.000157 0.000158 0.000155 0.000158
CookbookReviews 1.560657 1.491684 1.520183 1.486840
IEEE80211aa-GATS 0.030513 0.027608 0.029589 0.026295
Large-scale_Wave_Energy_Farm_Sydney_49 4504.285 4369.826 4369.902 47399.63
Superconductivty 10.53080 10.39726 10.64536 10.37837
archive2 388.8410 379.5506 390.6574 361.8616
bank8FM 0.028792 0.029047 0.028635 0.028826
communities_and_crime 0.143486 0.139045 0.136859 0.138984
fried 1.013058 1.013431 1.008568 1.010952
healthcare_insurance_expenses 4935.203 4743.384 4615.580 4698.968
house_16H_reg 31647.43 31333.32 30671.63 31017.80
kin8nm 0.068684 0.070110 0.067355 0.069678
mv 0.027143 0.019089 0.030948 0.027385
Table 10: Performance of TabPFN v2 and the extracted feature embeddings across 29 classification datasets. The table shows the average classification accuracy for each dataset when using different layers (Layer 6, Layer 9, Layer 12) of the transformer as feature embeddings, as well as the “combined” approach, where embeddings from up to three selected layers are concatenated. The “selected layers” column indicates the specific layers chosen for each dataset based on validation set performance. “Vanilla” refers to the embeddings extracted using the vanilla strategy, which utilizes only the 12th layer of the transformer.
PFN-v2 Vanilla layer-6 layer-9 layer-12 combined selected layers
FOREX_audchf-day-High 77.38 72.48 68.39 73.57 74.11 77.11 (5, 9, 11)
taiwanese_bankruptcy_prediction 96.99 96.70 97.14 96.77 97.07 97.14 (6)
rl 85.51 85.41 66.90 69.52 86.72 87.53 (11, 12)
pc3 89.46 89.78 90.10 88.82 88.82 88.82 (8)
eye_movements_bin 61.83 62.75 59.72 59.40 62.16 62.16 (6, 9, 12)
BNG(breast-w) 98.43 98.51 98.34 98.46 98.67 98.51 (6, 9)
FOREX_cadjpy-hour-High 69.53 70.77 62.12 64.87 70.66 70.88 (4, 5, 6)
dis 99.34 99.07 98.41 98.28 99.34 99.47 (4, 5, 6)
sylvine 97.46 97.27 92.49 93.95 97.27 96.49 (1, 11)
BNG(tic-tac-toe) 78.04 78.75 73.96 73.71 78.75 79.03 (5, 10, 12)
online_shoppers 90.59 90.11 90.02 90.11 90.63 90.02 (8)
Cardiovascular-Disease-dataset 72.84 73.08 72.96 73.06 73.14 73.09 (5, 8, 12)
credit 78.04 77.06 77.62 77.80 77.95 77.59 (4, 6, 9)
FOREX_audsgd-hour-High 67.26 69.30 57.24 61.06 69.62 70.41 (7, 10, 12)
waveform-5000 86.00 85.30 85.60 85.60 86.40 86.90 (1, 6, 11)
jungle_chess 85.65 85.30 78.55 80.44 86.66 86.85 (10, 11, 12)
BNG(cmc) 57.40 57.78 56.19 56.72 57.72 57.88 (9, 10, 12)
page-blocks 97.35 97.17 96.07 96.71 97.17 97.35 (6, 7, 12)
segment 93.07 93.29 91.99 88.10 93.51 92.64 (1, 12)
website_phishing 90.77 88.93 85.98 87.08 91.88 91.88 (7, 10)
baseball 93.66 94.03 93.28 94.03 93.66 95.15 (10, 11)
pendigits 99.50 99.36 92.81 93.04 99.41 99.45 (3, 4, 12)
Gender_Gap_in_Spanish_WP 60.84 55.79 59.68 60.32 60.53 60.84 (2, 12)
wine-quality-white 62.35 62.04 54.08 55.31 63.57 64.39 (8, 11, 12)
satimage 91.21 92.07 88.72 88.65 91.91 91.91 (8, 11, 12)
mfeat-fourier 90.00 89.75 77.75 82.25 89.50 89.50 (2, 7, 12)
VulNoneVul 98.95 98.95 98.95 98.95 98.95 98.95 (1)
law-school-admission-bianry 100.0 100.0 100.0 100.0 100.0 100.0 (6)
KDD 80.34 78.95 79.34 78.35 81.23 79.94 (1, 8, 10)
Appendix D Influence of Key Modules
We investigate the influence of two key components in TabPFN v2, i.e., the feature engineering that pre-process the raw features of a given tabular dataset and the post-hoc ensembleing.
Feature engineering. TabPFN v2 pre-processes the features of a given tabular datset with various strategies, such as quantile, category shuffling, SVD, and power transform. Specifically, we examine the effects of adding fingerprint features (add_fingerprint_feature) and polynomial features (polynomial_features) to the raw tabular data. The results indicate that TabPFN v2 performs well even without the use of these engineered features, suggesting that, for the benchmark datasets of Ye et al. (2024b), these specific feature engineering techniques do not provide a significant improvement. This finding highlights the robustness of TabPFN v2 and its ability to handle raw features effectively, without the need for extensive pre-processing or feature construction. We show the influence of this step in Figure 9.
Figure 9: Scatter plot comparing the normalized Accuracy/R2 scores. The x-axis represents the normalized Accuracy/R2 scores without Feature Engineering, while the y-axis represents the normalized Accuracy/R2 scores with Feature Engineering. The red dashed line ($y=x$) serves as a reference, indicating equal performance.
Model ensemble. Post hoc ensembling (PHE) involves applying TabPFN v2 to the datasets multiple times with different perturbations and aggregating the predictions of these base models at different temperatures. We show the change of performance of TabPFN v2 w.r.t. the number of ensemble numbers (i.e., the number of base models) in Figure 10. On the benchmark of Ye et al. (2024b), we observe that, overall, ensemble methods improve performance, with larger ensemble sizes yielding better results. However, we also note that even without ensembling, TabPFN v2 performs exceptionally well, and the relative performance gain from ensembling is limited. This suggests that while ensembling can provide further improvements, the base TabPFN v2 model is already highly effective on its own. The equivariant property described in (Arbel et al., 2025) provides insight into this phenomenon. Since TabPFN v2 introduces random tokens to handle heterogeneous features, the model becomes less sensitive to the arbitrary ordering of features, effectively enforcing equivariance in this aspect. As a result, the benefits of ensembling through feature order permutations are less pronounced compared to TabPFN v1.
Figure 10: Box plot of relative performance improvements of TabPFN v2 with post hoc ensembling (PHE) across different ensemble sizes (2, 4, 8, and 16 base models). The relative improvement is calculated as the performance gain over the non-ensemble model, where higher values indicate stronger performance. The box plots show the median, interquartile range (IQR), and outliers for each ensemble size.

### v2

@@
-A Closer Look at TabPFN v2: Strength, Limitation, and Extension
+A Closer Look at TabPFN v2: Understanding Its Strengths and Extending Its Capabilities
 Abstract
-Tabular datasets are inherently heterogeneous, posing significant challenges for developing pre-trained foundation models.
-The recently introduced transformer-based Tabular Prior-data Fitted Network v2 (TabPFN v2) achieves unprecedented in-context learning accuracy across multiple tabular datasets, marking a pivotal advancement in tabular foundation models.
-In this paper, we comprehensively evaluate TabPFN v2 on over 300 datasets, confirming its exceptional generalization capabilities on small- to medium-scale tasks.
-Our analysis identifies randomized feature tokens as a key factor behind TabPFN v2’s success, as they unify heterogeneous datasets into a fixed-dimensional representation, enabling more effective training and inference.
-To further understand TabPFN v2’s predictions, we propose a leave-one-fold-out approach, transforming TabPFN v2 into a feature extractor and revealing its capability to simplify data distributions and boost accuracy.
-Lastly, to address TabPFN v2’s limitations in high-dimensional, large-scale, and many-category tasks, we introduce a divide-and-conquer mechanism inspired by Chain-of-Thought prompting, enabling scalable inference.
-By uncovering the mechanisms behind TabPFN v2’s success and introducing strategies to expand its applicability, this study provides key insights into the future of tabular foundation models.
-Machine Learning, ICML
+Tabular datasets are inherently heterogeneous, presenting significant challenges for developing pre-trained foundation models.
+The recently introduced transformer-based Tabular Prior-data Fitted Network v2 (TabPFN v2) achieves unprecedented in-context learning performance across diverse downstream datasets, marking a pivotal advancement in tabular foundation models.
+In this paper, we take a closer look at TabPFN v2 to examine how it effectively handles heterogeneity and achieves high predictive accuracy, and to explore how its limitations in high-dimensional, many-category, and large-scale tasks can be mitigated.
+We find that TabPFN v2 can infer attribute relationships even when provided with randomized attribute token inputs, eliminating the need to explicitly learn dataset-specific attribute embeddings to address heterogeneity.
+We further show that TabPFN v2 can be transformed into a feature extractor, revealing its ability to construct a highly separable feature space for accurate predictions.
+Lastly, we demonstrate that TabPFN v2’s limitations can be addressed through a test-time divide-and-conquer strategy, enabling scalable inference without requiring re-training.
+By uncovering the mechanisms behind TabPFN v2’s success and introducing strategies to extend its applicability, this study offers key insights into the design of future tabular foundation models.
 1 Introduction
-Tabular data is ubiquitous across diverse applications, including healthcare (Hyland et al., 2020), finance (Kovalerchuk & Vityaev, 2005), and scientific research (Ivanciuc et al., 2007; Hyland et al., 2020).
-In this format, each instance is represented as a vector of attributes, and the task of a machine learning model is to map these vector inputs to their corresponding labels (Borisov et al., 2024).
-Traditionally, tree-based models (Prokhorenkova et al., 2018; Chen & Guestrin, 2016) have dominated the tabular domain, but deep tabular models are increasingly showing potential to advance the field and close the performance gap (Gorishniy et al., 2021; Holzmüller et al., 2024; Ye et al., 2025).
-However, unlike vision and language data, where pre-trained foundation models have driven significant progress (Kirillov et al., 2023; Zhou et al., 2024), tabular data is still desperately awaiting a similar breakthrough (Somepalli et al., 2022; Hollmann et al., 2023; Onishi et al., 2023; Zhu et al., 2023; Ye et al., 2023; van Breugel & van der Schaar, 2024).
+Tabular data is ubiquitous across a wide range of applications, including healthcare [33], finance [44], and scientific research [34, 33].
+In this format, each instance (e.g., a patient’s record) is represented as a vector of attributes, and the goal of a machine learning model is to map these vectors to their corresponding labels [7].
+Traditionally, tree-based models [57, 12] have dominated this domain, but recent advances in deep tabular models are increasingly closing the performance gap [23, 30, 77].
+However, unlike vision and language domains, where pre-trained foundation models have driven significant progress [41, 82], tabular data is still desperately awaiting a similar breakthrough [64, 28, 55, 83, 78, 67].
 A primary challenge arises from the inherent heterogeneity of tabular datasets, which often vary in dimensionality and attribute meanings, making the development of effective and versatile foundation models difficult.
 Additionally, there is an urgent need for such models, as many tabular datasets are small-scale—such as medical data with limited patient numbers.
-Training individual models from scratch for these datasets is highly sensitive to hyperparameters and struggles with insufficient data for generalization (Feurer et al., 2015; Guyon et al., 2019; Han et al., 2024).
-Recently, the Tabular Prior-Fitted Network v2 (TabPFN v2), proposed by Hollmann et al.
-(2025), has emerged as a significant step forward.
-Built on transformer architectures (Vaswani et al., 2017) and pre-trained on gigantic synthetic datasets (Hollmann et al., 2023, 2025), TabPFN v2 can be directly applied to downstream tasks without the need for additional tuning, effectively addressing the heterogeneity of tabular data.
+Training individual models from scratch for these datasets is highly sensitive to hyperparameter choices and often fails to generalize due to limited data [19, 25, 26].
+Recently, the Tabular Prior-Fitted Network v2 (TabPFN v2) [29] has emerged as a significant step forward.
+Built on transformer architectures [68] and pre-trained on gigantic synthetic datasets [28, 29], TabPFN v2 can be directly applied to diverse downstream tasks without additional tuning.
 Specifically, TabPFN v2 takes both a labeled training set and an unlabeled test instance as input, predicting the test label in an “in-context learning” manner.
-When evaluated on multiple datasets covering both classification and regression tasks, TabPFN v2 achieved unprecedented accuracy compared to previous methods.
-Motivated by the remarkable performance of TabPFN v2, this paper pursues four timely objectives to deepen our understanding and advance the field.
-First, we aim to expand the evaluation of TabPFN v2 across a diverse set of scenarios to investigate its generalizability.
-Second, we seek to understand the underlying mechanisms by which TabPFN v2 effectively handles data heterogeneity.
-Third, we strive to uncover how TabPFN v2 achieves its high prediction accuracy.
-Last but not least, we aim to extend TabPFN v2’s applicability beyond the comfort zone defined in Hollmann et al.
-(2025)—datasets with no more than 10,000 samples, 500 dimensions, and 10 classes.
-TabPFN v2 consistently outperforms existing methods on small- to medium-scale datasets.
-We evaluate TabPFN v2 extensively on over 300 tabular datasets spanning various domains, scales, dimensionalities, and tasks (Grinsztajn et al., 2022; McElfresh et al., 2023; Ye et al., 2024b; Rubachev et al., 2025).
-Our results demonstrate TabPFN v2’s exceptional generalizability on small- to medium-scale datasets, outperforming both tree-based models and state-of-the-art deep tabular models.
-However, TabPFN v2’s performance diminishes on large-scale and high-dimensional datasets, indicating areas where further improvements are needed.
-TabPFN v2 effectively handles data heterogeneity through randomized feature tokens.
-A key strength of TabPFN v2 is its ability to manage heterogeneous feature spaces, enabling applications to diverse downstream datasets without additional tuning.
-Given a $d$-dimensional input vector $\in\mathbb{R}^{d}$, TabPFN v2 transforms it into a $d\times k$ matrix and leverages a transformer architecture to handle variability in $d$, following (Song et al., 2019; Gorishniy et al., 2021; Yan et al., 2024).
-The critical innovation lies in the vector-to-matrix transformation: TabPFN v2 uses a “shared” $k$-dimensional vector to lift each attribute value into a $k$-dimensional space.
-To differentiate attributes, a random $k$-dimensional perturbation is added to the shared vector for each attribute, consistent within a dataset but distinct across datasets [^1] This detail was identified through the supplementary material and code of (Hollmann et al., 2025)..
-We view this transformation as a variant of token-based tabular methods, where TabPFN v2 removes the need for dataset- and attribute-specific feature token learning, a limitation that has hindered the transferability of previous methods to diverse downstream datasets.
-Our analysis confirms that these feature tokens require no semantic meaning as long as they are distinct across attributes, demonstrating their efficacy.
-TabPFN v2 simplifies feature distributions among categories.
-To deepen our understanding of TabPFN v2, we investigate whether its superior in-context learning capability is accompanied by the creation of more separable feature representations.
-This analysis is challenging due to the distinct roles of labeled training data and unlabeled test data in TabPFN v2’s in-context learning process, resulting in non-comparable embeddings.
-To overcome this, we propose a leave-one-fold-out strategy to extract features from the training data that align more closely with the test data.
-Our findings reveal that TabPFN v2 effectively transforms tabular datasets into a nearly separable embedding space.
-Remarkably, training a linear model on TabPFN v2’s extracted features achieves accuracy comparable to its in-context learner, underscoring TabPFN v2’s potential as a feature encoder.
-This approach not only offers valuable insights into TabPFN v2 but also paves the way for broader applications and further analysis of its capabilities.
-Divide-and-conquer extends TabPFN v2’s applicability.
-As noted in Hollmann et al.
-(2025) and confirmed by our study, TabPFN v2 encounters challenges with large-scale, high-dimensional, and many-category datasets.
-To address these limitations without extensive retraining, we introduce post-hoc divide-and-conquer strategies, reminiscent of Chain-of-Thought (CoT) prompting used in large-language models (Wei et al., 2022).
-For large-scale datasets, we split the training data into support and query sets, using the support set to extract TabPFN v2’s features for the query and test sets.
-The resulting features then form a new tabular dataset compatible with linear models, akin to representation learning strategies in the early days (Vincent et al., 2010).
-For high-dimensional datasets, we subsample attributes into subsets (Breiman, 2001), apply TabPFN v2 to each subset in parallel, and ensemble the predictions, similar to random forests.
-For a multi-class task with over 10 categories, we introduce decimal encoding inspired by Error Correcting Output Codes (ECOC) (Dietterich & Bakiri, 1995), decomposing it into multiple 10-class problems.
-Empirical results validate these strategies, demonstrating significant gains in TabPFN v2’s accuracy across these data regimes and highlighting the potential of advanced post-hoc methods to further expand the capabilities of tabular foundation models.
-Contributions.
-This paper provides a timely and in-depth investigation into TabPFN v2’s strengths, limitations, and potential extensions, aiming to offer valuable insights to advance tabular foundation models.
-In particular:
-  • We confirm TabPFN v2’s remarkable in-context learning capability across diverse small- to medium-scale tasks, positioning it as a new go-to method for these data regimes.
-  • We highlight randomized feature tokens as the key driver behind TabPFN v2’s handling of heterogeneous datasets and show their applicability to other deep tabular models.
-  • We introduce a simple leave-one-fold-out approach to turn TabPFN v2 into a feature encoder, enhancing its understanding and suggesting broader applications (e.g., for visualization and error diagnosis).
-  • We demonstrate the effectiveness of post-hoc mechanisms to extend TabPFN v2 beyond its designated data regimes, similar to prompting strategies like CoT for LLMs.
+When evaluated across both classification and regression tasks, TabPFN v2 consistently outperforms prior tabular methods, achieving state-of-the-art accuracy.
+Motivated by the remarkable performance of TabPFN v2, we aim to take a step further to understand the mechanisms behind its success [^1] We note that TabPFN v2 [29], like many recent large language models (LLMs) and foundation models, does not release its training data or training recipe.
+Accordingly, our focus is on understanding the properties of the released pre-trained model and exploring ways to extend its applicability.—specifically, how it effectively handles dataset heterogeneity and achieves high predictive accuracy.
+In addition, we investigate how to overcome its current limitations—namely, its suggested data regime of no more than 10,000 samples, 500 dimensions, and 10 classes [29]—ideally without requiring model re-training.
+We outline major insights as follows.
+  1. 1.
+    TabPFN v2 internalizes attribute token learning to handle data heterogeneity.
+Given an instance with $d$ attributes, TabPFN v2 transforms it into a set of fixed-dimensional tokens and uses a transformer architecture to handle variability in $d$, following [65, 23, 75].
+In sharp contrast to prior methods that rely on known attribute semantics (e.g., word vectors) or learn dataset-specific attribute tokens, TabPFN v2 instead employs randomized attribute tokens—resampled at each inference.
+This design “syntactically” allows TabPFN v2 to be directly applied to new downstream datasets with varying dimensionalities and attribute meanings without additional tuning, but raises a fundamental question: how does it still make accurate predictions?
+Our analysis shows that, regardless of the randomness, TabPFN v2 can consistently infer attribute relationships through in-context learning, essentially integrating attribute token learning into the inference itself.
+In short, TabPFN v2 unifies representation learning and prediction within a single forward pass.
+  2. 2.
+    TabPFN v2 can be repurposed as a feature extractor for downstream tasks.
+The exceptional predictive performance of TabPFN v2 suggests that it produces instance-level feature representations that are highly discriminative.
+However, verifying this is non-trivial, as TabPFN v2’s in-context learning mechanism assigns distinct roles to labeled training and unlabeled test instances, resulting in embeddings that are not directly comparable.
+To overcome this, we propose a leave-one-fold-out strategy that enables the extraction of instance features more closely aligned across training and test data.
+Our findings reveal that TabPFN v2 effectively maps tabular instances into a nearly linearly separable embedding space.
+Remarkably, training a linear model on these features yields accuracy comparable to that of TabPFN v2’s in-context learner, highlighting its potential as a powerful feature encoder.
+This not only offers insights into TabPFN v2’s inner workings but also opens the door to broader applications (e.g., visualization and error analysis).
+  3. 3.
+    Test-time divide-and-conquer effectively mitigates TabPFN v2’s limitations.
+As noted in [29], TabPFN v2 faces challenges when applied to high-dimensional, many-category, or large-scale datasets.
+Rather than resorting to model re-training, we show that these limitations can be effectively addressed through carefully designed post-hoc divide-and-conquer strategies, reminiscent of test-time scaling techniques developed for large language models [71, 52].
+Empirical results show significant accuracy gains across these challenging data regimes, highlighting the potential of advanced post-hoc methods to further extend the capabilities of tabular foundation models.
+Remark.
+This paper presents a timely and in-depth investigation into TabPFN v2, offering valuable insights for advancing tabular foundation models.
+While we do not propose a new architecture or training scheme, our contribution lies in the novel analysis and principled extension of TabPFN v2.
+This reflects a growing trend in foundation model research, where understanding, evaluating, and adapting powerful models is increasingly seen as being as impactful as designing new ones.
 2 Related Work
-Foundation Tabular Models.
-Pre-trained models have revolutionized vision and language domains (Kirillov et al., 2023; Zhou et al., 2024), but their adoption in tabular data remains limited due to the inherent heterogeneity of tabular datasets.
-Differences in feature spaces, dimensionalities, and class distributions pose significant challenges for joint training and transferability.
-One solution is to utilize semantic meanings of features, as demonstrated by methods that transform tabular examples into textual representations for large language models (LLMs) (Hegselmann et al., 2023; Zhang et al., 2023a; Wang et al., 2023; Wen et al., 2024).
-Pre-computing feature tokens based on semantic embeddings has also been explored to improve transferability (Yan et al., 2024; Kim et al., 2024).
-However, semantic meanings are often unavailable due to privacy concerns, annotation costs, or lack of describability.
-To address this, Ye et al.
-(2023) propose representing examples through similarity relationships with a fixed number of reference examples, mapping them into a latent space with a fixed dimension.
-The TabPFN family of models (Hollmann et al., 2023, 2025) leverages the in-context learning capabilities of transformers, directly predicting labels by positioning test instances within the context of training examples.
-While TabPFN v1 pads features to a fixed dimension, TabPFN v2 introduces a specialized feature tokenizer to better handle heterogeneity.
-Meta-learning has also been explored to obtain the weights of tabular models for downstream tasks with limited examples (Iwata & Kumagai, 2020; Bonet et al., 2024).
-Other pre-trained models require lightweight fine-tuning on downstream datasets to adapt to heterogeneous feature and class spaces (Liu et al., 2022; Zhang et al., 2023b; Shen et al., 2023; Zhu et al., 2023).
+Learning with Tabular Data.
+Tabular data is prevalent across diverse fields, including healthcare, finance, and education [44, 33, 59, 2].
+Tree-based models, such as XGBoost [12], LightGBM [39], and CatBoost [57], have long dominated this domain.
+However, recent advances in deep neural networks (DNNs) have demonstrated strong potential for tabular data [7].
+Popular architectures like multi-layer perceptrons [23, 38] and Transformers [32] have been adapted to tabular tasks, alongside custom architectures designed specifically for tabular data [42, 69].
+Deep tabular methods can be broadly categorized into two types.
+The first type directly processes raw features [30, 22, 77], sometimes incorporating feature-specific encoding strategies [21].
+The second type tokenizes features, transforming an example into a set of tokens [65, 32, 60].
+Comprehensive benchmarks have been developed to evaluate these methods across diverse datasets [24, 50, 76, 61], highlighting the strengths and weaknesses of deep tabular models in various scenarios.
+Tabular foundation models.
+Pre-trained models have revolutionized the vision and language domains [41, 82], but their adoption in tabular data remains limited due to the substantial heterogeneity across datasets.
+Variations in attribute spaces, dimensionalities, and label distributions present significant challenges for joint training and transferability.
+One solution is to leverage the semantic meanings of attributes, as demonstrated by methods that convert tabular instances into textual descriptions (e.g., “Age is 25, height is 165, …”) and apply large language models for prediction [27, 80, 70, 72].
+Alternatively, some approaches aim to improve transferability by pre-computing attribute tokens based on semantic embeddings [75, 40].
+In practical domains such as healthcare or scientific measurement, the semantic meanings of attributes are often inaccessible due to privacy constraints, annotation costs, or a lack of describability.
+To address this, [78] proposed representing each instance by its similarity profile to a fixed number of nearest neighbor examples, thereby mapping it into a consistent latent space with shared dimensional semantics.
+The TabPFN family [28, 29] leverages the in-context learning capabilities of transformers to directly predict labels by contextualizing test instances among training examples.
+This strategy inspired subsequent pre-trained tabular models such as [49, 15, 58].
+While TabPFN v1 pads attribute vectors to a fixed dimension, TabPFN v2 introduces a specialized attribute tokenizer to handle heterogeneous input spaces.
+Meta-learning has also been explored to generate model weights tailored for downstream tabular tasks with limited data [35, 6, 51].
+Other pre-trained models rely on lightweight fine-tuning to adapt to variations in attribute and label spaces [45, 81, 63, 83].
 Variants of TabPFN.
-TabPFN’s success stems from its pre-training on gigantic synthetic datasets, enabling strong in-context learning performance on small-scale classification tasks (Hollmann et al., 2023).
-Inspired by its capabilities, various applications have been investigated, including tabular data generation (Ma et al., 2024a), anomaly detection (Ruiz-Villafranca et al., 2024), and time series forecasting (Hoo et al., 2025).
-The efficiency of TabPFN is highly sensitive to context size, prompting strategies to enhance scalability and performance (Feuer et al., 2023; Xu et al., 2024).
-Some approaches enhance TabPFN’s performance on downstream tasks by adapting the context (Thomas et al., 2024), fine-tuning specific parts of the model (Feuer et al., 2024; Liu & Ye, 2025), or even pre-training the architecture on real-world datasets (Ma et al., 2024b).
-Nagler (2023) analyzes TabPFN from a bias-variance perspective, shedding light on its generalization capabilities.
-The recently introduced TabPFN v2 extends support to regression tasks and handles larger contexts (Hollmann et al., 2025).
-In this paper, we conduct a comprehensive evaluation of TabPFN v2, analyze its strengths and limitations, and propose solutions to overcome scalability and applicability challenges.
-3 Preliminary
-We formalize the problem of learning with tabular datasets and briefly describe the core mechanism of TabPFN v2.
-Figure 1: Illustration of the TabPFN v2 mechanism (Hollmann et al., 2025) for classification.
+TabPFN’s success stems from its pre-training on massive synthetic datasets, enabling strong in-context learning performance on small-scale classification tasks [28].
+Motivated by its capabilities, researchers have explored a variety of applications, including tabular data generation [48], anomaly detection [62], and time series forecasting [31].
+[54] provided a bias-variance analysis of TabPFN, offering insight into its generalization behavior.
+Another line of research focuses on improving scalability by addressing TabPFN’s sensitivity to context size [17, 74].
+Further strategies to enhance downstream performance include context adaptation with nearest neighbor [66], partial fine-tuning [18, 47], pre-training on real-world datasets [49], scalable ensemble [47], and more powerful and efficient pre-training on synthetic data [58].
+Most of these variants remain restricted to classification tasks due to limitations in TabPFN v1.
+The recently introduced TabPFN v2 [29] extends TabPFN to support regression tasks and accommodate larger context sizes.
+In this paper, we conduct a comprehensive evaluation of TabPFN v2, analyze its strengths, and introduce methods to overcome its scalability and applicability challenges.
+Figure 1: Left: Illustration of TabPFN v2’s mechanism for binary classification [29].
 $\{A_{1},\ldots,A_{d}\}$ denote $d$ attributes of the task.
-Training examples and test instances are combined into a tabular context set and transformed into a ${(N+1)\times d\times k}$ tensor using a combination of learnable and randomized tokens.
-Two types of self-attentions are applied alternately along rows (inter-sample) and columns (inter-feature).
+Training examples and a test instance are combined into a tabular context and transformed into a ${(N+1)\times(d+1)\times k}$ tensor using a combination of learnable and randomized tokens.
+Two types of self-attention are applied alternately across rows (inter-sample) and columns (inter-feature).
 The output token corresponding to the (dummy) label of the test instance is processed through an MLP to generate a 10-class logit.
+Right: Wilcoxon-Holm test at a significance level of 0.05 over 273 small- to medium-scale datasets.
+We omit the 27 datasets used to select TabPFN v2’s checkpoint from the 300 datasets in [76].
+3 Background
 Learning with a single tabular dataset.
 A tabular dataset ${\mathcal{D}}=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{N}$ contains $N$ training examples, corresponding to the rows in a table.
-Each instance ${\bm{x}}_{i}$ is represented by $d$ features or attributes (columns in the table); its label $y_{i}$ belongs to $[C]=\{1,\ldots,C\}$ for a classification task or is a numerical value for a regression task.
-We assume all attributes of an instance are numerical (continuous).
-If categorical (discrete) attributes are present, they are transformed with ordinal/one-hot encoding beforehand.
-The goal of tabular machine learning is to construct a mapping $f$ from an instance to its label.
+Each instance ${\bm{x}}_{i}$ is characterized by $d$ features or attributes (i.e., columns in the table), where $d$ typically varies across datasets.
+Its label $y_{i}$ belongs to $[C]=\{1,\ldots,C\}$ for a classification task or is a numerical value for a regression task.
+We assume that all attributes of an instance are numerical (continuous).
+Categorical (discrete) attributes, if present, are transformed using ordinal or one-hot encoding beforehand.
+The goal of tabular machine learning is to learn a mapping $f$ from instances to their labels.
 Specifically, given an unseen instance ${\bm{x}}^{*}\in{\mathbb{R}}^{d}$ sampled from the same distribution as ${\mathcal{D}}$, the learned mapping $f$ predicts its label as $\hat{y}^{*}=f({\bm{x}}^{*}\mid{\mathcal{D}})$.
-A lower prediction discrepancy between $\hat{y}^{*}$ and the true label $y^{*}$ indicates a stronger generalization ability of $f$.
-TabPFN Variants.
-The original TabPFN implements $f$ for classification via a Transformer-like architecture (Hollmann et al., 2023).
-Both training and test instances are padded to a fixed dimension $k^{\prime}$ (e.g., 100) with zeros.
-Then, ${\bm{x}}_{i}$ and $y_{i}$ are further mapped to $\tilde{{\bm{x}}}_{i}\in{\mathbb{R}}^{k}$ and $\tilde{{\bm{y}}}_{i}\in{\mathbb{R}}^{k}$ with linear transformations.
-TabPFN takes both a labeled training set and an unlabeled test instance as input, predicting the test label in an “in-context learning” manner.
-The task context is formulated as:
-${\mathcal{C}}=\{(\tilde{{\bm{x}}}_{1}+\tilde{{\bm{y}}}_{1}),\ldots,(\tilde{{\bm{x}}}_{N}+\tilde{{\bm{y}}}_{N}),(\tilde{{\bm{x}}}^{*})\}\in{\mathbb{R}}^{k\times(N+1)}\;,$
-with $N+1$ tokens of $k$ dimensions each.
-These tokens are processed through multiple transformer layers, which support variable-length inputs.
-The output token corresponding to the test instance is further mapped to a 10-class logit using a multi-layer perceptron (MLP).
-The recently introduced TabPFN v2 incorporates several modifications to the original TabPFN.
-For example, each attribute in ${\mathcal{D}}$ is mapped to a $k$-dimensional space through linear projection.
-Random perturbations are added to differentiate attributes in the $k$-dimensional space, similar to a kind of positional encoding.
-The core idea is illustrated in Figure 1.
-We view this strategy as a variant of the token-based method (Song et al., 2019; Gorishniy et al., 2021; Yan et al., 2024) and will discuss the construction of tokens and analyze how they handle the challenges of heterogeneous attribute spaces in section 5.
-Together with the mapped label embedding $\tilde{{\bm{y}}}_{i}$, an instance ${\bm{x}}_{i}$ is transformed into tokens with size $(d+1)\times k$.
-For a test instance ${\bm{x}}^{*}$, since its label is unknown, a dummy label (e.g., the average of training set labels) is used to generate the label embedding $\tilde{{\bm{y}}}^{*}$.
-Subsequently, the set of training and test instances is transformed into a ${(N+1)\times(d+1)\times k}$ tensor.
-Two types of self-attentions are alternately applied for inter-sample and inter-feature in-context learning.
-Finally, the output token corresponding to the test instance’s dummy label $\tilde{{\bm{y}}}^{*}$ is extracted and mapped to either a 10-class logit for classification or a single-value logit for regression.
-The weights in TabPFN v2 are pre-trained on synthetic datasets generated using structural causal models (SCMs); the checkpoint is selected based on real-world datasets.
-This pre-training enables TabPFN v2 to be directly applied to downstream tasks without additional tuning, effectively addressing the heterogeneity of tabular data.
-However, the complexity of transformers with respect to the input size limits TabPFN v2’s scalability to small and medium datasets (e.g., $N<10,000$) and its ability to handle classification with fewer than 10 classes.
-Please refer to Hollmann et al.
-(2025) for further details, including feature pre-processing, feature grouping, acceleration, and post-hoc ensembling.
+A smaller discrepancy between $\hat{y}^{*}$ and the true label $y^{*}$ indicates stronger generalizability of $f$.
+TabPFN.
+The original TabPFN implements $f$ for classification using a transformer-like architecture [28].
+Both training and test instances are first zero-padded to a fixed dimension $k^{\prime}$ (e.g., 100).
+Then, ${\bm{x}}_{i}$ and $y_{i}$ are linearly projected to $\tilde{{\bm{x}}}_{i}\in{\mathbb{R}}^{k}$ and $\tilde{{\bm{y}}}_{i}\in{\mathbb{R}}^{k}$, respectively.
+TabPFN processes both a labeled training set and an unlabeled test instance jointly, predicting the test label in an in-context learning manner.
+The task context is defined as ${\mathcal{C}}=\{(\tilde{{\bm{x}}}_{1}+\tilde{{\bm{y}}}_{1}),\ldots,(\tilde{{\bm{x}}}_{N}+\tilde{{\bm{y}}}_{N}),(\tilde{{\bm{x}}}^{*})\}\in{\mathbb{R}}^{(N+1)\times k}$, consisting of $N+1$ tokens, each of dimension $k$.
+These tokens are processed by multiple transformer layers, which accommodate variable-length inputs (i.e., variable $N$).
+The output token corresponding to the test instance is passed through a multi-layer perceptron (MLP) to produce a 10-class logit.
+TabPFN v2.
+The recently proposed variant [29] introduces several key modifications.
+First, each of the $d$ attributes in ${\mathcal{D}}$ is embedded into a $k$-dimensional space, with random perturbations added to differentiate attributes.
+Together with the label embedding $\tilde{{\bm{y}}}_{i}\in\mathbb{R}^{k}$, each training instance ${\bm{x}}_{i}$ is represented by $(d+1)$ tokens with dimension $k$.
+For a test instance ${\bm{x}}^{*}$, where the label is unknown, a dummy label (e.g., the average label of the training set) is used to generate the label embedding $\tilde{{\bm{y}}}^{*}$.
+The full input to TabPFN v2—comprising the training set and the test instance—is thus represented as a tensor of shape ${(N+1)\times(d+1)\times k}$.
+Two types of self-attention are applied in alternation: one over samples (among the $N{+}1$ instances) and the other over attributes (among the $d{+}1$ dimensions), enabling in-context learning along both axes.
+Finally, the output token corresponding to the test instance’s dummy label $\tilde{{\bm{y}}}^{*}$ is extracted and mapped to a 10-class logit for classification or a single-value logit for regression.
+An overview of this process is illustrated in Figure 1 (left).
+The weights in TabPFN v2 are pre-trained on diverse synthetic datasets generated using structural causal models (SCMs), with the checkpoint selected based on real-world datasets.
+For additional details, including feature pre-processing, acceleration, and post-hoc ensembling, please refer to [29].
+Remark.
+In the tabular domain, years of research into deep and foundation models have culminated in TabPFN v2 [29]—a breakthrough that, for the first time, enables deep models to consistently outperform traditional methods without fine-tuning.
+However, due to venue constraints, many technical details were omitted from the main paper.
+For example, the use of randomized tokens was documented in the supplementary material and code.
+In light of this, we aim to systematically analyze TabPFN v2, as we believe such a study is more impactful than proposing yet another architecture.
 4 Comprehensive Evaluation of TabPFN v2
-Building upon the remarkable performance of TabPFN v2, we conduct an extensive evaluation across diverse scenarios to assess its generalizability.
-Specifically, we evaluate TabPFN v2 on over 300 tabular datasets spanning a wide range of domains, scales, dimensionalities, and tasks.
-Figure 2: Wilcoxon-Holm test at a significance level of 0.05 over 261 small- to medium-scale datasets.
-We omit the 12 datasets with more than 10 classes and the 27 datasets used to select the TabPFN v2 checkpoint from the 300 datasets in (Ye et al., 2024b).
+Before presenting our core studies, we first extend TabPFN v2’s evaluation beyond the original set of datasets to over 300, covering a much broader range of domains, attributes, scales, dimensionalities, and tasks [24, 50, 76, 61], aiming to more thoroughly assess its generalizability and limitations.
 4.1 Setups
-To ensure a comprehensive evaluation, we adopt the benchmark proposed in Ye et al.
-(2024b), which includes 300 datasets covering 120 binary classification tasks, 80 multi-class classification tasks, and 100 regression tasks across various domains, dataset sizes, and feature types.
-This benchmark addresses common issues, such as mislabeled datasets and dataset redundancies due to multi-version overlaps (Kohli et al., 2024), ensuring high-quality evaluations.
-Out of the 300 datasets, 12 classification datasets contain more than 10 classes, and 27 datasets overlap with the validation set used for checkpoint selection during TabPFN v2’s pre-training (Hollmann et al., 2025).
-To avoid evaluation bias, we exclude these datasets and report comparison results on the remaining 261 datasets.
-Following the protocol in Gorishniy et al.
-(2021, 2024), each dataset is randomly split into training, validation, and test partitions in proportions to 64%/16%/20%.
-TabPFN v2 predicts the test set labels directly in an in-context learning fashion without further parameter and hyper-parameter tuning.
-Compared (deep) tabular methods perform hyper-parameter tuning using Optuna (Akiba et al., 2019) with 100 trials over the training set, and are early stopped on the validation set.
-All methods are evaluated with 15 random seeds, and the average performance across seeds is reported.
-For classification tasks, we use accuracy (higher is better) as the evaluation metric, while regression tasks are evaluated using Root Mean Square Error (RMSE, lower is better).
-Figure 3: Probability of Achieving the Best Accuracy/RMSE (PAMA) across 261 small- to medium-scale datasets.
-The values inside the rectangles represent the percentage of datasets where each method achieves the best accuracy.
-4.2 Results
-We compare TabPFN v2 against 26 representative (deep) tabular methods on 261 datasets (detailed references are in the appendix).
-To assess statistical significance, we apply the Wilcoxon-Holm test at a 0.05 significance level (Demsar, 2006).
-The critical difference analysis results are presented in Figure 2.
-The results demonstrate that TabPFN v2 consistently outperforms both tree-based methods, such as CatBoost (Prokhorenkova et al., 2018), and deep tabular models, including RealMLP (Holzmüller et al., 2024), ModernNCA (Ye et al., 2025), TabR (Gorishniy et al., 2024), and FT-Transformer (Gorishniy et al., 2021), with statistically significant differences.
-To further analyze performance, we consider the Probability of Achieving the Best Accuracy/RMSE (PAMA) (Delgado et al., 2014), which measures the percentage of datasets where a method achieves the best performance.
-As shown in Figure 3, TabPFN v2 achieves the highest PAMA score, with 25.06% of datasets yielding the best results, outperforming other methods like ModernNCA (with a percentage of 13.73%) and CatBoost (9.48%).
-These results validate TabPFN v2’s exceptional generalizability on small- to medium-scale tabular datasets, further enhanced by its efficiency and lack of reliance on hyper-parameter tuning.
-Table 1: Average rank (lower is better) of TabPFN v2 and representative comparison methods on 18 high-dimensional datasets and 18 large-scale datasets.
-Full results with our extensions are in Figure 6 and Figure 8.
-$\downarrow$ TabPFN v2 CatBoost MNCA RealMLP LR
+We first adopt the benchmark from [76], comprising 120 binary classification, 80 multi-class classification, and 100 regression tasks.
+It resolves common issues such as mislabeled data and redundancies from overlapping dataset versions [43], enabling more reliable evaluations.
+Out of the 300 datasets, 27 belong to the validation set used for checkpoint selection in TabPFN v2 [29].
+To avoid evaluation bias, we exclude these datasets and report results on the remaining 273 datasets.
+Following the protocol in [22, 23], each dataset is randomly split into training, validation, and test partitions in a 64%/16%/20% ratio.
+TabPFN v2 predicts test set labels directly using in-context learning, without any additional parameter or hyperparameter tuning.
+Baseline tabular methods—both deep and traditional—perform hyperparameter tuning using Optuna [1], with 100 trials on the training set and early stopping based on validation performance.
+All methods are evaluated using 15 random seeds, and we report the average performance across seeds.
+For classification tasks, we report accuracy (higher is better), while regression tasks are evaluated using Root Mean Square Error (RMSE; lower is better).
+For tasks with more than 10 classes, we adopt the built-in Error-Correcting Output Codes (ECOC) [16] strategy for TabPFN v2.
+Figure 2: Probability of Achieving the Maximum Accuracy or Minimum RMSE across 273 datasets. Values inside rectangles show the percentage of datasets on which a method achieves the best result.
+$\downarrow$ TabPFN v2 CatB MNCA R-MLP LR
 High-Dim 3.36 2.82 4.41 2.14 2.27
 Large-Scale 3.97 1.89 2.27 1.94 4.47
-4.3 Additional Evaluations
-To examine TabPFN v2’s performance on larger and more complex datasets, we conduct additional evaluations on 18 large-scale datasets with $N\times d>1,000,000$ and 18 high-dimensional datasets with $d\geq 2,000$ (Jiang et al., 2024).
-For high-dimensional datasets, we follow the same evaluation protocol as before.
-For large-scale datasets, due to the prohibitive cost of hyper-parameter tuning, default hyper-parameters are used for all methods.
-The average ranks of some representative methods are summarized in Table 1.
-Full results are in section 7 with our extensions.
-While TabPFN v2 performs superiorly on small- to medium-scale datasets, its performance degrades on large-scale and high-dimensional datasets.
-For instance, on large-scale datasets, TabPFN v2 ranks lower than both CatBoost and RealMLP, highlighting its scalability limitations.
-On high-dimensional datasets, TabPFN v2 fails to outperform even the simple Logistic Regression (LR) model, further emphasizing the need for enhancements to better handle these challenging scenarios.
-These results highlight TabPFN v2’s limitations in handling high-dimensional and large-scale datasets, which may result from two key factors.
-First, TabPFN v2 is pre-trained exclusively on small- and medium-scale synthetic datasets, creating a mismatch when applied to large-scale datasets.
-Second, the complexity of transformers in TabPFN v2 scales with both $N$ (the number of instances) and $d$ (the number of features), meaning that increasing either value significantly can lead to a substantial rise in computational complexity, which in turn may reduce the model’s effectiveness.
-In summary, while TabPFN v2 demonstrates exceptional performance on small- to medium-scale datasets, scaling it up to handle larger datasets, higher-dimensional feature spaces, and multi-class tasks exceeding 10 classes is a crucial and promising direction for future research.
-5 Heterogeneity Mechanisms Analysis
-As described in section 3, TabPFN v2 handles data heterogeneity by embedding a $d$-dimensional input vector into a $k$-dimensional space with size $d\times k$.
-This enables TabPFN v2 to adapt to diverse downstream datasets with varying feature dimensions seamlessly.
-In this section, we analyze this mechanism and revisit it as a variant of the token-based deep tabular method.
-Further analyses of the key components in TabPFN v2 are provided in Appendix D.
-5.1 Analysis of the Heterogeneity Handling Strategy
-Traditional deep tabular models often struggle with dataset heterogeneity due to their reliance on fixed feature semantics.
-For example, token-based approaches like AutoInt (Song et al., 2019) and FT-Transformer (FT-T) (Gorishniy et al., 2021) learn dataset- and attribute-specific feature tokens during training, which become ineffective when applied to new datasets with different feature meanings.
-These methods typically generate $d$ attribute-specific tokens $[{\bm{r}}_{1},\ldots,{\bm{r}}_{d}]\in{\mathbb{R}}^{k\times d}$, one for each dimension, and transform an instance ${\bm{x}}_{i}$ into $[{x_{i}^{1}}\cdot{\bm{r}}_{1},\ldots,{x_{i}^{d}}\cdot{\bm{r}}_{d}]\in{\mathbb{R}}^{k\times d}$, where $x_{i}^{j}$ is the $j$-th element of ${\bm{x}}_{i}$.
-However, when transitioning to a new dataset with $d^{\prime}$ features, these learned tokens cannot be reused, requiring feature tokens to be learned anew.
-Previous methods address this by relying on semantic descriptions of attributes (Yan et al., 2024) or assuming overlap between feature sets to construct a mapping (Onishi et al., 2023; Zhou et al., 2023).
-In contrast, TabPFN v2 eliminates this dependency through randomized feature tokens.
-Specifically, TabPFN v2 applies a shared $k$-dimensional vector ${\bm{u}}\in\mathbb{R}^{k}$ to lift raw attribute values into a $k$-dimensional space.
-To differentiate attributes, TabPFN v2 adds a random $k$-dimensional perturbation for each attribute, consistent within a dataset but distinct across datasets.
-Formally, the randomized tokens are denoted as:
-${\bm{R}}={\bm{W}}{\bm{P}}\in{\mathbb{R}}^{k\times d},$ (1)
-where ${\bm{P}}\in{\mathbb{R}}^{k^{\prime}\times d}$ is randomly generated, and ${\bm{W}}\in{\mathbb{R}}^{k\times k^{\prime}}$ is a learnable pre-condition projection matrix that maps the random values to the required dimensionality ($k^{\prime}<k$).The $j$-th column of ${\bm{R}}$, ${\bm{r}}_{j}\in{\mathbb{R}}^{k}$, serves as the specific token for the $j$-th attribute and is combined with the shared vector ${\bm{u}}$.
-Then, an instance ${\bm{x}}_{i}$ is transformed through
-$[{x_{i}^{1}}\cdot({\bm{u}}+{\bm{r}}_{1}),\ldots,{x_{i}^{d}}\cdot({\bm{u}}+{\bm{r}}_{d}),\tilde{{\bm{y}}}_{i}]\in{\mathbb{R}}^{k\times(d+1)}\;.$ (2)
-This randomized tokenization ensures that distinct attributes—irrespective of their semantic meanings—are mapped to unique, nearly orthogonal directions in the high-dimensional space.
-This is akin to one-hot encoding but operates in a continuous over-parameterized regime, where random vectors naturally approximate orthogonality.
-Through its transformer layers, TabPFN v2 dynamically combines these basis vectors during inference, weighting their contributions based on the input context.
-This process is analogous to inverse PCA: while PCA decomposes data into orthogonal components, TabPFN v2 reconstructs heterogeneous features by mixing randomized bases scaled by their attribute values.
-This approach avoids the need for pre-defined feature semantics, while the shared vector ${\bm{u}}$ introduces a stable inductive bias that supports cross-dataset generalization.
-Pre-training on diverse and abundant synthetic datasets further equips TabPFN v2 with the capability to adapt these bases across diverse tasks, unifying disparate tabular datasets into a coherent latent framework.
-This randomized token strategy effectively removes the need for dataset- and attribute-specific feature token learning, addressing a significant limitation of prior methods and improving transferability to diverse downstream datasets.
-Figure 4: The change of ranks between FT-T, MNCA, and the variants with randomized tokens over 45 tabular datasets.
-(The rank is calculated together with the 26 compared baselines.) Values in the parentheses denote the average rank.
-5.2 Effectiveness of Randomized Feature Tokens
-To further evaluate the effectiveness of randomized tokens, we integrate this strategy with two representative methods: FT-Transformer (FT-T) by Gorishniy et al.
-(2021) and ModernNCA (MNCA) by Ye et al.
-(2025).
-FT-T with Randomized Tokens (FT-T).
-FT-T is a token-based method that learns specific tokens $\{{\bm{r}}_{1},\ldots,{\bm{r}}_{d}\}\in{\mathbb{R}}^{d\times k}$ for each attribute.
-After transforming an instance into a set of tokens, a transformer processes these tokens together with a learned class token.
-The output token corresponding to the class token is then fed into a linear classifier.
-We replace FT-T’s learned tokens with the randomized tokens used in TabPFN v2, retrain FT-T (denoted as FT-T), and evaluate the resulting performance.
+>10 classes 3.33 2.75 3.17 1.42 4.33
+Figure 3: Average rank (lower is better) of TabPFN v2 and representative baselines on 18 high-dimensional, 18 large-scale, and 12 datasets with more than 10 classes.
+Full results with our extensions are in Figure 6.
+4.2 Empirical Results: Strengths of TabPFN v2
+We compare TabPFN v2 against 26 representative tabular methods (see Appendix A for full references).
+To assess statistical significance, we apply the Wilcoxon-Holm test with a significance level of 0.05 [14].
+As shown in the critical difference diagram in Figure 1 (right), TabPFN v2 consistently outperforms both tree-based methods, such as CatBoost [57], and deep tabular models, including RealMLP [30], ModernNCA [77], TabM [20], TabR [22], and FT-Transformer [23].
+To further assess performance, we report the Probability of Achieving the Maximum Accuracy or Minimum RMSE (PAMA) [13], which measures the proportion of datasets on which a method achieves the best performance.
+As shown in Figure 3, TabPFN v2 attains the highest score, delivering the top results on 26.02% of the datasets—outperforming other methods such as ModernNCA (11.79%) and TabM (9.78%).
+These results underscore TabPFN v2’s strong generalizability.
+4.3 Empirical Results: Limitations of TabPFN v2
+The above evaluation focuses on small- to medium-scale datasets, specifically those with fewer than 10,000 examples.
+However, as noted in [29], the computational complexity of transformers constrains TabPFN v2’s ability to scale effectively to datasets with larger sample sizes or higher dimensionality.
+To verify this, we conduct additional evaluations on 18 high-dimensional datasets with $d\geq 2{,}000$ [37] and 18 large-scale datasets where $N\times d>1{,}000{,}000$.
+For high-dimensional datasets, we follow the same protocol as before.
+For large-scale datasets, due to the prohibitive cost of hyperparameter tuning, default hyperparameters are used for all methods.
+The average ranks of several representative methods are summarized in Figure 3.
+The full results—along with our extensions—are in Section 7.
+As shown, TabPFN v2’s performance degrades on both large-scale and high-dimensional datasets.
+On large-scale datasets, it ranks below both CatBoost and RealMLP; on high-dimensional datasets, it even falls behind the simple Logistic Regression (LR) model.
+Beyond these two limitations, Figure 3 also reports results on the 12 datasets in Section 4.2 that contain more than 10 categories, where the ECOC strategy currently used by TabPFN v2 appears ineffective in achieving high accuracy.
+While increased computational complexity may contribute to this reduced effectiveness, we hypothesize that TabPFN v2 was pre-trained exclusively on small- to medium-scale synthetic datasets with fewer than 10 categories, leading to a mismatch when applied to larger or more complex real-world data.
+These results underscore the limitations of TabPFN v2, suggesting areas for further improvement.
+5 How Does TabPFN v2 Effectively Handle Data Heterogeneity?
+Section 4 demonstrates TabPFN v2’s excellent generalizability to heterogeneous downstream tasks while also highlighting its current limitations.
+In the rest of the paper, we first examine the mechanisms behind its strengths, followed by methods to overcome its limitations.
+5.1 Diving into TabPFN v2’s Mechanisms for Heterogeneous Input
+Revisiting the problem.
+As noted in Sections 3 and 2, tabular datasets often differ in both the number of attributes (i.e., $d$) and the semantics of those attributes.
+Even when dimensionalities match, the dimensional semantics from different datasets are typically not directly comparable.
+A robust tabular foundation model must therefore handle such heterogeneity effectively, enabling it to learn from diverse pre-training datasets and transfer its capabilities to new downstream tasks.
+Tokenization as a feasible solution.
+Among prior approaches, the most relevant to TabPFN v2 are token-based methods [65, 23, 75].
+The core idea is to convert a $d$-dimensional instance ${\bm{x}}\in\mathbb{R}^{d}$ into a set of $d$ fixed-dimensional tokens (each of dimension $k$), with one token per attribute.
+This enables the use of transformer architectures, which naturally accommodate variability in $d$ across datasets.
+To embed each attribute into a shared $k$-dimensional space, prior work either uses pre-defined semantic embeddings [75] (e.g., word vectors of attribute names) or learns dataset-specific embeddings [65, 23].
+Given $d$ attribute-specific tokens $[{\bm{r}}_{1},\ldots,{\bm{r}}_{d}]\in{\mathbb{R}}^{d\times k}$, each instance ${\bm{x}}_{i}\in{\mathbb{R}}^{d}$ can then be transformed into $\left[x_{i}^{1}\cdot{\bm{r}}_{1},\ldots,x_{i}^{d}\cdot{\bm{r}}_{d}\right]\in{\mathbb{R}}^{d\times k}$, where $x_{i}^{j}$ denotes the $j$-th element of ${\bm{x}}_{i}$.
+By embedding all attributes into a shared, fixed-dimensional feature space, this approach allows the transformer to learn transferable patterns and knowledge from heterogeneous datasets.
+Difficulty in direct generalization.
+While appealing, the aforementioned methods face a notable challenge when applied to downstream tasks: attribute names or semantics are not always accessible, as discussed in Section 2.
+Although it is possible to learn dataset-specific attribute tokens, doing so incurs additional computational cost and prohibits the reuse of previously learned tokens.
+Consequently, this limits the direct generalization of the foundation model to new tasks.
+TabPFN v2’s mechanisms.
+TabPFN v2 builds on prior token-based methods by representing each instance ${\bm{x}}_{i}$ as a sequence of tokens.
+However, rather than assigning a deterministic token to each attribute, TabPFN v2 samples random tokens at inference time.
+Specifically, it learns a shared vector ${\bm{u}}\in\mathbb{R}^{k}$ that lifts each element of ${\bm{x}}_{i}$ into a $k$-dimensional space.
+To distinguish attributes, TabPFN v2 adds a random perturbation to each one.
+[^2] This detail was identified from the supplementary material and code of [29].
+For the $j$-th attribute (i.e., $x_{i}^{j}$), the representation becomes $x_{i}^{j}\cdot{\bm{u}}+{\bm{r}}_{j}$, where ${\bm{r}}_{j}={\bm{W}}{\bm{p}}_{j}$.
+Here, ${\bm{p}}_{j}\in\mathbb{R}^{k^{\prime}}$ is a randomly generated vector, and ${\bm{W}}\in\mathbb{R}^{k\times k^{\prime}}$ is a learned projection matrix that conditions the perturbation.
+The full instance ${\bm{x}}_{i}$ is then represented as:
+$[{x_{i}^{1}}\cdot{\bm{u}}+{\bm{r}}_{1},\ldots,{x_{i}^{d}}\cdot{\bm{u}}+{\bm{r}}_{d},\tilde{{\bm{y}}}_{i}]\in{\mathbb{R}}^{k\times(d+1)},$ (1)
+where the last token $\tilde{{\bm{y}}}_{i}$ encodes the label information (see Figure 1 for an illustration).
+(b) Layer 3
+(c) Layer 6
+(d) Layer 9
+(e) Layer 12
+(f) Legend
+Figure 4: Attribute relationships inferred by TabPFN v2.
+The first and third rows show PCA projections of the $d$ attribute tokens from all $N$ training instances at various layers for the churn and bank datasets.
+Colors indicate different attributes (see legend on the right).
+The second and fourth rows display the attribute-wise attention maps.
+Each matrix cell represents the average attention weight between attributes; the last element along each axis (e.g., the last column and row) corresponds to the label.
+The first plots in the second and fourth rows summarize the cosine similarity of attention maps across random seeds.
+See text for details.
+The abbreviations of feature names of these two datasets are explained in Table 4 and Table 5.
+5.2 TabPFN v2 Internalizes Attribute Token Learning
+TabPFN v2’s randomized tokenization scheme eliminates the need to define attribute- or dataset-specific tokens across tasks, thereby syntactically enabling direct application of the pre-trained model.
+At first glance, this may appear to disregard the valuable semantic meaning of attributes.
+However, we show that through in-context learning, TabPFN v2 can consistently infer relationships among attributes within a dataset—despite the randomness introduced during tokenization.
+Specifically, we analyze the behavior of attribute tokens from three perspectives, as illustrated in Figure 4, using two representative downstream datasets: churn and bank.
+First, we visualize the attribute token embeddings (i.e., the first $d$ tokens in Equation 1) across all $N$ training instances.
+The first and third rows of Figure 4 present PCA projections of these $N\times d$ tokens at the input stage and after transformer layers $\{3,6,9,12\}$, with colors indicating different attributes.
+Initially, tokens from different attributes appear randomly scattered.
+However, as the input progresses through the transformer layers, these tokens become increasingly structured.
+For example, in the bank dataset (which predicts term deposit subscriptions), attributes such as “job,” “education,” and “balance” eventually cluster into semantically coherent groups.
+Second, we examine attribute-wise attention patterns across layers, including attention to the label token.
+The second and fourth rows of Figure 4 show heatmaps of attention weights averaged over heads and training instances.
+Each row in the heatmap represents the attention distribution from one attribute to all others; the last element along each axis (e.g., the last column and row) corresponds to the label.
+Darker shades indicate stronger attention.
+We observe a consistent pattern across datasets: in early layers, attributes predominantly attend to the label token, likely to absorb task-specific signals.
+In intermediate layers, attention becomes more uniformly distributed, facilitating information exchange across attributes.
+In deeper layers, attention concentrates on semantically relevant attributes, suggesting the model has inferred inter-attribute relationships useful for prediction.
+Lastly, to assess robustness against random token initialization, we compute attribute-wise attention weights across 10 runs.
+The cosine similarities and variances of these attention patterns are summarized in the first plots of the second and fourth rows in Figure 4.
+The results confirm that attention patterns remain stable across runs, except for the first layer.
+Remark.
+The above results suggest that TabPFN v2 can reliably infer meaningful attribute relationships through in-context learning.
+Although input embeddings are randomized, they consistently differentiate attributes across instances—functionally akin to one-hot encodings.
+Pre-training on diverse tasks thus enables the model to extract predictive patterns (e.g., co-occurrence across attributes, value distributions, and relative magnitudes) directly from the statistical structure of each dataset, without relying on pre-defined attribute semantics.
+As a result, the model effectively internalizes attribute token learning within the inference process.
+6 TabPFN v2 Can Be Transformed into an Effective Feature Encoder
+In Section 5, we show that TabPFN v2’s in-context learning process infers meaningful attribute relationships. Here, we examine whether TabPFN v2 also produces separable instance representations.
 (a) Raw Feature
-(b) Vanilla Extraction
+(b) Vanilla
 (c) Layer 3
@@
 (f) Layer 12
-Figure 5: Visualization of extracted embeddings for four datasets: churn (first row, two classes), bank (second row, two classes), KDD (third row, two classes), and website_phishing (fourth row, three classes).
-We use crosses to denote training examples and circles to denote test examples.
-(a) shows the raw features, while (b) presents the embeddings extracted using the vanilla strategy.
-(c)-(f) depict the embeddings obtained using our proposed methods across different layers.
-The accuracy value is calculated by training a linear model (logistic regression) over the extracted embeddings on the training set and predicting on the test set.
-MNCA with Randomized Tokens (MNCA).
-ModernNCA is a neighborhood-based method that predicts an instance based on its similarity to training instances in a latent space.
-While ModernNCA typically operates on raw features, we adapt it into a token-based method by applying ModernNCA to the token output from FT-T before the final linear layer, incorporating the randomized tokens as above.
-We evaluate these FT-T and MNCA variants over the tiny-bechmark2 from Ye et al.
-(2024b), which contains 45 representative tabular datasets.
-The results in Figure 4 demonstrate how the performance rank changes when replacing the specific tokens with the randomized tokens.
-Remark.
-Our experiments confirm that feature tokens do not significantly rely on semantic meanings.
-Using randomized tokens only slightly increases the average ranks as long as they are distinct across attributes.
-This highlights the robustness and universality of the randomized token strategy, demonstrating its effectiveness in handling heterogeneous datasets and its potential to improve generalization across diverse tasks once pre-trained.
-6 TabPFN v2 as a Feature Encoder
-To better understand TabPFN v2, we examine whether its in-context learning capability also produces separable and meaningful feature representations.
-6.1 Challenges in Embedding Extraction
-As described in section 3, the output tokens from the multiple transformer layers in TabPFN v2 correspond one-to-one with the input tokens, forming a tensor of size $(N+1)\times(d+1)\times k$.
-The final token, corresponding to the (dummy) label embedding $\tilde{{\bm{y}}}^{*}$ of the test instance, is passed through an MLP block to produce the output.
-An intuitive idea is to treat the output tokens associated with the training label embeddings $\{{\bm{y}}_{i}\}_{i=1}^{N}$—prior to the MLP block—as the extracted embeddings for the training data.
-However, this straightforward method has one fundamental limitation due to the distinct roles of labeled training and unlabeled test data in TabPFN v2’s in-context learning process.
-Specifically, the label embeddings for the training instances are derived from their true labels, while those for the test instances rely on dummy labels.
-This discrepancy results in output embeddings that are inherently non-comparable between the training and test instances.
+Figure 5: Visualization of the extracted instance features from two datasets: churn (first row, binary) and bank (second row, binary).
+Blue and red indicate classes; darker crosses and lighter circles denote training and test samples.
+(a) shows the raw input features (e.g., ${\bm{x}}_{i}$), while (b) presents embeddings from the vanilla strategy.
+(c)-(f) display embeddings produced by our method at different layers.
+Classification accuracy is reported by training a linear logistic regression model on the training embeddings and evaluating on the test set.
+6.1 Naive Feature Extraction Fails
+As shown in Figure 1 (left), TabPFN v2 makes predictions based on the output token corresponding to the (dummy) label embedding $\tilde{{\bm{y}}}^{*}$ of the test instance.
+This output token can thus be interpreted as the instance embedding for the test example.
+A natural extension to obtain embeddings for the training instances is to extract the output tokens corresponding to the training label embeddings $\{\tilde{{\bm{y}}}_{i}\}_{i=1}^{N}$.
+However, as shown in Figure 5 (b), this naive approach leads to surprisingly discrepant feature distributions between training (darker cross) and test (lighter circle) examples.
+As a result, a linear classifier trained on these embeddings performs poorly on the test set.
+We attribute this discrepancy to the distinct roles of labeled training data and unlabeled test data in TabPFN v2’s in-context learning process.
+Specifically, the label embeddings for the training instances are derived from true labels, whereas those for the test instances rely on dummy labels.
+This mismatch renders the resulting output embeddings non-comparable between training and test instances.
 6.2 Leave-one-fold-out Feature Extraction
 To address this challenge, we propose a leave-one-fold-out strategy that enables the extraction of comparable embeddings for training and test data.
-Within the TabPFN v2 framework, we define the support set ${\mathcal{S}}$ as the examples with true labels and the query set ${\mathcal{Q}}$ as the examples with dummy labels.
-Under the standard configuration, ${\mathcal{S}}$ corresponds to the labeled training set, while ${\mathcal{Q}}$ corresponds to the unlabeled test instances.
-A key trade-off arises: ${\mathcal{S}}$ needs to include as much labeled data as possible to propagate knowledge from ${\mathcal{S}}$ to ${\mathcal{Q}}$ effectively in the in-context learning process.
-However, to ensure that embeddings for training and test instances are comparable, training instances must also be included in ${\mathcal{Q}}$ and paired with dummy label embeddings.
-To reconcile these requirements, we split the training set into multiple folds (e.g., 10 folds).
-One fold is used as ${\mathcal{Q}}$ for embedding extraction, while the remaining folds constitute ${\mathcal{S}}$ with true training labels.
-This ensures that ${\mathcal{S}}$ retains sufficient label information while allowing embeddings to be extracted for the training data in ${\mathcal{Q}}$.
-An extreme version of this strategy operates in a leave-one-out fashion, with only one training instance placed in ${\mathcal{Q}}$ at a time.
-Results in Figure 5 (c)-(f) demonstrate that embeddings extracted using this strategy (with 10 folds) more effectively reveal dataset properties.
-We observe that TabPFN v2 simplifies tabular data distributions and transforms datasets into nearly separable embedding spaces, particularly in the embeddings after intermediate transformer layers.
-Table 2: Average rank (lower is better) of TabPFN v2 and a linear classifier trained on the extracted embeddings across 29 classification datasets.
-“Combined” refers to an approach where embeddings from up to three layers (from the 12 available layers) are selected based on the validation set performance and concatenated.
+In the TabPFN v2 framework, we treat examples with true labels as the support set ${\mathcal{S}}$, and those with dummy labels as the query set ${\mathcal{Q}}$.
+Under the standard configuration, ${\mathcal{S}}$ corresponds to the labeled training set and ${\mathcal{Q}}$ to the unlabeled test instances.
+To extract comparable embeddings for the training examples, they must also be included in ${\mathcal{Q}}$ with dummy label embeddings.
+This, however, creates a dilemma: effective in-context learning relies on maximizing the size of ${\mathcal{S}}$ to ensure sufficient knowledge transfer to ${\mathcal{Q}}$.
+Including training examples in ${\mathcal{Q}}$ thus competes with the need to keep ${\mathcal{S}}$ as large as possible.
+To overcome this dilemma, we partition the training set into multiple folds (e.g., 10).
+In each round, one fold serves as ${\mathcal{Q}}$—with dummy labels used for embedding extraction—while the remaining folds form ${\mathcal{S}}$ with true labels.
+This setup preserves sufficient label supervision in ${\mathcal{S}}$ while enabling the extraction of embeddings for training instances in ${\mathcal{Q}}$.
+Results in Figure 5 (c)-(f) show that embeddings extracted by this strategy (with 10 folds) more faithfully capture dataset structure.
+We observe that TabPFN v2 simplifies the original tabular data distributions, transforming datasets into nearly linearly separable embedding spaces—especially after intermediate transformer layers.
+6.3 Validation of Embedding Quality
+To validate the quality of the extracted embeddings, we train a logistic regression on embeddings derived from the training set and evaluate it on test set embeddings.
+Table 1: Average rank (lower is better) of TabPFN v2 and linear classifiers trained on the extracted embeddings across 29 classification datasets.
+Combined: embeddings from up to three layers (from the 12 available layers) are selected and concatenated, based on the validation set performance.
 $\downarrow$ TabPFN v2 Vanilla Layer 6 Layer 9 Layer 12 Combined
-Rank 3.12 3.43 5.03 4.72 2.45 2.24
-6.3 Validation of Embedding Quality
-To validate the quality of the extracted embeddings, we train a logistic regression on top of the extracted embeddings.
-Specifically, the classifier is trained on embeddings derived from the training set and evaluated on test set embeddings.
-The average rank across 29 classification datasets from the tiny benchmark2 in Ye et al.
-(2024b) is reported in Table 2.
-Remarkably, training a linear classifier on these extracted embeddings yields performance comparable to TabPFN v2’s in-context learner.
-Embeddings from intermediate layers (or their selected concatenations) sometimes achieve even better results.
-These findings highlight TabPFN v2’s potential as a robust feature encoder, offering valuable insights into its architecture and paving the way for broader applications in tabular data analysis and representation learning.
-7 Extensions of TabPFN v2
-We aim to expand TabPFN v2’s applicability beyond the boundaries outlined in Hollmann et al.
-(2025), which restricts its use to datasets with no more than 10,000 samples, 500 dimensions, and 10 classes.
-To address these limitations without requiring extensive retraining, we propose post-hoc divide-and-conquer strategies inspired by Chain-of-Thought (CoT) prompting in large-language models (Wei et al., 2022).
-We reformulate challenging tasks into multiple simpler subtasks that TabPFN v2 can effectively handle.
-Figure 6: Mean accuracy on 18 high-dimensional datasets. “*” denote our extension of TabPFN v2, and “-PCA” is another variant using PCA to reduce dimensions.
+Rank 2.69 5.97 4.28 4.00 2.12 1.94
+The average rank across 29 classification datasets from the tiny benchmark2 in [76] is reported in Table 1.
+Remarkably, training a simple linear classifier on the extracted embeddings achieves performance comparable to that of TabPFN v2’s in-context learner.
+Furthermore, concatenating embeddings from multiple layers (e.g., both output and intermediate representations) can sometimes lead to even better results.
+These findings underscore TabPFN v2’s potential as a strong and versatile feature encoder, suggesting broader applicability in downstream tasks such as tabular data analysis.
+7 Improving TabPFN v2 via Test-Time Divide-and-Conquer
+This section addresses the limitations discussed in Section 4.3, aiming to extend TabPFN v2’s applicability beyond the boundaries.
+Specifically, we propose post-hoc divide-and-conquer strategies inspired by Chain-of-Thought (CoT) prompting [71], which decompose challenging tasks into simpler subtasks that TabPFN v2 can effectively handle.
 7.1 High Dimension Datasets
-High-dimensional datasets (Jiang et al., 2024) present a unique challenge due to the quadratic complexity of TabPFN v2 with respect to the number of dimensions.
-To mitigate this, we propose subsampling the feature space into smaller subsets, processing each subset independently, and combining the predictions in an ensemble (bagging) fashion, similar to random forests (Breiman, 2001).
-In detail, we iteratively sample $m$ subsets (we set $m=4$ in experiments), each containing $d^{\prime}<d$ randomly selected attributes.
-For each subset, we leverage the TabPFN v2’s ability to handle lower-dimensional data to obtain predictions.
-The final result aggregates these outputs using averaging (for regression) or majority voting (for classification).
-Figure 6 summarizes the results 18 high-dimensional classification datasets.
-TabPFN v2 with the proposed divide-and-conquer and then ensemble strategy significantly increases the mean accuracy (the second-highest) against the vanilla TabPFN v2, effectively extending TabPFN v2’s scalability to datasets with $d\geq 2000$.
-Another variant that utilizes PCA to reduce the dimensionality together with bagging also resolves the dimensionality issue to some extent.
-Figure 7: Mean accuracy on 12 datasets with more than 10 classes.
-“PFNv2” denotes our extension, while “PFNv2-ECOC” denotes the multi-class ECOC strategy implemented by Hollmann et al.
-(2025) in their February 2025 code update.
+High-dimensional datasets [37] present a unique challenge due to the quadratic complexity of TabPFN v2 with respect to the number of dimensions.
+To mitigate this, we propose subsampling the feature space into smaller subsets, processing each subset independently, and combining the predictions in an ensemble (bagging) fashion, similar to random forests [8].
+In detail, we iteratively sample $m$ subsets, each containing $d^{\prime}<d$ randomly selected attributes.
+For each subset, we leverage TabPFN v2’s ability to handle lower-dimensional data to obtain predictions.
+We denote this divide-and-conquer and then ensemble strategy as TabPFN v2, which aggregates outputs using averaging (for regression) or majority voting (for classification).
+Figure 6 (left) summarizes the results on 18 high-dimensional classification datasets.
+A variant that utilizes PCA to reduce the dimensionality, together with bagging, resolves the dimensionality issue to some extent.
+TabPFN v2 with $d^{\prime}=500$ and $m=\left\lceil d/{d^{\prime}}\right\rceil$ significantly increases the mean accuracy (to the highest), effectively extending TabPFN v2’s scalability to datasets with $d\geq 2000$.
 7.2 Multi-Class Problems with More Than 10 Classes
@@
 For each digit position $j\in\{1,\ldots,t\}$, we train a separate TabPFN v2 model $f_{j}$ to predict the $j$-th digit.
-For example, in a 15-class problem, we decompose it into two subtasks: the first predicts the tens digit (with classes ${0,1}$), while the second predicts the ones digit (with classes ${0,\dots,9}$).
 During inference, the predicted digits are reconstructed to obtain the final class label.
-To improve robustness, we permute the class order $\sqrt{C}$ times, leveraging the efficiency of TabPFN v2 to predict across all permutations.
-The final prediction is obtained by ensembling these results.
-As shown in Figure 7, this approach achieves the second-best mean accuracy on 12 datasets with more than 10 classes while preserving computational efficiency.
-Figure 8: Average rank on 18 large-scale datasets.
-“*” indicates our extension.
+This strategy is also developed in [49], and we denote it as TabPFN v2-DPT.
+As the decimal encoding inherently introduces artificial correlations among classes — classes that share the same digit at a given position are grouped together, even if they are semantically unrelated.
+To mitigate this effect, our TabPFN v2 randomly permutes the class-to-digit mapping $\sqrt{C}$ times, leading to different groupings in each run, and the prediction results are ensembles to improve robustness.
+As shown in Figure 6 (middle), this approach achieves the second-best mean accuracy on 12 datasets with more than 10 classes while preserving computational efficiency.
+Figure 6: “*” indicates our extension.
+Left: Mean accuracy on 18 high-dimensional datasets.
+“-PCA” is another variant using PCA to reduce dimensions.
+Middle: Mean accuracy on 12 datasets with more than 10 classes.
+“-ECOC” denotes the multi-class ECOC strategy implemented by [29].
+Right: Average rank on 18 large-scale datasets.
 “-B” refers to the variant that randomly subsamples 10,000 training examples four times and aggregates their predictions.
@@
 All variants improve TabPFN v2.
+Detailed results can be found in Appendix E.
 7.3 Large-Scale Datasets
@@
 We extract their embeddings to form a new tabular dataset, on which a logistic regression classifier is trained to make predictions on the test set embeddings.
-This process is repeated four times, and the final predictions are aggregated, denoted as TabPFN v2-SQ.
+This process is repeated four times, and the final predictions are aggregated.
+We denote this version as TabPFN v2-SQ.
 We also investigate integrating TabPFN v2 with decision trees to handle large-scale tasks.
-Specifically, we first train a shallow decision tree, setting its maximum depth to 5 layers for datasets with fewer than 100,000 samples, and to 10 layers for larger datasets.
+We note that a similar strategy was mentioned in [29] to handle within-dataset heterogeneity for a drastically different purpose.
+Specifically, we sample 32 subsets from the original training set, each containing 60% of the original data (sampled without replacement).
+For each subset, we first train a shallow decision tree by setting the minimum number of samples required to split an internal node to 10,000.
 The decision tree partitions the training set into smaller, more manageable subsets.
-Then, we apply TabPFN v2 to each leaf node, treating the examples in each node as a separate training set for TabPFN v2.
-During inference, a test instance is first passed through the shallow decision tree to a leaf node and then predicted by the corresponding TabPFN v2 model.
-We denote this extension as TabPFN v2-DT.
-We note that a similar strategy was mentioned in Hollmann et al.
-(2025) but for a drastically difference purpose.
-Specifically, Hollmann et al.
-(2025) use decision trees to handle within-dataset heterogeneous for some highly challenging tabular datasets, while we use decision trees for scalability.
-Figure 8 shows the average rank results, including TabPFN v2-SQ and TabPFN v2-DT alongside variants using bagging and KMeans-based sampling.
-We observe that all variants improve upon the vanilla TabPFN v2 on large-scale datasets, with TabPFN v2-DT and TabPFN v2-SQ achieving the most significant improvement.
+During inference, a test instance is first passed through each of the shallow decision tree to a leaf node and then predicted by the corresponding TabPFN v2 model.
+The predictions from all 32 models are aggregated.
+We denote this extension as TabPFN v2-DF.
+Figure 6 (right) shows the average rank results, including TabPFN v2-SQ and TabPFN v2-DF alongside variants using bagging and KMeans-based sampling.
+We observe that all variants improve upon the vanilla TabPFN v2 on large-scale datasets, with TabPFN v2-DF and TabPFN v2-SQ achieving the most significant improvement.
 8 Conclusion
-Our analysis highlights TabPFN v2’s potential as a pre-trained foundation model for tabular data.
-By leveraging randomized feature tokens, it effectively handles dataset heterogeneity and excels in in-context learning, outperforming other models on small- to medium-scale datasets.
-Additionally, its ability to create nearly separable embeddings suggests its potential role as a feature encoder.
-To address TabPFN v2’s limitations on high-dimensional, large-scale, and many-category datasets, we propose post-hoc divide-and-conquer strategies that improve scalability, expanding applicability while preserving its core strengths.
-Appendix A Additional Related Work
-Learning with Tabular Data.
-Tabular data is prevalent across diverse fields, including healthcare, finance, and education (Kovalerchuk & Vityaev, 2005; Hyland et al., 2020; Romero & Ventura, 2010; Amatriain et al., 2010).
-Tree-based models, such as XGBoost (Chen & Guestrin, 2016), LightGBM (Ke et al., 2017), and CatBoost (Prokhorenkova et al., 2018), have long dominated this domain.
-However, recent advances in deep neural networks (DNNs) have demonstrated strong potential for tabular data (Borisov et al., 2024).
-Popular architectures like multi-layer perceptrons (Gorishniy et al., 2021; Kadra et al., 2021) and Transformers (Huang et al., 2020) have been adapted to tabular tasks, alongside custom architectures designed specifically for tabular data (Klambauer et al., 2017; Wang et al., 2021).
-Deep tabular methods can be broadly categorized into two types.
-The first type directly processes raw features (Holzmüller et al., 2024; Gorishniy et al., 2024; Ye et al., 2025), sometimes incorporating feature-specific encoding strategies (Gorishniy et al., 2022).
-The second type tokenizes features, transforming an example into a set of tokens (Song et al., 2019; Huang et al., 2020; Rubachev et al., 2022).
-Comprehensive benchmarks have been developed to evaluate these methods across diverse datasets (Grinsztajn et al., 2022; McElfresh et al., 2023; Ye et al., 2024b; Rubachev et al., 2025), highlighting the strengths and weaknesses of deep tabular models in various scenarios.
-Appendix B Evaluation Details
-Please refer Ye et al.
-(2024b) for details of the 300 small to medium datasets.
+We present a timely investigation into TabPFN v2, a groundbreaking foundation model for tabular tasks.
+Our analysis uncovers the core mechanism behind TabPFN v2’s strong performance across heterogeneous tabular datasets: it can infer attribute relationships on-the-fly—even from randomly initialized token inputs—without relying on pre-defined semantics or learning dataset-specific representations.
+We also demonstrate that TabPFN v2 can be repurposed as a powerful feature encoder, enabling broader applications such as data visualization and diagnostic analysis.
+To address its limitations in more complex data regimes, we introduce post-hoc divide-and-conquer strategies that extend TabPFN v2’s utility without requiring model re-training.
+Together, these contributions offer fresh insights into advancing the development and application of foundation models for tabular data.
+Appendix A Evaluation Details
+Experimental Compute Resources. All experiments were conducted using 4 NVIDIA RTX 6000 Ada GPUs and 2 Intel(R) Xeon(R) Platinum 8352V CPUs.
+Please refer [76, 46] for details of the 300 small to medium datasets.
 For high-dimensional datasets, we selected 18 datasets with more than 2000 features from the scikit-feature repository.
 Detailed statistics of high-dimensional datasets and large-scale datasets are reported in Table 2 and Table 3.
-We follow Ye et al.
-(2024b) and use different colors to represent various categories of methods in the result figures, ensuring clarity and easy comparison.
-In Figure 2 and Figure 3, we compare the following methods:
-  • Classical Methods ( ): The classical methods include Dummy, Logistic Regression (LR), K-Nearest Neighbors (KNN), Support Vector Machines (SVM), Naive Bayes, Linear Regression, and DNNR (Nader et al., 2022), which serve as basic baselines for classification and regression tasks.
-  • Tree-based Methods ( ): Tree-based methods such as Random Forest (Breiman, 2001), XGBoost (Chen & Guestrin, 2016), LightGBM (Ke et al., 2017), and CatBoost (Prokhorenkova et al., 2018) are known for their high performance on tabular data.
-  • MLP variants ( ): MLP variants, including vanilla MLP, MLP-PLR, Self-Normalizing Neural Networks (Klambauer et al., 2017), and Residual Network (Gorishniy et al., 2021), enhance the flexibility and generalization of traditional MLP architectures through advanced regularization and residual connections.
-  • Special Architectures ( ): Methods with specially designed architectures, such as DCNv2 (Wang et al., 2021), DANets (Chen et al., 2022), and TabCaps (Chen et al., 2023), focus on improving feature interaction and abstraction to capture complex relationships in tabular data.
-  • Token-based Methods ( ): Token-based methods like AutoInt (Song et al., 2019), TabTransformer (Huang et al., 2020), FT-Transformer (Gorishniy et al., 2021), and ExcelFormer (Chen et al., 2024) represent features as tokens, enabling models to capture higher-order interactions through attention mechanisms.
-  • Regularization-based Methods ( ): Regularization-based methods, including TANGOS (Jeffares et al., 2023), SwitchTab (Wu et al., 2024), and PTaRL (Ye et al., 2024a), aim to improve model generalization by incorporating regularization techniques during training to enhance the robustness of predictions.
-  • Tree-mimic Methods ( ): Tree-mimic methods, such as NODE (Popov et al., 2020), GrowNet (Badirli et al., 2020), and TabNet (Arik & Pfister, 2021), combine the interpretability of decision trees with the power of deep learning, employing attention mechanisms to select important features.
-  • Context-based Methods ( ): Context-based methods like TabR (Gorishniy et al., 2024) and ModernNCA (Ye et al., 2025) leverage contextual information from the training data to improve predictions by utilizing neighborhood-based and in-context learning strategies.
+The abbreviations of feature names shown in Figure 4 are explained in Tables 4 and 5.
+We follow [76] and use different colors to represent various categories of methods in the result figures, ensuring clarity and easy comparison.
+In Figure 1 (right) and Figure 3, we compare the following methods:
+  • Classical Methods ( ): The classical methods include Dummy, Logistic Regression (LR), K-Nearest Neighbors (KNN), Support Vector Machines (SVM), Naive Bayes, Linear Regression, and DNNR [53], which serve as basic baselines for classification and regression tasks.
+  • Tree-based Methods ( ): Tree-based methods such as Random Forest [8], XGBoost [12], LightGBM [39], and CatBoost [57] are known for their high performance on tabular data.
+  • MLP variants ( ): MLP variants, including vanilla MLP, MLP-PLR, Self-Normalizing Neural Networks [42], Residual Network [23], and TabM [20] enhance the flexibility and generalization of traditional MLP architectures through advanced regularization and residual connections.
+  • Special Architectures ( ): Methods with specially designed architectures, such as DCNv2 [69], DANets [10], and TabCaps [9], focus on improving feature interaction and abstraction to capture complex relationships in tabular data.
+  • Token-based Methods ( ): Token-based methods like AutoInt [65], TabTransformer [32], FT-Transformer [23], and ExcelFormer [11] represent features as tokens, enabling models to capture higher-order interactions through attention mechanisms.
+  • Regularization-based Methods ( ): Regularization-based methods, including TANGOS [36], SwitchTab [73], and PTaRL [79], aim to improve model generalization by incorporating regularization techniques during training to enhance the robustness of predictions.
+  • Tree-mimic Methods ( ): Tree-mimic methods, such as NODE [56], GrowNet [5], and TabNet [4], combine the interpretability of decision trees with the power of deep learning, employing attention mechanisms to select important features.
+  • Context-based Methods ( ): Context-based methods like TabR [22] and ModernNCA [77] leverage contextual information from the training data to improve predictions by utilizing neighborhood-based and in-context learning strategies.
 In addition to the aforementioned methods, for other experimental results, we will demonstrate the performance of TabPFN v2 and its variants, which are represented by emerald teal ( ), ensuring that their experimental effects are clearly distinguished from the other methods.
+Remark.
+The standard checkpoint released by TabPFN employs a feature grouping size of 2, which complicates the analysis of individual feature embeddings and inter-feature relationships.
+To facilitate such analysis, we use a modified checkpoint with group size=1 for the experiments in Figure 4, which is available at HuggingFace.
 Table 2: Dataset Information for High-Dimensional Data Experiments: A collection of 18 datasets with varying numbers of instances, features, and classes used in our high-dimensional experiments.
@@
 microsoft 1,200,192 136 - yahoo 709,877 699 -
-Appendix C Detailed Results
+Table 4: Explanation of the feature name abbreviations in the bank dataset used in Figure 4.
+Feature Description
+age Age of the client.
+bal Average yearly balance in euros.
+day Last contact day of the month.
+dur Last contact duration in seconds.
+camp Number of contacts performed during this campaign.
+pdays Number of days since the client was last contacted (999 means never contacted).
+prev Number of contacts performed before this campaign.
+job Job type of the client (e.g., admin., technician, unemployed).
+mar Marital status of the client.
+edu Education level of the client.
+def Whether the client has credit in default.
+house Whether the client has a housing loan.
+loan Whether the client has a personal loan.
+cont Communication type used for the last contact (e.g., cellular, telephone).
+mon Last contact month of the year.
+pout Outcome of the previous marketing campaign.
+Table 5: Explanation of the feature name abbreviations in the second dataset used in Figure 4.
+Feature Description
+state The U.S. state associated with the customer’s account.
+al The number of days or months the account has been active.
+pn The customer’s phone number (may serve as an identifier).
+nvmm The number of voicemail messages received.
+tdm Total number of minutes the customer talked during the day.
+tdc Total number of calls made during the day.
+tdc2 Total charges incurred during the day.
+tem Total number of minutes talked during the evening.
+tec Total number of evening calls made.
+tec2 Total charges incurred during the evening.
+tnm Total number of minutes talked at night.
+tnc Total number of night calls made.
+tnc2 Total charges incurred at night.
+tim Total number of minutes talked internationally.
+tic Total number of international calls made.
+tic2 Total charges for international calls.
+ac The area code of the customer’s phone number.
+ip Whether the customer has an international calling plan (yes/no).
+vmp Whether the customer has a voicemail plan (yes/no).
+ncs Number of times the customer contacted customer service.
+Appendix B Details of Comparison Baselines
+This section provides implementation details of the TabPFN v2 variants used in our analysis for high-dimensional, many-class, and large-scale tasks (cf. Figure 6).
+High-Dimensional Classification.
+To address high-dimensional tasks, we introduce a baseline variant, TabPFN v2-PCA, which incorporates dimensionality reduction.
+Specifically, TabPFN v2-PCA reduces the feature dimension to 500 using PCA to satisfy the input constraints of TabPFN v2.
+This process is repeated multiple times with different PCA projections, and the resulting predictions are aggregated via bagging to improve robustness.
+Many-Class Classification. We consider the following variants of TabPFN v2 to address the 10-class limit in classification tasks:
+  • TabPFN v2-ECOC: We use the implementation provided in the official TabPFN extensions repository, which applies Error-Correcting Output Codes (ECOC).
+  • TabPFN v2-DPT: We encode each class label as a $t$-digit decimal string and train a separate TabPFN v2 to predict each digit.
+For instance, a 15-class problem is decomposed into two subproblems: one for the tens digit (classes $\{0,1\}$) and one for the ones digit (classes $\{0,\dots,9\}$).
+The predicted digits are then decoded to recover the final class label.
+The same strategy is also proposed in [49].
+We implement TabPFN v2 based on TabPFN v2-DPT for efficiency.
+Specifically, TabPFN v2 permutes the class-to-digit mapping $\sqrt{C}$ times.
+For fair comparison, we also increase the number of ensembles in TabPFN v2-DPT to $\sqrt{C}$ per digit to match the total number of predictions.
+Large-Scale Tasks. As the size of training data increases, the benefit of pre-trained knowledge may diminish. We consider the following variants of TabPFN v2 to scale to larger datasets:
+  • TabPFN v2-DT: A shallow decision tree is trained with a minimum split size of 10,000.
+The tree partitions the dataset into smaller subsets, and a separate TabPFN v2 model is applied to each leaf node.
+At inference, a test instance is routed through the tree to a corresponding leaf, where it is predicted by the respective TabPFN v2 model.
+  • TabPFN v2-B: A bagging-based variant that randomly samples 10,000 training examples four times and aggregates their predictions.
+  • TabPFN v2-K: Selects a representative subset of 10,000 training examples based on proximity to KMeans-derived prototypes.
+Our TabPFN v2 -DF is an extension of TabPFN v2-DT to a forest-based ensemble.
+Specifically, we sample 32 subsets from the original training data, each containing 60% of the samples (without replacement), and train a separate TabPFN v2-DT model on each subset.
+During inference, predictions from all 32 models are aggregated—e.g., by majority voting or averaging—depending on the task type.
+Appendix C Additional Feature Extraction Results
+In this section, we provide further results to validate the effectiveness of our leave-one-fold-out strategy for feature extraction with TabPFN v2.
+We also compare this supervised approach to two unsupervised embedding extraction methods.
+C.1 Additional Results for Figure 5
+Figure 7 visualizes instance embeddings extracted using our leave-one-fold-out strategy on two datasets: website_phishing (first row, three classes) and KDD (second row, binary).
+Training and test examples are shown as darker crosses and lighter circles, respectively, and colored by class.
+These visualizations support our claim that TabPFN v2, when combined with our proposed strategy, can serve as an effective feature encoder.
+(a) Raw Feature
+(b) Vanilla
+(c) Layer 3
+(d) Layer 6
+(e) Layer 9
+(f) Layer 12
+Figure 7: Visualization of the extracted instance features from two datasets: website_phishing (first row, three classes) and KDD (second row, binary).
+Colors indicate classes, where Darker crosses and lighter circles denote training and test examples, respectively.
+(a) shows the raw input features (e.g., ${\bm{x}}_{i}$), while (b) presents embeddings from the vanilla strategy.
+(c)-(f) display embeddings produced by our method at different layers.
+Classification accuracy is reported by training a linear logistic regression model on the training embeddings and evaluating on the test set.
+Table 6: Average rank (lower is better) of TabPFN v2 and a linear classifier trained on the extracted embeddings across 29 classification datasets.
+In addition to the supervised feature extraction strategy considered in Section 6 (including the vanilla one, our leave-one-fold-out, and the version based on the combined features), we compare with two unsupervised embedding extraction approaches by appending a column of dummy labels with zero values and permuting each column as labels, respectively.
+$\downarrow$ TabPFN v2 Vanilla Dummy Permute Ours Combined
+Rank 2.66 5.72 4.90 3.69 2.16 1.88
+C.2 Comparison Between Supervised and Unsupervised Feature Extraction
+Our leave-one-fold-out strategy explicitly incorporates label information and accounts for the distinct roles of training and test instances.
+To further understand the nature of the extracted embeddings, we compare this supervised strategy to two unsupervised alternatives provided by the TabPFN extensions repository.
+Let $X\in\mathbb{R}^{n\times d}$ denote a dataset with $n$ samples and $d$ features:
+  • Unsupervised-Dummy: Each sample is paired with a constant pseudo-target $\mathbf{y}=\mathbf{0}\in\mathbb{R}^{n}$, forming a regression task.
+The embedding $E_{\text{D}}\in\mathbb{R}^{n\times h}$ is obtained by extracting the output tokens of the TabPFN regressor.
+  • Unsupervised-Permute: For each feature $j\in\{1,\ldots,d\}$, we treat $\mathbf{x}^{(j)}=X_{:,j}$ as a pseudo-target and use the remaining features $X^{(-j)}$ as input.
+Depending on the type of $\mathbf{x}^{(j)}$, TabPFN is applied in classification or regression mode to obtain $E^{(j)}\in\mathbb{R}^{n\times h}$.
+These embeddings are concatenated to a high-dimensional form:
+    $E_{\text{P}}=\text{concat}(E^{(1)},E^{(2)},\ldots,E^{(d)})\in\mathbb{R}^{n\times(d\cdot h)}.$
+We compare these unsupervised methods with our supervised leave-one-fold-out strategy in Table 6.
+Overall, unsupervised approaches underperform compared to the supervised ones and fail to recover the classification ability of TabPFN v2.
+This performance gap arises because label information is introduced only post hoc via a linear classifier rather than during embedding extraction.
+Among the unsupervised methods, the permutation-based approach performs better, likely due to its ability to encode attribute-specific structure.
+Figure 8 presents a visual comparison of embeddings produced by the three methods using the same color and marker scheme as in Figure 7.
+Since unsupervised methods lack label supervision during embedding generation, their embeddings tend to scatter broadly without forming well-separated clusters by class.
+These results further highlight the contrasting goals of supervised and unsupervised strategies—class separation versus feature distribution coverage, respectively.
+(a) Raw Feature
+(b) Vanilla
+(c) Dummy
+(d) Permute
+(e) Ours
+Figure 8: Comparison between unsupervised and supervised (ours) feature extraction.
+Visualization of extracted embeddings for four datasets: churn (first row, two classes), bank (second row, two classes), KDD (third row, two classes), and website_phishing (fourth row, three classes).
+We use crosses to denote training examples and circles to denote test examples.
+(a) shows the raw features, while (b) presents the embeddings extracted using the vanilla strategy.
+(c) and (d) refer to unsupervised embedding extraction approaches by appending a column of dummy labels with zero values and permuting each column as labels, respectively.
+(e) depicts the embeddings obtained using our proposed methods.
+The accuracy value is calculated by training a linear model (logistic regression) over the extracted embeddings on the training set and predicting on the test set.
+Appendix D Influence of Key Modules and Meta-Feature Analysis
+We investigate the influence of two key components in TabPFN v2, i.e., the feature engineering that pre-processes the raw features of a given tabular dataset and the post-hoc ensembleing.
+In addition, we analyze the conditions under which TabPFN v2 performs well or poorly through a meta-feature-based classification analysis.
+Feature engineering.
+TabPFN v2 pre-processes the features of a given tabular dataset with various strategies, such as quantile, category shuffling, SVD, and power transform.
+Specifically, we examine the effects of adding fingerprint features (add_fingerprint_feature) and polynomial features (polynomial_features) to the raw tabular data.
+The results indicate that TabPFN v2 performs well even without the use of these engineered features, suggesting that, for the benchmark datasets of [76], these specific feature engineering techniques do not provide a significant improvement.
+This finding highlights the robustness of TabPFN v2 and its ability to handle raw features effectively, without the need for extensive pre-processing or feature construction.
+We show the influence of this step in Figure 9.
+Figure 9: Scatter plot comparing the normalized Accuracy/R2 scores.
+The x-axis represents the normalized Accuracy/R2 scores without Feature Engineering, while the y-axis represents the normalized Accuracy/R2 scores with Feature Engineering.
+The red dashed line ($y=x$) serves as a reference, indicating equal performance.
+Model ensemble.
+Post hoc ensembling (PHE) involves applying TabPFN v2 to the datasets multiple times with different perturbations and aggregating the predictions of these base models at different temperatures.
+We show the change of performance of TabPFN v2 w.r.t.
+the number of ensemble numbers (i.e., the number of base models) in Figure 10.
+On the benchmark of [76], we observe that, overall, ensemble methods improve performance, with larger ensemble sizes yielding better results.
+However, we also note that even without ensembling, TabPFN v2 performs exceptionally well, and the relative performance gain from ensembling is limited.
+This suggests that while ensembling can provide further improvements, the base TabPFN v2 model is already highly effective on its own.
+The equivariant property described in [3] provides insight into this phenomenon.
+Since TabPFN v2 introduces random tokens to handle heterogeneous features, the model becomes less sensitive to the arbitrary ordering of features, effectively enforcing equivariance in this aspect.
+As a result, the benefits of ensembling through feature order permutations are less pronounced compared to TabPFN v1.
+Figure 10: Box plot of relative performance improvements of TabPFN v2 with post hoc ensembling (PHE) across different ensemble sizes (2, 4, 8, and 16 base models).
+The relative improvement is calculated as the performance gain over the non-ensemble model, where higher values indicate stronger performance.
+The box plots show the median, interquartile range (IQR), and outliers for each ensemble size.
+Meta-Feature Analysis of TabPFN v2 Performance.
+To better understand the conditions under which TabPFN v2 performs well or poorly, we conducted a meta-learning-based classification analysis using 300 datasets.
+Specifically, we used the average rank of TabPFN v2 across datasets as a performance indicator.
+The threshold for classification was set at the mean rank, 6.31.
+Datasets where TabPFN v2 achieved a rank lower than or equal to 6.31 were labeled as “Good”, while those with a higher rank were labeled as “Bad”.
+We extracted meta-features from each dataset and used them to train a decision tree classifier, aiming to distinguish between the “Good” and “Bad” categories.
+All the meta-features utilized in this task are detailed in Table 7, accompanied by their respective explanations.
+A visual depiction of a simple depth-3 decision tree is shown in Figure 11, and the tree reveals key factors that influence the effectiveness of TabPFN v2.
+The decision tree visualizes how to predict whether TabPFN v2 performs well (“Good”) or poorly (“Bad”) on a given dataset, based on dataset meta-features: The root node splits on the number of instances (nr_inst), indicating that TabPFN v2 tends to perform better on datasets with fewer than 24,350 samples.
+For these smaller datasets, the left subtree further splits on the mean joint entropy (joint_ent.mean), where higher values (greater than 3.028) are associated with improved performance.
+For datasets with lower mean joint entropy ($\leq$ 3.028), TabPFN v2 also tends to perform well when the number of rows is relatively small ($\leq$ 4,862).
+In contrast, the right subtree, which represents larger datasets, reveals that a low standard deviation of the interquartile range (iq_range.std) across features ($\leq$ 0.657) is linked to poorer model performance.
+Figure 11: Decision tree for predicting TabPFN v2 performance (Good vs.
+Bad) based on dataset characteristics, constructed from experiments on 300 datasets.
+The tree splits on meta-features such as number of instances (nr_inst), joint entropy (joint_ent.mean), number of numerical features (nr_num), distance between minority and majority classes’ center of mass (gravity) and interquartile range (iq_range) statistics.
+Each split is chosen to maximize information gain.
+The leaf nodes indicate the predicted performance class, the Gini impurity, and class distribution.
+This tree provides insights into the types of datasets where TabPFN v2 is expected to perform well.
+Table 7: Meta-features used in the meta-feature analysis of TabPFN v2 performance.
+Meta-Feature Explanation
+attr_conc The concentration coef. of each pair of distinct attributes.
+class_conc The concentration coefficient between each attribute and class.
+class_ent The target attribute Shannon’s entropy.
+inst_to_attr The ratio between the number of instances and attributes.
+mean The mean value of each attribute.
+sd The standard deviation of each attribute.
+var The variance of each attribute.
+range The range (max - min) of each attribute.
+iq_range The interquartile range (IQR) of each attribute.
+nr_attr The total number of attributes.
+sparsity The (possibly normalized) sparsity metric for each attribute.
+t_mean The trimmed mean of each attribute.
+nr_bin The number of binary attributes.
+nr_cat The number of categorical attributes.
+nr_num The number of numeric features.
+nr_norm The number of attributes normally distributed based in a given method.
+nr_cor_attr The number of distinct highly correlated pair of attributes.
+gravity The distance between minority and majority classes’ center of mass.
+nr_class The number of distinct classes.
+joint_ent The joint entropy between each attribute and class.
+attr_ent Shannon’s entropy for each predictive attribute.
+cov The absolute value of the covariance of distinct dataset attribute pairs.
+eigenvalues The eigenvalues of covariance matrix from dataset.
+eq_num_attr The number of attributes equivalent for a predictive task.
+max The maximum value from each attribute.
+min The minimum value from each attribute.
+median The median value from each attribute.
+freq_class The relative frequency of each distinct class.
+mad The Median Absolute Deviation (MAD) adjusted by a factor.
+mad The Median Absolute Deviation (MAD) adjusted by a factor.
+mut_inf The mutual information between each attribute and target.
+nr_inst The number of instances (rows) in the dataset.
+nr_outliers The number of attributes with at least one outlier value.
+ns_ratio The noisiness of attributes.
+imblance_ratio The ratio of the number of instances in the minority to the majority class.
+attr_to_inst The ratio between the number of attributes.
+Appendix E Detailed Results
 We list the detailed results of TabPFN v2 and our extensions on various benchmarks.
@@
 Notably, we excluded 27 datasets from these results in LABEL:tab:main_results, as they were used by TabPFN v2 to select the best checkpoint.
-These excluded datasets, which are not shown in Figure 2 and Figure 3, include:
+These excluded datasets, which are not shown in Figure 1 (right) and Figure 3, include:
     • (1)
@@
 The table provides the mean accuracy of models like KNN, TabPFN-v2*, XGBoost (XGB), CatBoost (CatB), Random Forest (RForest), ModernNCA (MNCA), MLP, Logistic Regression (LogReg), and RealMLP, showcasing how they perform on multi-class tasks with a larger number of classes.
-Additionally, we compare PFN-v2-ECOC, a multi-class classification solution provided by Hollmann et al.
-(2025).
-This method extends TabPFN-v2 by leveraging Error-Correcting Output Codes (ECOC) to enhance multi-class classification performance [^2] https://github.com/PriorLabs/tabpfn-community/blob/main/src/tabpfn_extensions/many_class/many_class_classifier.py.
+Additionally, we compare PFN-v2-ECOC, a multi-class classification solution provided by [29].
+This method extends TabPFN-v2 by leveraging Error-Correcting Output Codes (ECOC) to enhance multi-class classification performance.
+[^3] https://github.com/PriorLabs/tabpfn-community/blob/main/src/tabpfn_extensions/many_class/many_class_classifier.py
   • In Table 11, we compare the performance of various models on 18 large-scale datasets.
@@
 This illustrates the models’ performance on large-scale datasets.
-  • We present a comparison of FT-T (Gorishniy et al., 2021) and MNCA (Ye et al., 2025) with and without randomized feature tokens on tiny benchmark datasets from Ye et al.
-(2024b) in Table 9.
-The first two columns represent FT-T and MNCA with randomized feature tokens, while the last two show the original FT-T and MNCA methods.
-The table demonstrates the effect of randomized feature tokens on model performance, with accuracy for classification tasks and RMSE for regression tasks.
-FT-T and MNCA implementations, including their hyperparameter search space, follow the approach described in Liu et al.
-(2024), and the results were obtained after 100 rounds of hyperparameter tuning.
   • In Table 12, we show the performance of TabPFN v2 and the extracted feature embeddings across 29 classification datasets.
@@
 Dataset MNCA MLP KNN RealMLP XGB RForest LogReg PFN-v2 v2-pca v2* ProtoG CatB
-CLL_SUB_111 62.90 72.46 57.39 70.43 73.04 70.14 73.91 70.14 57.68 73.17 65.51 71.59
-BASEHOCK 96.31 97.01 71.88 97.46 95.29 96.73 96.99 69.09 97.41 97.47 96.32 95.87
-Prostate_GE 81.27 86.98 80.00 87.94 89.52 87.94 91.43 95.24 88.57 94.60 84.13 94.92
-PCMAC 88.21 88.53 66.48 90.15 91.64 92.20 87.15 92.70 90.76 86.86 88.21 92.01
+CLL_SUB_111 62.90 72.46 57.39 70.43 73.04 70.14 73.91 70.14 57.68 71.59 65.51 71.59
+BASEHOCK 96.31 97.01 71.88 97.46 95.29 96.73 96.99 69.09 97.41 97.36 96.32 95.87
+Prostate_GE 81.27 86.98 80.00 87.94 89.52 87.94 91.43 95.24 88.57 94.29 84.13 94.92
+PCMAC 88.21 88.53 66.48 90.15 91.64 92.20 87.15 92.70 90.76 90.14 88.21 92.01
 GLI_85 81.57 85.49 76.47 89.80 82.35 83.92 90.59 80.39 86.27 92.55 81.96 80.78
-RELATHE 88.18 90.54 75.03 90.23 87.11 87.30 90.49 86.36 87.65 86.81 89.92 90.35
-SMK_CAN_187 63.51 66.84 69.47 69.82 66.49 70.70 72.11 71.05 71.75 73.28 70.71 71.40
-warpPIE10P 98.41 99.05 92.38 100.0 94.92 98.57 100.0 100.0 100.0 99.84 97.79 98.89
-leukemia 90.22 95.11 86.67 94.67 97.78 92.00 96.00 92.44 93.33 95.33 94.00 94.22
-orlraws10P 97.67 98.33 92.00 99.00 84.33 99.00 99.00 92.00 99.33 98.33 92.67 99.00
-GLIOMA 58.00 60.67 68.00 67.33 66.67 64.00 64.00 62.67 69.33 68.00 69.91 66.67
-warpAR10P 83.08 85.64 53.08 97.44 81.28 87.18 97.69 90.77 95.38 92.05 90.04 87.44
-TOX_171 76.00 88.19 70.86 90.48 78.10 78.67 90.29 80.95 82.48 86.86 85.52 83.05
-lung 91.54 95.45 93.66 95.28 93.66 92.68 95.12 95.28 93.50 95.45 95.43 93.01
-ALLAML 87.56 95.56 81.33 96.89 96.00 96.44 92.00 92.89 93.78 92.44 91.14 94.67
-colon 78.46 78.97 76.92 83.08 74.87 82.56 86.15 81.54 78.46 80.00 78.46 77.95
-gisette 97.21 97.57 95.04 97.86 97.55 96.82 97.51 97.35 97.26 97.37 97.18 97.78
-arcene 81.67 85.50 84.50 81.00 75.00 86.83 88.00 83.67 88.33 90.67 85.33 85.00
-Mean 82.86 87.11 77.29 88.83 84.76 86.87 89.36 85.25 87.29 88.95 86.37 87.48
+RELATHE 88.18 90.54 75.03 90.23 87.11 87.30 90.49 86.36 87.65 89.95 89.92 90.35
+SMK_CAN_187 63.51 66.84 69.47 69.82 66.49 70.70 72.11 71.05 71.75 72.10 70.71 71.40
+warpPIE10P 98.41 99.05 92.38 100.0 94.92 98.57 100.0 100.0 100.0 100.0 97.79 98.89
+leukemia 90.22 95.11 86.67 94.67 97.78 92.00 96.00 92.44 93.33 96.00 94.00 94.22
+orlraws10P 97.67 98.33 92.00 99.00 84.33 99.00 99.00 92.00 99.33 99.67 92.67 99.00
+GLIOMA 58.00 60.67 68.00 67.33 66.67 64.00 64.00 62.67 69.33 68.67 69.91 66.67
+warpAR10P 83.08 85.64 53.08 97.44 81.28 87.18 97.69 90.77 95.38 96.67 90.04 87.44
+TOX_171 76.00 88.19 70.86 90.48 78.10 78.67 90.29 80.95 82.48 87.24 85.52 83.05
+lung 91.54 95.45 93.66 95.28 93.66 92.68 95.12 95.28 93.50 95.61 95.43 93.01
+ALLAML 87.56 95.56 81.33 96.89 96.00 96.44 92.00 92.89 93.78 94.67 91.14 94.67
+colon 78.46 78.97 76.92 83.08 74.87 82.56 86.15 81.54 78.46 79.49 78.46 77.95
+gisette 97.21 97.57 95.04 97.86 97.55 96.82 97.51 97.35 97.26 97.23 97.18 97.78
+arcene 81.67 85.50 84.50 81.00 75.00 86.83 88.00 83.67 88.33 92.00 85.33 85.00
+Mean 82.86 87.11 77.29 88.83 84.76 86.87 89.36 85.25 87.29 89.73 86.37 87.48
 Table 10: Performance of various models on 12 multi-class classification tasks with more than 10 classes.
@@
 The performance is evaluated on 12 multi-class datasets with more than 10 classes, with accuracy values presented for each model on the respective datasets.
-Dataset KNN PFN-v2* PFN-v2-ECOC XGB CatB RForest MNCA MLP LogReg RealMLP
-100-plants-texture 79.69 90.94 84.92 77.06 89.73 82.65 80.52 83.92 86.88 88.35
-100-plants-margin 77.50 88.56 79.40 74.25 84.06 82.79 77.60 80.44 79.69 83.58
-100-plants-shape 60.31 79.52 63.38 56.15 65.19 64.33 70.10 47.33 65.94 72.08
-UJI_Pen_Characters 36.26 45.71 44.20 30.35 38.88 34.24 44.03 37.75 19.41 46.37
-texture 98.45 100.0 100.0 98.55 99.13 96.76 99.68 99.40 99.64 99.95
-letter 94.90 97.57 97.78 96.26 96.75 91.56 97.96 96.40 75.80 98.31
-walking-activity 60.29 61.22 61.92 65.06 64.92 61.74 64.85 60.64 27.02 65.13
-helena 28.94 33.31 19.20 32.42 37.90 33.91 36.58 37.91 33.40 38.55
-internet_usage 30.17 54.34 50.86 51.08 37.90 33.91 52.09 43.00 37.73 52.23
-kropt 71.22 77.96 77.11 86.95 79.26 71.77 78.27 64.45 28.08 92.03
-kr-vs-k 70.78 78.46 76.29 87.26 74.81 71.60 76.83 65.03 28.03 91.85
-ASP-POTASSCO 34.75 43.50 45.27 42.24 41.08 42.86 37.45 29.63 35.14 41.70
-Mean 61.94 70.93 66.69 66.47 67.47 64.01 68.00 62.16 51.40 72.51
+Dataset KNN PFN-v2* PFN-v2-DPT PFN-v2-ECOC XGB CatB RForest MNCA MLP LogReg RealMLP
+100-plants-texture 79.69 90.94 82.67 84.92 77.06 89.73 82.65 80.52 83.92 86.88 88.35
+100-plants-margin 77.50 88.56 81.94 79.40 74.25 84.06 82.79 77.60 80.44 79.69 83.58
+100-plants-shape 60.31 79.52 72.15 63.38 56.15 65.19 64.33 70.10 47.33 65.94 72.08
+UJI_Pen_Characters 36.26 45.71 33.38 44.20 30.35 38.88 34.24 44.03 37.75 19.41 46.37
+texture 98.45 100.0 99.98 100.0 98.55 99.13 96.76 99.68 99.40 99.64 99.95
+letter 94.90 97.57 96.69 97.78 96.26 96.75 91.56 97.96 96.40 75.80 98.31
+walking-activity 60.29 61.22 57.28 61.92 65.06 64.92 61.74 64.85 60.64 27.02 65.13
+helena 28.94 33.31 28.54 19.20 32.42 37.90 33.91 36.58 37.91 33.40 38.55
+internet_usage 30.17 54.34 50.51 50.86 51.08 37.90 33.91 52.09 43.00 37.73 52.23
+kropt 71.22 77.96 71.44 77.11 86.95 79.26 71.77 78.27 64.45 28.08 92.03
+kr-vs-k 70.78 78.46 71.54 76.29 87.26 74.81 71.60 76.83 65.03 28.03 91.85
+ASP-POTASSCO 34.75 43.50 41.88 45.27 42.24 41.08 42.86 37.45 29.63 35.14 41.70
+Mean 61.94 70.93 65.67 66.69 66.47 67.47 64.01 68.00 62.16 51.40 72.51
 Table 11: Performance of various models on 18 large-scale datasets.
-The results show the mean accuracy/RMSE of different models, including MLP, Logistic Regression/Linear Regression (LR), KNN, XGBoost (XGB), Random Forest (RForest), CatBoost (CatB), ModernNCA (MNCA), RealMLP, and various versions of TabPFN v2: original TabPFN v2 (PFNv2), TabPFN v2 with K-means (PFNv2-K), TabPFN v2 with Bagging (PFNv2-B), and PFNv2* (TabPFNv2*).
-Dataset MLP LR KNN XGB RForest CatB MNCA RealMLP PFNv2 PFNv2-K PFNv2-B PFNv2* PFNv2-DT
-BNG(credit-a) 90.07 85.98 87.41 90.21 89.25 91.13 89.98 90.91 89.55 89.01 89.66 89.89 89.96
-CDC_Indicators 86.79 86.55 86.39 86.76 86.60 86.78 86.76 86.76 86.65 86.68 86.69 86.74 86.22
-Higgs 75.53 64.29 65.16 73.33 71.87 74.81 73.28 75.36 71.64 71.56 72.01 72.13 72.44
-Smoking_signal 73.90 72.53 72.36 73.87 73.08 73.99 73.63 74.00 73.47 73.37 73.55 73.69 72.94
-nomao 96.19 94.59 95.20 96.92 96.07 97.03 96.68 96.37 96.08 96.29 96.12 96.18 96.51
-sf-police-incidents 87.84 87.84 85.87 87.68 87.84 87.87 - 87.84 87.84 87.84 87.84 87.84 87.84
-Data_Crowdfunding 96.48 67.04 93.70 96.89 95.29 96.81 96.53 96.71 94.59 91.81 94.96 95.07 96.93
-Fashion-MNIST 89.54 85.69 86.00 90.03 86.57 90.24 89.36 90.25 68.40 82.82 83.89 86.26 78.58
-covertype 94.01 72.54 92.76 96.30 78.30 90.77 97.31 97.38 83.54 82.95 84.16 86.85 96.89
-jannis 71.99 64.60 65.67 71.83 69.19 72.26 72.57 73.00 70.24 70.26 70.59 71.31 71.90
-poker-hand 99.99 50.12 54.01 99.51 64.63 97.69 76.31 99.88 41.97 38.86 36.80 54.12 85.60
-volkert 69.85 58.75 67.41 69.74 62.71 70.88 77.18 73.76 62.82 62.15 62.81 64.84 68.53
-Airlines_DepDelay (×$\times 10^{1}$) 2.905 2.933 3.170 2.891 2.907 2.881 - 2.482 2.937 2.933 2.937 2.915 2.906
-Wave_Energy_Farm (×$\times 10^{3}$) 8.199 13.19 32.29 6.917 7.294 7.173 6.148 59.05 7.214 8.375 7.063 10.506 6.489
-UJIndoorLoc (×$\times 10^{0}$) 9.958 $\infty$ 9.004 10.47 23.19 9.139 5.990 65.34 66.49 7.825 7.435 9.538 7.024
-blogfeedback (×$\times 10^{1}$ ) 2.387 $\infty$ 2.410 2.093 2.026 2.044 1.953 2.105 3.073 2.687 2.700 2.014 2.011
-microsoft (×$\times 10^{-1}$ ) 7.577 7.782 8.284 7.514 7.566 7.453 7.573 5.077 7.735 7.981 7.720 7.612 7.952
-yahoo (×$\times 10^{-1}$ ) 7.692 7.997 8.504 7.629 - 7.514 - 5.671 8.148 8.332 8.132 7.961 8.118
-Table 9: Performance comparison of FT-T and MNCA with and without randomized feature tokens across tiny benchmark datasets of Ye et al.
-(2024b).
-The first two columns (FT-T and MNCA) show results using randomized tokens, while the last two columns (FT-T and MNCA) represent the original methods.
-All values are the mean of 15 runs with different random seeds.
-For classification tasks, the performance is measured by accuracy, while for regression tasks, the metric used is Root Mean Squared Error (RMSE).
-The table demonstrates the effect of randomized feature tokens on model performance in both types of tasks.
-Dataset FT-T* MNCA* FT-T MNCA
-BNG(breast-w) 98.62 98.66 98.60 98.59
-BNG(cmc) 58.35 58.39 58.50 58.56
-BNG(tic-tac-toe) 80.92 81.10 81.31 81.08
-Cardiovascular-Disease-dataset 73.35 73.29 73.22 73.20
-FOREX_audchf-day-High 52.90 72.93 51.79 75.44
-FOREX_audsgd-hour-High 69.53 66.86 61.84 71.09
-FOREX_cadjpy-hour-High 69.62 70.99 68.80 71.18
-Gender_Gap_in_Spanish_WP 59.16 59.16 59.86 59.25
-KDD 80.15 80.12 80.28 79.58
-VulNoneVul 98.93 98.81 98.93 98.79
-baseball 93.56 93.71 93.76 92.74
-credit 75.83 75.39 75.88 72.98
-dis 98.29 98.31 98.76 98.68
-eye_movements_bin 59.01 61.31 59.45 90.88
-jungle_chess_2pcs_raw_endgame_complete 97.11 99.67 97.61 99.50
-law-school-admission-bianry 99.00 100.0 98.02 100.0
-mfeat-fourier 86.28 85.78 85.23 86.48
-online_shoppers 89.70 89.65 90.19 90.53
-page-blocks 96.89 97.23 96.71 96.68
-pc3 88.54 89.16 89.78 88.75
-pendigits 99.20 99.39 99.35 99.40
-rl 66.99 79.74 71.96 83.75
-satimage 89.61 90.77 89.59 91.04
-segment 91.49 91.40 91.37 92.77
-sylvine 94.48 94.22 94.71 95.88
-taiwanese_bankruptcy_prediction 96.90 96.68 96.74 96.74
-waveform-5000 85.89 86.02 86.11 86.24
-website_phishing 89.45 89.94 90.14 87.53
-wine-quality-white 54.98 64.41 55.17 63.27
-Ailerons 0.000157 0.000158 0.000155 0.000158
-CookbookReviews 1.560657 1.491684 1.520183 1.486840
-IEEE80211aa-GATS 0.030513 0.027608 0.029589 0.026295
-Large-scale_Wave_Energy_Farm_Sydney_49 4504.285 4369.826 4369.902 47399.63
-Superconductivty 10.53080 10.39726 10.64536 10.37837
-archive2 388.8410 379.5506 390.6574 361.8616
-bank8FM 0.028792 0.029047 0.028635 0.028826
-communities_and_crime 0.143486 0.139045 0.136859 0.138984
-fried 1.013058 1.013431 1.008568 1.010952
-healthcare_insurance_expenses 4935.203 4743.384 4615.580 4698.968
-house_16H_reg 31647.43 31333.32 30671.63 31017.80
-kin8nm 0.068684 0.070110 0.067355 0.069678
-mv 0.027143 0.019089 0.030948 0.027385
+The results show the mean accuracy/RMSE of different models, including MLP, Logistic Regression/Linear Regression (LR), KNN, XGBoost (XGB), Random Forest (RForest), CatBoost (CatB), ModernNCA (MNCA), RealMLP, and various versions of TabPFN v2: original TabPFN v2 (PFNv2), TabPFN v2 with K-means (PFNv2-K), TabPFN v2 with Bagging (PFNv2-B), PFNv2* (TabPFNv2*), PFNv2-DT (TabPFN-DT), and PFNv2-DF (TabPFN-DF).
+Dataset MLP LR KNN XGB RForest CatB MNCA RealMLP PFNv2 PFNv2-K PFNv2-B PFNv2* PFNv2-DT PFNv2-DF
+BNG(credit-a) 90.07 85.98 87.41 90.21 89.25 91.13 89.98 90.91 89.55 89.01 89.66 89.89 90.45 90.43
+CDC_Indicators 86.79 86.55 86.39 86.76 86.60 86.78 86.76 86.76 86.65 86.68 86.69 86.74 86.75 86.70
+Higgs 75.53 64.29 65.16 73.33 71.87 74.81 73.28 75.36 71.64 71.56 72.01 72.13 73.53 73.62
+Smoking_signal 73.90 72.53 72.36 73.87 73.08 73.99 73.63 74.00 73.47 73.37 73.55 73.69 73.74 73.84
+nomao 96.19 94.59 95.20 96.92 96.07 97.03 96.68 96.37 96.08 96.29 96.12 96.18 96.75 96.34
+sf-police-incidents 87.84 87.84 85.87 87.68 87.84 87.87 - 87.84 87.84 87.84 87.84 87.84 87.84 87.84
+Data_Crowdfunding 96.48 67.04 93.70 96.89 95.29 96.81 96.53 96.71 94.59 91.81 94.96 95.07 96.90 96.83
+Fashion-MNIST 89.54 85.69 86.00 90.03 86.57 90.24 89.36 90.25 68.40 82.82 83.89 86.26 78.91 78.82
+covertype 94.01 72.54 92.76 96.30 78.30 90.77 97.31 97.38 83.54 82.95 84.16 86.85 97.38 97.44
+jannis 71.99 64.60 65.67 71.83 69.19 72.26 72.57 73.00 70.24 70.26 70.59 71.31 72.57 72.50
+poker-hand 99.99 50.12 54.01 99.51 64.63 97.69 76.31 99.88 41.97 38.86 36.80 54.12 91.13 92.33
+volkert 69.85 58.75 67.41 69.74 62.71 70.88 77.18 73.76 62.82 62.15 62.81 64.84 68.66 67.76
+Airlines_DepDelay (×$\times 10^{1}$) 2.905 2.933 3.170 2.891 2.907 2.881 - 2.482 2.937 2.933 2.937 2.915 2.900 2.897
+Wave_Energy_Farm (×$\times 10^{3}$) 8.199 13.19 32.29 6.917 7.294 7.173 6.148 59.05 7.214 8.375 7.063 10.506 6.616 6.785
+UJIndoorLoc (×$\times 10^{0}$) 9.958 $\infty$ 9.004 10.47 23.19 9.139 5.990 65.34 66.49 7.825 7.435 9.538 14.404 7.472
+blogfeedback (×$\times 10^{1}$ ) 2.387 $\infty$ 2.410 2.093 2.026 2.044 1.953 2.105 3.073 2.687 2.700 2.014 1.914 1.944
+microsoft (×$\times 10^{-1}$ ) 7.577 7.782 8.284 7.514 7.566 7.453 7.573 5.077 7.735 7.981 7.720 7.612 7.944 7.728
+yahoo (×$\times 10^{-1}$ ) 7.692 7.997 8.504 7.629 - 7.514 - 5.671 8.148 8.332 8.132 7.961 16.409 8.069
 Table 12: Performance of TabPFN v2 and the extracted feature embeddings across 29 classification datasets.
@@
 “Vanilla” refers to the embeddings extracted using the vanilla strategy, which utilizes only the 12th layer of the transformer.
-PFN-v2 Vanilla layer-6 layer-9 layer-12 combined selected layers
-FOREX_audchf-day-High 77.38 72.48 68.39 73.57 74.11 77.11 (5, 9, 11)
-taiwanese_bankruptcy_prediction 96.99 96.70 97.14 96.77 97.07 97.14 (6)
-rl 85.51 85.41 66.90 69.52 86.72 87.53 (11, 12)
-pc3 89.46 89.78 90.10 88.82 88.82 88.82 (8)
-eye_movements_bin 61.83 62.75 59.72 59.40 62.16 62.16 (6, 9, 12)
-BNG(breast-w) 98.43 98.51 98.34 98.46 98.67 98.51 (6, 9)
-FOREX_cadjpy-hour-High 69.53 70.77 62.12 64.87 70.66 70.88 (4, 5, 6)
-dis 99.34 99.07 98.41 98.28 99.34 99.47 (4, 5, 6)
-sylvine 97.46 97.27 92.49 93.95 97.27 96.49 (1, 11)
-BNG(tic-tac-toe) 78.04 78.75 73.96 73.71 78.75 79.03 (5, 10, 12)
-online_shoppers 90.59 90.11 90.02 90.11 90.63 90.02 (8)
-Cardiovascular-Disease-dataset 72.84 73.08 72.96 73.06 73.14 73.09 (5, 8, 12)
-credit 78.04 77.06 77.62 77.80 77.95 77.59 (4, 6, 9)
-FOREX_audsgd-hour-High 67.26 69.30 57.24 61.06 69.62 70.41 (7, 10, 12)
-waveform-5000 86.00 85.30 85.60 85.60 86.40 86.90 (1, 6, 11)
-jungle_chess 85.65 85.30 78.55 80.44 86.66 86.85 (10, 11, 12)
-BNG(cmc) 57.40 57.78 56.19 56.72 57.72 57.88 (9, 10, 12)
-page-blocks 97.35 97.17 96.07 96.71 97.17 97.35 (6, 7, 12)
-segment 93.07 93.29 91.99 88.10 93.51 92.64 (1, 12)
-website_phishing 90.77 88.93 85.98 87.08 91.88 91.88 (7, 10)
-baseball 93.66 94.03 93.28 94.03 93.66 95.15 (10, 11)
-pendigits 99.50 99.36 92.81 93.04 99.41 99.45 (3, 4, 12)
-Gender_Gap_in_Spanish_WP 60.84 55.79 59.68 60.32 60.53 60.84 (2, 12)
-wine-quality-white 62.35 62.04 54.08 55.31 63.57 64.39 (8, 11, 12)
-satimage 91.21 92.07 88.72 88.65 91.91 91.91 (8, 11, 12)
-mfeat-fourier 90.00 89.75 77.75 82.25 89.50 89.50 (2, 7, 12)
-VulNoneVul 98.95 98.95 98.95 98.95 98.95 98.95 (1)
-law-school-admission-bianry 100.0 100.0 100.0 100.0 100.0 100.0 (6)
-KDD 80.34 78.95 79.34 78.35 81.23 79.94 (1, 8, 10)
-Appendix D Influence of Key Modules
-We investigate the influence of two key components in TabPFN v2, i.e., the feature engineering that pre-process the raw features of a given tabular dataset and the post-hoc ensembleing.
-Feature engineering.
-TabPFN v2 pre-processes the features of a given tabular datset with various strategies, such as quantile, category shuffling, SVD, and power transform.
-Specifically, we examine the effects of adding fingerprint features (add_fingerprint_feature) and polynomial features (polynomial_features) to the raw tabular data.
-The results indicate that TabPFN v2 performs well even without the use of these engineered features, suggesting that, for the benchmark datasets of Ye et al.
-(2024b), these specific feature engineering techniques do not provide a significant improvement.
-This finding highlights the robustness of TabPFN v2 and its ability to handle raw features effectively, without the need for extensive pre-processing or feature construction.
-We show the influence of this step in Figure 9.
-Figure 9: Scatter plot comparing the normalized Accuracy/R2 scores.
-The x-axis represents the normalized Accuracy/R2 scores without Feature Engineering, while the y-axis represents the normalized Accuracy/R2 scores with Feature Engineering.
-The red dashed line ($y=x$) serves as a reference, indicating equal performance.
-Model ensemble.
-Post hoc ensembling (PHE) involves applying TabPFN v2 to the datasets multiple times with different perturbations and aggregating the predictions of these base models at different temperatures.
-We show the change of performance of TabPFN v2 w.r.t.
-the number of ensemble numbers (i.e., the number of base models) in Figure 10.
-On the benchmark of Ye et al.
-(2024b), we observe that, overall, ensemble methods improve performance, with larger ensemble sizes yielding better results.
-However, we also note that even without ensembling, TabPFN v2 performs exceptionally well, and the relative performance gain from ensembling is limited.
-This suggests that while ensembling can provide further improvements, the base TabPFN v2 model is already highly effective on its own.
-The equivariant property described in (Arbel et al., 2025) provides insight into this phenomenon.
-Since TabPFN v2 introduces random tokens to handle heterogeneous features, the model becomes less sensitive to the arbitrary ordering of features, effectively enforcing equivariance in this aspect.
-As a result, the benefits of ensembling through feature order permutations are less pronounced compared to TabPFN v1.
-Figure 10: Box plot of relative performance improvements of TabPFN v2 with post hoc ensembling (PHE) across different ensemble sizes (2, 4, 8, and 16 base models).
-The relative improvement is calculated as the performance gain over the non-ensemble model, where higher values indicate stronger performance.
-The box plots show the median, interquartile range (IQR), and outliers for each ensemble size.
+“S” and “P” refer to unsupervised embedding extraction approaches by appending a column of dummy labels with zero values and permuting each column as labels, respectively, as described in Section C.2.
+PFN-v2 Vanilla S P layer-6 layer-9 layer-12 combined selected layers
+FOREX_audchf-day-High 77.38 50.68 56.95 69.48 68.39 73.57 74.11 77.11 (5, 9, 11)
+taiwanese_bankruptcy_prediction 96.99 56.45 96.77 95.75 97.14 96.77 97.07 97.14 (6)
+rl 85.51 50.00 60.56 70.82 66.90 69.52 86.72 87.53 (11, 12)
+pc3 89.46 10.22 89.78 86.90 90.10 88.82 88.82 88.82 (8)
+eye_movements_bin 61.83 50.00 55.12 57.95 59.72 59.40 62.16 62.16 (6, 9, 12)
+BNG(breast-w) 98.43 69.51 97.60 98.51 98.34 98.46 98.67 98.51 (6, 9)
+FOREX_cadjpy-hour-High 69.53 51.79 66.55 71.12 62.12 64.87 70.66 70.88 (4, 5, 6)
+dis 99.34 85.43 98.41 98.54 98.41 98.28 99.34 99.47 (4, 5, 6)
+sylvine 97.46 85.66 72.78 95.71 92.49 93.95 97.27 96.49 (1, 11)
+BNG(tic-tac-toe) 78.04 34.71 71.41 73.79 73.96 73.71 78.75 79.03 (5, 10, 12)
+online_shoppers 90.59 84.51 85.93 89.46 90.02 90.11 90.63 90.02 (8)
+Cardiovascular-Disease-dataset 72.84 50.86 68.73 72.60 72.96 73.06 73.14 73.09 (5, 8, 12)
+credit 78.04 62.31 75.86 78.31 77.62 77.80 77.95 77.59 (4, 6, 9)
+FOREX_audsgd-hour-High 67.26 51.48 65.49 70.14 57.24 61.06 69.62 70.41 (7, 10, 12)
+waveform-5000 86.00 80.60 55.70 87.10 85.60 85.60 86.40 86.90 (1, 6, 11)
+jungle_chess 85.65 39.60 64.12 72.14 78.55 80.44 86.66 86.85 (10, 11, 12)
+BNG(cmc) 57.40 42.62 52.48 55.16 56.19 56.72 57.72 57.88 (9, 10, 12)
+page-blocks 97.35 94.25 95.43 96.35 96.07 96.71 97.17 97.35 (6, 7, 12)
+segment 93.07 72.29 69.26 87.23 91.99 88.10 93.51 92.64 (1, 12)
+website_phishing 90.77 36.90 82.66 90.04 85.98 87.08 91.88 91.88 (7, 10)
+baseball 93.66 78.73 92.54 92.16 93.28 94.03 93.66 95.15 (10, 11)
+pendigits 99.50 59.75 72.40 98.18 92.81 93.04 99.41 99.45 (3, 4, 12)
+Gender_Gap_in_Spanish_WP 60.84 33.68 59.47 60.84 59.68 60.32 60.53 60.84 (2, 12)
+wine-quality-white 62.35 10.51 49.29 55.10 54.08 55.31 63.57 64.39 (8, 11, 12)
+satimage 91.21 82.04 84.99 89.19 88.72 88.65 91.91 91.91 (8, 11, 12)
+mfeat-fourier 90.00 55.50 46.75 85.75 77.75 82.25 89.50 89.50 (2, 7, 12)
+VulNoneVul 98.95 1.05 98.95 98.33 98.95 98.95 98.95 98.95 (1)
+law-school-admission-bianry 100.0 99.83 79.76 98.82 100.0 100.0 100.0 100.0 (6)
+KDD 80.34 78.45 62.36 76.76 79.34 78.35 81.23 79.94 (1, 8, 10)